### v1

# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

Abstract.
Cryptocurrencies and central bank digital currencies (CBDCs) are reshaping the monetary landscape, offering transparency and efficiency while raising critical concerns about user privacy and regulatory compliance. This survey provides a comprehensive and technically grounded overview of privacy-preserving digital currencies, covering both cryptocurrencies and CBDCs. We propose a taxonomy of privacy goals—including anonymity, confidentiality, unlinkability, and auditability—and map them to underlying cryptographic primitives, protocol mechanisms, and system architectures. Unlike previous surveys, our work adopts a design-oriented perspective, linking high-level privacy objectives to concrete implementations. We also trace the evolution of privacy-preserving currencies through three generations, highlighting shifts from basic anonymity guarantees toward more nuanced privacy-accountability trade-offs. Finally, we identify open challenges at the intersection of cryptography, distributed systems, and policy definition, which motivate further investigation into the primitives and design of digital currencies that balance real-world privacy and auditability needs.
1. Introduction
The rise of cryptocurrencies and the rapid development of central bank digital currencies (CBDCs) have fueled renewed interest in digital currencies. This momentum stems from the advances in decentralized technologies (i.e., blockchains and smart contracts), the introduction of decentralized finance, and the increasing concerns over data privacy and financial surveillance. While cryptocurrencies have demonstrated that permissionless, cryptographically secure value transfer is feasible, central banks are interested in modernizing monetary infrastructures with digital-native capabilities. However, these innovations raise fundamental questions about user privacy, institutional control, and regulatory compliance (e.g., (Auer et al., 2025; Ghesmati et al., 2022; Herskind et al., 2020)). Public blockchains typically offer pseudo-anonymity, using addresses decoupled from real identities. However, the transparency of public blockchains enables transactions tracing, address clustering, and ultimately users de-anonymization, thus exposing financial behavior and undermining confidentiality (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). As a result, privacy-preserving digital currencies have emerged as a critical research field, attracting attention from cryptographers, system designers, economists, and policymakers.
Researchers have developed a variety of privacy-enhancing technologies (PETs), including mixing protocols (e.g., (Maxwell, 2013; Ziegeldorf et al., 2018; Ruffing and Moreno-Sanchez, 2017)) and confidential transactions (e.g., (Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022)), and privacy-native cryptocurrencies (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Tomescu et al., 2022; Noether et al., 2016)). These solutions implement various notions of privacy, such as anonymity, confidentiality, and unlinkability. However, privacy alone is not sufficient: compliance with regulatory frameworks—such as Anti-Money Laundering (AML) and Combating the Financing of Terrorism (CFT)—requires mechanisms for auditability and selective disclosure. Therefore, recent work has explored cryptographic protocols such as zero-knowledge proofs (ZKPs), threshold encryption, and homomorphic commitments to reconcile these conflicting goals.
This survey proposes a unifying, design-oriented perspective on privacy-preserving digital currencies, spanning both decentralized cryptocurrencies and CBDCs. Unlike previous surveys that focus on broad overviews (e.g., (Genkin et al., 2018; Auer et al., 2025; Herskind et al., 2020)), specific technologies (e.g., (Almashaqbeh and Solomon, 2022; Ghesmati et al., 2022; Sun et al., 2021)), or do not consider auditability (e.g., (Peng et al., 2021; Herskind et al., 2020; Feng et al., 2019)), our work aims to bridge these dimensions. We introduce a systematic taxonomy of privacy and auditability properties and map them to concrete implementations across a broad range of systems. Structured as a hitchhiker’s guide, our survey defines key design goals and illustrates how they have been implemented in the literature through cryptographic primitives, protocols, and system architectures. Our contributions are threefold. First, we introduce a taxonomy that classifies privacy-preserving digital currencies along multiple dimensions, including system architecture, trust model, value representation, privacy guarantees, and auditability mechanisms. Second, we systematize how privacy properties are formalized and realized in digital currency systems, mapping high-level goals to technical mechanisms. We trace the evolution of privacy-preserving digital currencies across three generations. Third, we identify research challenges stemming from the design of privacy-preserving digital currencies, which lie at the intersection of cryptography, distributed systems, and privacy engineering. Our survey aims to organize this diverse and technically rich landscape, supporting both foundational research and practical protocol design in the rapidly evolving domain of digital currencies.
This paper is organized as follows. Section 2 reviews related surveys and outlines how our work complements and differs from them. Section 3 introduces the key cryptographic primitives used in digital currencies and cryptocurrencies. Section 4 presents a detailed analysis of privacy-preserving currency designs along key architectural and cryptographic dimensions. Section 5 outlines open challenges and future research directions, and Section 6 concludes the paper.
2. Related Surveys
A large body of literature has explored privacy-preserving digital currencies, motivated by both the technological challenges and the societal implications of financial surveillance. This includes surveys on PETs, systematization of cryptocurrencies, and privacy investigations for CBDCs.
Surveys on Privacy in Cryptocurrencies
Among the earliest works, Bonneau et al. (Bonneau et al., 2015) reviewed extensions of Bitcoin and introduced altcoins (i.e., alternatives to Bitcoin), including early privacy-native protocols. Genkin et al. (Genkin et al., 2018) offered a broad and accessible overview of techniques for anonymity in decentralized cryptocurrencies, while emphasizing the absence of a unified definition of privacy, which hinders fair comparisons between systems. More recent surveys have attempted to formalize anonymity notions. For instance, Amarasinghe et al. (Amarasinghe et al., 2019) developed a taxonomy of anonymity features, and evaluated Bitcoin-derived solutions against it. Our work draws from and extends this analysis by further considering the trade-offs between anonymity, confidentiality, and auditability across a broader range of systems, including CBDCs. Feng et al. (Feng et al., 2019) offered a layered classification of privacy-preserving mechanisms from network-level obfuscation to cryptographic primitives. While technically detailed, their coverage of cryptocurrency-specific protocols and privacy-accountability trade-offs is limited. Similarly, a systematic literature review by Herskind et al. (Herskind et al., 2020) catalogs techniques such as stealth addresses, confidential transactions, and network anonymity. However, it does not analyze core system features such as value representation models or system architectures. Peng et al. (Peng et al., 2021) identify PETs and illustrate how various protocols integrate them, but do not investigate protocol evolution or design trade-offs.
Other related surveys take a narrower or orthogonal perspective. Almashaqbeh and Solomon (Almashaqbeh and Solomon, 2022) focus primarily on privacy in blockchain computation, i.e., smart contracts, but do not consider concerns of accountability in payments. Zhang (Zhang, 2023) and Wang et al. (Wang et al., 2020) provide useful overviews of privacy in deployed cryptocurrencies, but lack systematic analysis of technical contributions. Alsalami and Zhang (Alsalami and Zhang, 2019) present a classification of PETs but do not address accountability aspects.
Privacy in CBDCs
A distinct body of work addresses privacy in CBDCs, driven by the need to balance individual privacy with institutional oversight. Auer et al. (Auer et al., 2022, 2025) distinguish between soft privacy, based on access controls and trust in intermediaries, and hard privacy, based on cryptographic techniques. While they offer a high-level classification and conceptual framework, their discussion does not consider low-level protocol mechanisms or specific cryptographic solutions. Darbha and Arora (Bank of Canada)111https://www.bankofcanada.ca/2020/06/staff-analytical-note-2020-9/ and Pocher and Veneris (Pocher and Veneris, 2022) explore PETs from a compliance and governance perspective. These works offer blueprints but lack detailed cryptographic analysis. Our work bridges this gap by providing a comparative review of cryptographic primitives, protocols, and system architectures that spans both CBDCs and cryptocurrencies.
Specialized Surveys
Several surveys focus on specific technologies or techniques. Examples include Bitcoin (Ghesmati et al., 2022), smart contracts (Benarroch et al., 2024; Li et al., 2022), DeFi (Baum et al., 2023), mixers (Arbabi et al., 2023), accountability mechanisms (Chatzigiannis et al., 2021), and ZKPs (Sun et al., 2021; Almashaqbeh and Solomon, 2022; Liang et al., 2025). These contributions do not consider the full design space of privacy-preserving digital currencies as we do, nor do they explore how different privacy features interact—particularly in settings requiring auditability. Koerhuis et al. (Koerhuis et al., 2020) present a forensic analysis of Monero and Verge, and highlight how implementation flaws threaten privacy; a concern we also emphasize.
Contribution of This Survey
Compared to previous works, our survey offers a comprehensive framework for analyzing privacy-preserving digital currencies. We propose a unified taxonomy that integrates both decentralized cryptocurrencies and CBDCs. Specifically, we present a systematic analysis of existing solutions to identify key value representation models, privacy features, and auditability guarantees. We focus on how such features are realized through cryptographic protocols and specific design choices, resulting in a classification of contributions that connects design goals to technical implementations.
3. Background
We first introduce key cryptographic primitives, including commitments, encryption schemes, and ZKPs. Then, we briefly review key cryptographic protocols used in privacy-preserving currencies, such as anonymity-enhanced signatures, multi-party computation, and one-time public keys.
3.1. Commitments and Homomorphic Commitments
A commitment scheme is a cryptographic primitive that lets someone commit to a value in a way that he cannot change it later. The scheme consists of two phases: (i) commit phase, where a commitment is made to a value , but the value is kept secret from third parties; (ii) opening phase, where the committed value is revealed. A commitment is hiding if an adversary cannot distinguish which of two values corresponds to the commit; it is binding if the committer cannot generate the same commit from two different values. These properties can be either computational, if the adversary is computationally bounded, or perfect, if not. A commitment scheme cannot be both perfectly hiding and perfectly binding, due to fundamental cryptographic limitations (Goldreich, 2004).
Homomorphic commitments, such as Pedersen (Pedersen, 1992) or ElGamal (Elgamal, 1985), preserve the addition or multiplication of corresponding values. For example, let and be elements of a prime group such that is not feasible to compute, let be a random value over and let be the committer’s value. Perdersen commitments are defined as . The homomorphic properties of Pedersen commitments guarantee that namely, the sum of the commitments of and is equal to the commitment of the sum of the two values under the sum of the two randomness and , allowing multiple operations on the committed messages. Pedersen commitments are perfectly hiding, computationally binding, and additively homomorphic; ElGamal commitments are computationally hiding, perfectly binding, and multiplicatively homomorphic.
3.2. Encryption Schemes
3.2.1. Identity-based Encryption
Identity-based encryption (IBE) schemes (Shamir, 1985) are public-key encryption schemes in which the public key is a string that uniquely identifies a user (e.g., an email). The corresponding private keys for decryption are derived from a master private key, by a trusted central authority. IBE can be used to simplify key management and recover, and access control within an organization. Anonymous variants make the public key indistinguishable from the ciphertext, effectively hiding the identity of the recipient of the message.
3.2.2. Threshold Encryption
Threshold encryption allows data to be encrypted in such a way that it can only be decrypted when a predefined quorum of participants agrees to reveal it. Threshold encryption can help maintain privacy while removing single point of failures. A popular implementation uses the ElGamal encryption scheme (Elgamal, 1985). A threshold-issuance variant of Boneh-Franklin IBE is used in UTT (Tomescu et al., 2022), to distribute the master secret key among multiple authorities.
3.2.3. Homomorphic Encryption
Homomorphic encryption schemes (Rivest et al., 1978) allow to compute additions, multiplications or both, without revealing the plaintext. Operations can be performed directly on ciphertexts such that the result corresponds to the encryption of the operation applied to the plaintexts, i.e., , where represent an encryption function and denotes either addition or multiplication. This allows to delegate computation to a third party while maintaining confidentiality. In payment systems (e.g., (Ma et al., 2021; Bünz et al., 2020)), it enables confidential transactions when used as a building block for ZKPs, like to homomorphic commitments.
3.3. Zero-knowledge Proofs
A ZKP is a protocol enabling one party (the prover) to convince another party (the verifier) that a statement is true, without disclosing any information beyond the validity of the statement itself. A dishonest prover can convince the verifier of a false statement, only with negligible probability (soundness). If the prover is computationally bounded (computational soundness), the proof is known as argument (Bitansky et al., 2012). ZKPs can be either interactive, requiring the parties to exchange messages, or non-interactive (NIZK), when the verifier is convinced through a single message. ZKPs can also be proofs of knowledge when the prover can convince the verifier that he knows a certain value, e.g., a secret key (or arguments of knowledge, if the prover is computationally bounded). If the ZK proof is short and fast to verify, it is succinct. If it requires no trusted setup, it is transparent.
A conceptual framework for ZK protocols is the -protocol (or Sigma-protocol), which is a type of 3-move interactive proof system: the prover commits a value, the verifier provides a challenge, and the prover provides a proof based on the commit and the challenge, so that they cannot choose a “malicious” proof. Using the Fiat-Shamir transform, -protocols can be made non-interactive (Fiat and Shamir, 1987). Building on this, various NIZK protocols have been designed, including zk-SNARK (Bitansky et al., 2012), Bulletproofs (Bünz et al., 2018), and—to some extent—zk-STARK (Ben-Sasson et al., 2019). Zk-SNARKs are succinct non-interactive arguments of knowledge. Many zk-SNARK constructions have been proposed in past years (e.g.,(Bitansky et al., 2012; Boneh et al., 2020; Groth, 2016; Chen et al., 2023)), with different approaches also for setup. Among them, the most popular approach is Groth16 (Groth, 2016), an instance optimized for performance but with some trade-offs like requiring a trusted setup. Bulletproof is optimized for range proof and does not require trusted setup. zk-STARK is a recent approach offering efficient proving times and no need for trusted setup. STARKs are succinct transparent arguments of knowledge. However, in the current state, zk-STARK generates large proof size (a few hundred kilobytes), so they are not as succinct as SNARKs. A recent survey on ZKPs can be found, e.g., in (Christ et al., 2024; El-Hajj and Oude Roelink, 2024). We usually refer to NIZK as ZK.
3.4. Anonymity-enhanced Signatures
3.4.1. Blind Signatures
Blind signatures (Chaum, 1983) are a form of digital signature in which the signer produces a valid signature on a message without learning its content: before signing, the message is blinded, i.e., transformed through a random element that makes it unrecognizable. Once the contents of the signed message are unblinded, they can be publicly verified as usual.
3.4.2. Randomizable Signatures
Randomized signatures provide a way to randomize a previously generated signature, transforming it into a new valid signature, but making it impossible to link the two instances together. This enables a level of anonymity by unlinking user identities from their transaction history, as each transaction appears distinct even if performed by the same user. The original idea was proposed by Camenisch and Lysyanskaya (Camenisch and Lysyanskaya, 2003, 2004) but suffered from a linear size in the number of messages to be signed. Later, Pointcheval-Sanders proposed more efficient signature and verification algorithms (Pointcheval and Sanders, 2016). Existing works (e.g., (Baudet et al., 2023; Wüst et al., 2019; Sarencheh et al., 2025; Tomescu et al., 2022)) adopt either Pointcheval-Sanders (Pointcheval and Sanders, 2016) or Coconut (Sonnino et al., 2020), a threshold scheme derived from it. The Pointcheval-Sanders scheme has the ability to produce signatures of a message directly from its commitment. This effectively makes it a blind signature, thanks to the hiding property of commitments.
3.4.3. Group Signatures
Group signature schemes (Chaum and van Heyst, 1991) enable any member of a designated group to generate a valid signature without revealing who signed (ensuring anonymity). A verifier uses a group public key to confirm that a valid signature was produced by some group member. This property also implies unlinkability: it is not possible to determine whether two different signatures were produced by the same user (Perera et al., 2022). However, group signatures can be made linkable to prevent double spending in transactions (e.g., (Zhang et al., 2019)). Group signatures also support user traceability: a designated group manager holds a group secret key that enables him to revoke a signer’s anonymity and reveal his identity when necessary.
3.4.4. Ring Signatures
Ring signatures (Rivest et al., 2001) are similar to group signatures, as they hide the signer’s identity behind a group. Differently from group signatures, user anonymity is permanent, because there is no revoke mechanism. Indeed, the signer simply joins their public key with a set of other public keys during the signing phase. A verifier needs the whole list of public keys to verify the validity of the signature, and cannot discern which one of the participants corresponding to the public keys is the signer, but can only guess with probability . There are some variant ring signatures, where anonymity is somewhat limited and the signer can be traced under special circumstances, e.g., after double spending. These are called traceable ring signatures (Fujisaki and Suzuki, 2007; Fujisaki, 2011). Ring signatures can also be linkable (Liu and Wong, 2005), if they allow anyone to determine if two signatures were produced by the same signer (e.g., (Van Saberhagen, 2013; Li et al., 2021)).
3.4.5. Threshold Signatures
Threshold signatures are a type of signature that can be jointly produced by a committee of signers. A -of- signature scheme requires at least signers out of a total of participants to produce a valid signature. Works that use threshold signatures (Androulaki et al., 2020; Sarencheh et al., 2025; Kiayias et al., 2022; Baudet et al., 2023) use threshold blind signatures, based on Coconut (Sonnino et al., 2020).
3.5. Secure Multi-party Computation
Secure Multi-Party Computation (MPC) allows multiple parties to jointly compute a result over their private inputs without any party having to reveal its own input to the others. The key idea is that the parties collaborate to perform some computation and, at the end, they all learn the correct output of the function, but nothing else. This technique is used, e.g., in CoinParty (Ziegeldorf et al., 2018).
3.6. One-time Public Keys
3.6.1. Stealth Addresses
A stealth address hides the recipient’s identity by making each payment unlinkable. It is a one-time public key generated by the transaction sender with public key provided by the transaction receiver. Recipient does not update his public key, but the sender can derive fresh one-time keys using stealth addresses. Let be a generator of an elliptic curve group with scalar field . Each receiver has a public view key and a public spend key , where are secret scalars. A sender chooses a random scalar and computes the shared secret , with a hash-to-scalar function, and the one-time public key . Then, the sender includes and in the transaction. The recipient scans the blockchain and for each and computes . If then the output belongs to the recipient.
3.6.2. Updatable Public Keys
Updatable public keys allow to update the recipient’s public key in a decentralized manner, producing dynamic and evolving public keys. Quisquis (Fauzi et al., 2019) introduces this technique to update both sender and recipient public keys. Formally, given a prime-order group with the prime order of the group, and its generator, each user has a secret key and a public key , for some random value . To perform a transaction, the sender updates the public key by choosing a random value and computing . The updated key is indistinguishable from a fresh key, thanks to the Decisional Diffie-Hellman assumption. The recipient verifies whether a given updated public key corresponds to his secret key . This works because updated keys still verify correctly with the original secret key.
4. Design Dimensions of Privacy-Preserving Digital Currencies
Dimension of analysis along which we review privacy-preserving digital currencies.
The research community has explored various solutions to ensure anonymity, confidentiality, and auditability in digital currencies. We examine the most relevant contributions from the literature, analyzing how they address key design challenges. To provide a comprehensive overview of the diverse approaches studied so far, we identify key dimensions of analysis—protocol, value representation, privacy features, and auditability—and highlight the most significant design choices. These interconnected dimensions embed privacy features into the system’s design, linking them to the ledger structure, value representation, fund exchange, and auditability. Figure 1 outlines the dimensions that structure this section. Table LABEL:tab:sota-cryptocurrencies summarizes key contributions, whereas Table 1 provides a description of the labels used in our analysis.
| Dimension | Legend | |||
| Protocol | ||||
| Type | CD | CBDC | D | Digital currency |
| Blockchain Compatibility | B | Bitcoin | E | Ethereum |
| Value Representation | ||||
| Data Model | A | Account-based | CT | Chaumian token |
| T2B | Transaction per bank table | UT | Unspent token | |
| Other | ||||
| Auditability | AB | Anonymity budget | L | Operating limits |
| RA | Revocable anonymity | TR | Coin tracing | |
| ZD | Verification without disclosure | |||
| Cryptographic Primitives | BS | Blind signature | C | Commitment |
| DS | Digital signature | GS | Group signatures | |
| HC | Homomorphic commitment | IBE | Identity-based encryption | |
| MS | Multi signature | MT | Merkle tree | |
| O-RAM | Oblivious RAM | PE | Public key encryption | |
| RS | Ring signature | SA | Stealth address | |
| UP | Updatable public keys | ZK | Zero-knowledge proof | |
| Threshold-based version of | Randomizable version of | |||
| * | Custom variant of | |||
| Generics | ✓ | Yes | No | |
| ◐ | Partial | Alternative | ||
| ○ | Abstracted | |||
4.1. Protocol
Over time, different generations of approaches have emerged, each addressing specific features and evolving needs. These developments have been driven by key factors, including the introduction of blockchain technology, advancements in cryptographic primitives, and the growing interest in designing CBDCs. The protocol dimension provides a broad perspective on the diverse range of existing solutions.
4.1.1. Generations of Privacy-preserving Digital Currencies
The term digital currency encompasses a range of different forms of money that is managed, stored, and exchanged on digital computer systems over the Internet222https://en.wikipedia.org/wiki/Digital_currency. Cryptocurrency and CBDC represent two specific forms of digital currency. A cryptocurrency is defined, stored, and exchanged over peer-to-peer systems, such as the blockchains. A CBDC is digital currency issued by a Central Bank (Bech and Garratt, 2017), where the term is often used to indicate that the currency generally does not have a classical physical form of a fiat currency. We identify three main generations of solutions for privacy-preserving digital currencies.
First Generation
The first generation focuses on the design of digital currencies leveraging digital signatures and a centralized authority (i.e., the issuing bank) (Chaum, 1983). The fundamental problem relates to guaranteeing anonymity of sender (i.e., the payer) although he must interact with the central bank to obtain valid tokens (representing coins) to perform subsequent payments. Chaum (Chaum, 1983) proposes one of the first attempts to construct a digital currency, commonly referred to as ecash. Later, several works improved ecash with different features, including support for different coin denominations (e.g., (Dold, 2019)), partial spending (e.g., (Dold, 2019)), token transferability (e.g., (Hayes, 1990; Sai Anand and Madhavan, 2000; Baldimtsi et al., 2015; Bauer et al., 2021; Chase et al., 2014)), offline spending (e.g., (Chaum et al., 1990)), and auditability (e.g., (Camenisch et al., 2006; Dold, 2019)).
Second Generation
The second generation of works emerged with the introduction of Bitcoin (Nakamoto, 2008). Bitcoin and, later, Ethereum (Buterin et al., 2013) propose public and permissionless blockchains, meaning that the ledger contains transactions visible to everyone and anyone can participate in the process of updating the public ledger by following a specific (validation and consensus) protocol. Being public, these blockchains use addresses (acting as pseudonyms) to exchange money in a pseudo-anonymous manner, as addresses are not directly tied to user identities. Public blockchains cannot easily and fully satisfy two privacy-related features: untraceability and unlinkability (Rahalkar and Virgaonkar, 2024). Untraceability ensures that, for any given incoming transaction, all potential senders are equally likely. Unlinkability means that, for any two outgoing transactions, there is no way to prove they were sent to the same recipient. While generating a new address for each transaction can enhance these properties, various studies indicate that achieving true anonymity and unlinkability on public blockchains remains challenging. An analysis of the transaction graph can reveal clusters of pseudonyms or patterns that link with high probability two or more addresses (e.g., (Meiklejohn et al., 2016; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). The second generation of solutions mainly focuses on improving unlinkability, anonymity, and confidentiality, often considering decentralized architectures (e.g., (Miers et al., 2013; Van Saberhagen, 2013; Fauzi et al., 2019; Baudet et al., 2023; Friolo et al., 2024; Sarencheh et al., 2025; Androulaki et al., 2013; Herrera-Joancomartí, 2015; Meiklejohn et al., 2013)).
We classify this second generation of approaches into three main categories: privacy-native protocols, mixing protocols, and off-ledger solutions. Changing the protocol of existing blockchains is hard, as it involves forks. Therefore, privacy-native protocols define (mostly) novel currencies with improved confidentiality, anonymity, and unlinkability features (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Ben Sasson et al., 2014; Tomescu et al., 2022)). Conversely, mixing protocols propose supplementary services for permissionless blockchains aimed to collect a group of transactions, scramble senders and receivers, and publish them as a single group transaction (e.g., (Maxwell, 2013; Bissias et al., 2014; Bonneau et al., 2014; Valenta and Rowan, 2015; Ruffing and Moreno-Sanchez, 2017; Ziegeldorf et al., 2018; Qin et al., 2023; Wang et al., 2023)). Mixing protocols break the link between sender and receiver, thus reducing the possibility of linking together addresses of the blockchain (i.e., they enhance unlinkability of transactions). Off-ledger solutions are a specific approach related to blockchains (which serve as Layer 1) and the ability to build decentralized peer-to-peer Layer 2 systems on top of them. Layer 2 solutions process transactions off the main chain, reducing the need to record every transaction on the shared ledger (e.g., (Qin et al., 2023; Wang et al., 2023; Baudet et al., 2023; Seres et al., 2019)). Since we focus on on-ledger privacy, we omit a detailed analysis of Layer 2 privacy solutions.
Third Generation
While protecting the privacy and anonymity of payers and payees remains essential, the third generation of privacy-preserving digital currencies introduces mechanisms that balance these guarantees with auditability and regulatory compliance (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Narula et al., 2018; Li et al., 2021; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Yu et al., 2024; Xue et al., 2023)). These approaches are particularly motivated by concerns around AML and CFT. Auditability refers to the system’s ability to track, verify, or selectively disclose information about transactions, without fully compromising user privacy. This capability enables the enforcement of specific policies that apply to users, coins, or transactions behaviors over time. To achieve this, modern privacy-preserving systems integrate mechanisms such as revocable anonymity, trapdoor-based decryption, and ZKPs for correctness and compliance.
4.1.2. System Architecture
We examine the system architectures underlying privacy-preserving digital currencies and mixing protocols, which enhance transaction unlinkability.
Centralized Systems
The first generation of privacy preserving currencies typically rely on centralized architectures involving up to three or four trusted authorities, each responsible for distinct functions. They include: (i) a central bank, which issues digital coins and maintains monetary control; (ii) a tracing authority, responsible for auditing and tracing under specific conditions; (iii) a registration authority, in charge of user on-boarding and identity verification; and, in some cases, (iv) a technical support entity, which collaborates with the tracing authority to enable conditional de-anonymization. Most early proposals assume a single central bank (e.g., (Chaum, 1983; Dold, 2019; Cecchetti et al., 2017; Zhang et al., 2019; Beer et al., 2024)). For privacy, special cryptographic primitives are used, such as blind signatures (Chaum, 1983), which allow the bank to issue valid coins without learning user identities. To support auditability and regulatory compliance, systems often include a separate tracing authority that can decrypt transactions or enforce specific policies (e.g., check or bypass an anonymity budget). This entity operates independently from the central bank, thereby preserving some separation of duties (e.g., (Sander and Ta-Shma, 1999; Wüst et al., 2022)). Yu et al. (Yu et al., 2024) introduce a technical support authority that must cooperate with the tracing authority to extract sender identities. In the context of CBDCs, system architectures usually include a registration authority, responsible for enrolling user registration and managing identity records (e.g., (Zhang et al., 2019))—an essential requirement considering regulatory frameworks (e.g., AML and CFT).
Decentralized Systems
Blockchains introduced a fundamental shift in the design of digital currencies by leveraging decentralization and cryptographic protocols. Decentralization serves two main purposes. First, it enhances resilience against malicious actors or censorship by distributing trust across multiple entities. This is often supported by cryptographic techniques such as threshold encryption, secret sharing, and multi-signature schemes. Second, it enables intermediary-free value exchange, using commitments and digital signatures to ensure integrity, authenticity, and non-repudiation. Many privacy-native solutions adopt decentralized architectures in line with these principles (e.g., (Noether et al., 2016; Van Saberhagen, 2013; Poelstra, 2016; Androulaki et al., 2020; Narula et al., 2018; Ma et al., 2021; Xue et al., 2023)). Moreover, the programmability of blockchains has enabled privacy-preserving overlays (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)) and smart contracts333Although smart contracts are logically centralized, they are executed on each node; therefore, we consider smart contracts as a decentralized solution. implementing private tokens “pegged” to the coin of the primary chain (e.g., (Bünz et al., 2020; Baudet et al., 2023; Sarencheh et al., 2025)).
Architecture of Mixing Protocols
Different architectural styles have been explored for mixing protocols, including both centralized (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015; Heilman et al., 2016; Qin et al., 2023; Liu et al., 2018)) and decentralized solutions (e.g., (Ruffing et al., 2014; Ziegeldorf et al., 2018; Bissias et al., 2014; Ruffing and Moreno-Sanchez, 2017; Nosouhi et al., 2023; Wang et al., 2023)). In centralized solutions, a third party (often called mixer or tumbler) provides the mixing service for a fee. While such designs are simple and support large anonymity sets, they introduce risks such as availability, theft, and potential privacy breaches. Accordingly, different trust models have been proposed to mitigate these concerns (see Section 4.1.4). Decentralized mixers remove the need for a trusted party by using cryptographic protocols or peer-to-peer coordination, enhancing security and censorship resistance, but they often incur higher complexity, require user coordination, and typically support smaller anonymity sets. Notable examples include Xim (Bissias et al., 2014), CoinShuffle++ (Ruffing et al., 2016), and ValueShuffle (Ruffing and Moreno-Sanchez, 2017). Both centralized and decentralized approaches face challenges such as Sybil attacks and regulatory scrutiny. In a Sybil attack, an adversary inject fake participants into the mix to degrade anonymity.
4.1.3. Protocol Integration and Compatibility with Existing Blockchains
Although based on blockchain principles, most second-generation privacy-preserving cryptocurrencies are not fully compatible with baseline systems. We analyze how they interact with existing digital currencies, whether through compatibility, non-compatible extensions, or entirely different approaches.
Compatible Solutions
Mixing protocols propose a supplementary services for existing blockchains, in particular for Bitcoin, and are specifically designed to be fully compatible (e.g., (Ruffing et al., 2016; Valenta and Rowan, 2015; Qin et al., 2023; Ruffing and Moreno-Sanchez, 2017)). Another category of compatible solutions leverages blockchain programmability to introduce privacy-enhanced currencies through overlays (e.g., (Bünz et al., 2020; Baudet et al., 2023)). For instance, Zether (Bünz et al., 2020) employs a smart contract on Ethereum to store and manage private accounts, whereas Zef (Baudet et al., 2023) introduces opaque coins in the form of off-chain certificates exchanged between Ethereum user accounts.
Non-compatible Extensions
Most privacy-enhancing currency proposals result in non-compatible extensions of existing blockchain protocols. For example, Zerocoin (Miers et al., 2013) is an overlay over Bitcoin that allows user to convert bitcoins in zerocoins. A zerocoin can be proved (using ZK techniques) to originate from a valid and unspent bitcoin, without revealing the link to the original bitcoin. Zerocoin does not provide full anonymity: it obscures coin traceability, but it does not hide the number of transactions, recipient addresses, and balances. Building on Zerocoin, Zerocash (Ben Sasson et al., 2014) enhances privacy by concealing both sender and recipient addresses, supporting arbitrary transaction values, and enabling change handling. Both Zerocoin and Zerocash require new transaction types and payment semantics, and break compatibility with Bitcoin. Garman et al. (Garman et al., 2017) extends Zerocash with no backward compatibility, by introducing policy enforcement, selective user tracing, and tracing of tainted coins—features that further diverge from Bitcoin’s existing design. Yu et al. (Yu et al., 2024) propose an overlay on Bitcoin for anonymous yet regulated transactions using a coin backed by bitcoins. Their approach relies on three centralized entities that cooperate to enable batch linkability and payer tracing while preserving privacy. These changes are not backward-compatible with Bitcoin.
Novel Approaches
Most approaches present greenfield solutions for privacy-native cryptocurrencies (e.g., (Van Saberhagen, 2013; Poelstra, 2016; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020)), often extending open-source solutions (e.g., (Li et al., 2021; Papadoulis et al., 2024)). Both centralized (e.g., (Cecchetti et al., 2017; Zhang et al., 2019)) and decentralized architectures (e.g., (Narula et al., 2018; Androulaki et al., 2020; Lin et al., 2020; Ma et al., 2021; Goodell et al., 2023; Elfadul et al., 2024)) for privacy-native solutions have been proposed, with a sensible preference for the latter. Early approaches like CryptoNote (Van Saberhagen, 2013), MimbleWimble (Poelstra, 2016), and Monero (Noether et al., 2016) were among the first to address Bitcoin’s privacy limitations by still proposing fully decentralized alternatives. CryptoNote (Van Saberhagen, 2013) enhances sender and receiver anonymity through one-time ring signatures and NIZK proofs, inspiring cryptocurrencies like ByteCoin, DigitalNote, and Aeon (Amarasinghe et al., 2019). MimbleWimble (Poelstra, 2016) improves privacy and scalability by aggregating confidential transactions instead of storing individual transaction data. Cryptocurrencies building on MimbleWimble include Grin444https://grin.mw/ and Beam555https://www.beam.mw/. Monero (Noether et al., 2016) strengthens privacy with stealth addresses, ring signatures, and confidential transactions, while also integrating network anonymity protocols.
Other approaches propose partially decentralized solutions (e.g., (Wüst et al., 2019)), combine concepts from centralized mixers (e.g., (Fauzi et al., 2019)), or consider hierarchical bank-intermediated settings (e.g., (Cecchetti et al., 2017; Narula et al., 2018)). Wüst et al. (Wüst et al., 2019) propose PRCash to strike a trade-off between performance (fast payments), user privacy, and regulatory oversight. Basically, PRCash includes two centralized entities, which respectively issue money and oversee compliance, and a permissioned blockchain, which records MimbleWimble-like transactions. Quisquis (Fauzi et al., 2019) builds on the concept of mixing to redistribute wealth: each transaction updates a bunch of other accounts on chain, thus preventing addresses from appearing multiple times on the blockchain, reducing the risk of de-anonymization and linkability. Quisquis ensures integrity through ZKPs while maintaining privacy using commitments and updatable public keys. Cecchetti et al. (Cecchetti et al., 2017) propose Solidus, a hierarchical, bank-intermediated systems where a small number of banks manage transactions of on-chain assets on behalf of a large number of users. This setup ensures transaction-graph confidentiality: sender and receiver of a transaction cannot be publicly identified666While banks can identify their respective clients, other entities only learn the identities of the banks., even by pseudonyms.
4.1.4. Trust Assumptions
We now focus on the degree of trust required, such as reliance on a trusted third party, to facilitate secure transactions. Most solutions for digital and cryptocurrencies do not usually rely on trusted parties for confidentiality (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Li et al., 2021; Ma et al., 2021; Guan et al., 2022; Tomescu et al., 2022; Baudet et al., 2023; Nosouhi et al., 2023; Sarencheh et al., 2025; Wang et al., 2023)), even when they propose a centralized system (e.g., (Chaum, 1983; Chaum et al., 2021; Wüst et al., 2022)).
The solutions that introduce trust assumptions do it for different purposes. The approaches that use zk-SNARK as ZKP require a trusted setup to initialize and correctly share some cryptographic parameters (e.g., (Miers et al., 2013; Garman et al., 2017; Ben Sasson et al., 2014; Xue et al., 2023)). A few works require a trusted central bank to issue coins that can be later exchanged anonymously (e.g., (Goodell et al., 2023; Friolo et al., 2024)). Elfadul et al. (Elfadul et al., 2024) assume the presence of a regulatory supervision that grants permissions to users to perform transaction; this centralized authority could asymmetrically revoke such authorization. Contributions that explore special forms of auditability related to revocable anonymity (e.g. (Wüst et al., 2019; Lin et al., 2020; Xue et al., 2023; Yu et al., 2024)), trapdoor (e.g., (Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Elfadul et al., 2024)), or user tracing (e.g., (Garman et al., 2017)) require trust in the tracing authority that uncovers transaction details only when necessary. PEReDi (Kiayias et al., 2022) reduces trust requirements by relying on quorum of centralized entities that must agree to perform specific operations, such as decrypting transaction information.
As regards centralized mixing protocols, since the third party represents a single point of failure, different protocols focused on different threat models, where the mixer may be trusted and accountable (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015)) or untrusted (e.g., (Heilman et al., 2016)). For instance, Mixcoin offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Conversely, Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. The rise of off-chain solutions has let to the exploration of alternative mixer definitions. For instance, BlindHub (Qin et al., 2023) implements mixing within payment channel hubs, where the hub acts as the tumbler. It employs blind signature-based protocols to hide users’ addresses from the hub, ensuring privacy.
4.1.5. Interaction Models
Privacy-preserving digital currencies differ in the extent to which transactions require direct or indirect interaction between the payer and the payee. Interactive payment protocols require active cooperation between payer and payee during the transaction process. This is typical in many account-based models where the recipient must be involved in constructing or approving the updated account state, often through ZKPs or digital signatures. For instance, PEReDi (Kiayias et al., 2022) and Platypus (Wüst et al., 2022) enforce synchronized interaction: both parties participate in updating the ledger with proofs and commitments, ensuring mutual accountability and preventing attacks such as dusting or replay. Similarly, Chaumian token-based systems often require the payee to claim tokens immediately after receipt to prevent double spending (e.g., (Dold, 2019)). In contrast, non-interactive payment protocols allow the payer to unilaterally construct a transaction and broadcast it to the network. Many token-based systems, especially those based on unspent tokens, adopt this strategy (e.g., (Ben Sasson et al., 2014; Van Saberhagen, 2013; Noether et al., 2016; Poelstra, 2016)). This model improves user convenience and enables asynchronous payments, but may require periodic ledger scanning or local computation by the receiver777Off-chain communication can alleviate the burden of scanning the ledger, but current protocols lack this feature.. Some account-based protocols implement delayed claiming to approximate this non-interactive behavior. Sarencheh et al. (Sarencheh et al., 2025) implement a non-interactive variant of PEReDi, which works in two stages. In the transfer stage, the sender interacts with a set of issuers to update his account and provide information in ZK. The claim stage credits the receiver account, exploiting information enclosed in ZKP by the sender. While interactive designs enable rich policy enforcement and conditional privacy, non-interactive models better suit high-latency or asynchronous networks.
4.2. Value Representation
The value representation dimension examines the approaches used to represent the currency, the user balance, and the exchange of value. For the purpose of this survey, we investigate these aspects with a specific focus on their relation with privacy-related features.
4.2.1. Ledger Data Models
Digital currencies and cryptocurrencies follow three main approaches for representing and tracking ownership of digital assets: accounts, tokens, and tabular ledger.
Account-based
Account-based ledgers associate a balance with each account address, which is debited or credited by transactions. This simplifies state management and enables simple definitions of policies based on cumulative balances. Nonetheless, it requires safeguards against replay and double-spending attacks, often implemented by storing a sequence number (or nonce) in each account and updating it with every transaction. For example, Ethereum (Buterin et al., 2013) associates to each account888Ethereum also defines standards to represent fungible tokens (ERC-20) and non-fungible tokens (ERC-721). ERC-20 can be used to design privacy-preserving currencies (e.g., (Bünz et al., 2020)). a nonce (indicating the number of sent transactions), a balance, and two hashes referencing the account’s storage contents and associated code. Since the balance of an account is explicitly represented, obtaining anonymity and unlinkability is challenging, especially in systems where users cannot freely generate multiple addresses (e.g., (Guan et al., 2022)). Privacy-preserving account-based solutions replace plain-text balances with commitments, achieving confidentiality of balances and transactions (e.g., (Fauzi et al., 2019; Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022; Sarencheh et al., 2025)). Liu et al. (Liu et al., 2022) propose an account-based CBDC with sender and receiver anonymity but with no unlinkability. The system follows a two-tier model where the central bank can identify users via encrypted identifiers, while commercial banks cannot. To guarantee anonymity while ensuring unlinkability, most of the approaches change users’ account address at each transaction and use an authenticator to prove correctness. The authenticator can be implemented, e.g., using digital signatures by trusted or centralized authorities (e.g., (Kiayias et al., 2022)) or ZKPs (e.g., (Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)). Platypus (Wüst et al., 2022) rely on ZKPs; conversely, in PEReDi (Kiayias et al., 2022), account states must be (blindly) signed by a quorum of maintainers, and since each transaction reveals prior account snapshots to construct new authenticators, randomized signatures are used to prevent linkability. Quisquis (Fauzi et al., 2019) and AQQUA (Papadoulis et al., 2024) use updatable public keys to update accounts in a decentralized manner. The account-based model lends itself well to auditability, as it allows simple definition of policies on holding limits (e.g., (Gross et al., 2021; Wüst et al., 2022; Papadoulis et al., 2024)) or spending and receiving limits (e.g., (Papadoulis et al., 2024)). For example, the account state by Gross et al. (Gross et al., 2021) includes an accumulator of all spending in the current epoch, enabling policies to rely on this information. AQQUA (Papadoulis et al., 2024), instead, includes commitments to three different counters (i.e., balance, total amounts sent and received). Conversely, Platypus (Wüst et al., 2022) leaves the specification of the additional information to be application dependent.
Token-based
Token-based ledgers do not maintain explicit balances per address, but instead track individual tokens. The user’s balance can be derived from the sum of all owned tokens. This model enables parallel validation of user transactions and stronger privacy guarantees (as balances are not explicitly materialized) but introduces management complexity, particularly when users distribute their coins across multiple addresses for added privacy. There are two main types of token structures: Chaumian tokens and unspent tokens. Although both represent digital value, they differ in structure, spending mechanisms, privacy guarantees, and centralization.
Chaumian tokens (Chaum, 1983) use blind signatures and were first deployed in centralized, bank-issued electronic cash schemes. When issuing a token, the bank blindly signs a random value provided by the payer, allowing the payer to later unblind it for spending. Blind signatures prevent the bank from linking withdrawals to subsequent transactions, thus ensuring sender anonymity. These tokens rely on a trusted authority to issue and verify their validity; to this end, the authority must hold an ever increasing list of spent tokens that might introduce scalability challenges. Some implementations introduce expiration to limit circulation (e.g., (Dold, 2019; Lovejoy et al., 2023)). The Chaumian model inspired decentralized privacy-preserving cryptocurrencies (e.g., Monero (Noether et al., 2016), Zerocoin (Miers et al., 2013), and Zerocash (Ben Sasson et al., 2014)), which eliminate the need for a central issuer by leveraging a combination of different cryptographic techniques. For example, Zerocash supports transparent (like Bitcoin) and shielded transactions, with the latter hiding sender, recipient, and amount. Shielded payments encrypt these values with one-time keys and use zk-SNARK proofs to validate inputs-outputs equality, concealing all on-chain details.
Unspent tokens, introduced in Bitcoin (Nakamoto, 2008), represent transaction outputs that can be used as inputs for future transactions, which in turn create new unspent tokens. As such, they are inherently traceable unless PETs are used (Baldimtsi et al., 2024). Unspent tokens are issued in a decentralized manner, usually as result of the mining process. Their core attributes include an identifier, value (or its commitment), and ownership information. Their recording in a public, transparent ledger prevents double-spending but complicates privacy. Privacy-focused cryptocurrencies employ stealth addresses, ring signatures, confidential transactions, and ZKPs to obscure transaction details and unlink payments (e.g., (Noether et al., 2016; Poelstra, 2016; Miers et al., 2013; Ben Sasson et al., 2014))—these properties are discussed in more detail later. Tokens can also include auxiliary data to support the evaluation of specific policies (e.g., (Xue et al., 2023; Yu et al., 2024)). For example, Xue et al. (Xue et al., 2023) include information to prove in ZK that the total amount of the cryptocurrency transferred and the transaction frequency in a time period is limited. For a detailed discussion of auditability in token-based systems, see Section 4.4.
Tabular Ledger
Narula et al. (Narula et al., 2018) propose zkLedger, a third approach tailored to a specific use case: a bank-intermediated CBDC. It uses a tabular ledger where rows represent transactions and columns represent banks. Each cell contains a Pedersen commitment to hide values, while enabling cryptographic verification. Transactions include ZKPs to prove balance, asset ownership, and consistency with audit records. This structure supports rich privacy-preserving auditing: the system includes an auditor who can issue a rich set of auditing queries—using primitives like sums, moving averages, variance, standard deviation, ratios—and receive answers that are provably consistent with the ledger. As downside, this structure does not scale well with many banks and transactions, making it suitable for small, privacy-focused networks.
4.2.2. Ledger State Update
Updating the ledger is another operation that may reveal sensitive information to other parties. We review the approaches for account-based and token-based ledgers.
Account-based
A generic state update transaction for account-based ledger includes the sender and receiver addresses, the exchanged amount, and the sender digital signature. For privacy-enhanced account-based, a naïve state update would reveal the updated accounts. Therefore, privacy-preserving state updates in account-based ledger usually adopt three strategies (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)): to obtain confidentiality, they update a commitment of the new account state; to obtain unlinkability, they always update the public key associated with the (recipient) account; and to obtain anonymity, they update a group of accounts at the same time.
Token-based
Token-based ledgers store the list of valid (unspent) tokens or burnt (spent) tokens. A spending operation must invalidate an existing unspent token, to prevent double-spending. With Chaumian tokens (e.g., (Chaum, 1983; Dold, 2019; Lovejoy et al., 2023)), payment systems include a withdrawal operation, which creates a new valid token that the payer can spend later in time, and a spend operation. The state of the system is represented by the list of spent tokens, which is updated at spending time; unspent tokens are usually not explicitly stored (if not by payers). Privacy-preserving systems based on Chaumian tokens do not usually guarantee transaction confidentiality (e.g., (Chaum, 1983; Hayes, 1990; Sander and Ta-Shma, 1999; Dold, 2019; Goodell et al., 2023; Friolo et al., 2024)), therefore they still maintain the list of spent tokens. Unlike other approaches, Sander and Ta-Shma (Sander and Ta-Shma, 1999) represent valid tokens in a Merkle tree. The bank does not blind sign tokens; instead, valid tokens come with a hash chain proving membership in the Merkle tree. The Merkle root is publicly available to ensure transparency and auditability. Withdrawal coins can be invalidated if needed, addressing certain attack scenarios (e.g., blackmail). Friolo et al. (Friolo et al., 2024) employ a ZKP-based protocol to achieve unlinkability. Transactions are represented as commitments, and when spent, their opening is revealed along with a ZKP demonstrating that it is linked to a valid, unspent token. In contrast, UTT (Tomescu et al., 2022) ensures transaction confidentiality, and it recurs to nullifiers to track spent coins. Nullifiers enable the bank to detect double-spending without knowing transaction specifics. In (Tomescu et al., 2022), a nullifier is a pseudo-random value computed from the coin’s serial number (and a secret). The system checks their validity through ZKPs and, if valid, stores them in a list for future checks.
With unspent tokens, the payment operation simultaneously invalidates previous tokens and generate new ones. The state of the system is represented by the list of unspent tokens999Since blockchains store transactions in an append-only log, they trace the derivation history of each unspent token. This helps implement a decentralize yet scalable system.. Unspent tokens cannot be easily marked as spent without compromising sender’s anonymity. Most of the approaches, including CryptoNote (Van Saberhagen, 2013), Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), and their extensions (e.g., (Garman et al., 2017; Androulaki and Karame, 2014; Xue et al., 2023)), let the payer spend a token by using a serial number and a ZKP that indicates that the number is a valid opening to some commitment stored in an unspent token. Monero (Noether et al., 2016) (and its extension Traceable Monero (Li et al., 2021)) uses one-time signatures for unspent tokens, which can be used only once to spend the token. At spending time, the ledger registers a salted hash of the public key. MimbleWimble (and PRCash that reuses its transaction format (Wüst et al., 2019)) eliminates the idea of performing transactions between addresses; instead, it uses a binding and hiding commitment scheme (Pedersen, 1992), together with Bulletproof range proofs (Bünz et al., 2018). By using commitments as outputs, coins act as having their own private key, the knowledge of which, along with the knowledge of their value, would enable spending the funds. This guarantees anonymity and unlinkability.
4.2.3. Handling Change
Account-based ledgers use transactions to transfer exact value between payer and payee, therefore these kind of systems do not need to support change.
Instead, token-based ledgers propose different approaches for handling change. Systems relying on Chaumian tokens do not easily support change or partial spending (e.g., (Dold, 2019)). Tomescu et al. (Tomescu et al., 2022) allow multiple inputs and outputs for a single transaction, but requires the payer to prove value preservation in ZK. The proof consists in two parts: the first proves that the sum of inputs equals the sum of outputs; the second is a range proof on the output value (needed because ZKPs works with modulo arithmetic and the equality proof is not enough to guarantee value preservation). System based on unspent tokens inherit the change implementation strategy by Bitcoin, where the change is another unspent token owned by a (new) address managed by the payer (e.g., (Miers et al., 2013; Androulaki et al., 2020; Elfadul et al., 2024)) or a commitment controlled by the payer (e.g., (Noether et al., 2016; Wüst et al., 2019)).
4.2.4. Support for Different Denominations
Chaum (Chaum, 1983) considered a simple system where tokens have implicit value. A few extensions consider the possibility to emit tokens with different denominations (e.g., (Dold, 2019; Tomescu et al., 2022)). For example, Dold (Dold, 2019) in GNU Taler use multiple denomination key pairs for blind-signing coins of different monetary values. Conversely, Tomescu et al. (Tomescu et al., 2022) explicitly represent the token value by concealing it within a commitment. As regards unspent token-based ledgers, Zerocoin does not support arbitrary denominations (Miers et al., 2013). This design simplifies the ZKP used for anonymity but introduces usability limitations, as users must split or combine coins when making payments. Zerocash (Ben Sasson et al., 2014) supports arbitrary denominations by using homomorphic commitments, allowing users to transact any amount while preserving privacy. The remaining approaches allow transacting tokens of different denominations, as each unspent token internally holds its value (or a commitment to it). Among them, CryptoNote (Van Saberhagen, 2013) does not hide transaction amounts, making it vulnerable to statistical analysis—e.g., identifying change as the smaller output. Using multiple change addresses can obscure amounts but leads to issues like proliferation of dust transactions (Noether et al., 2016). Hiding values with commitments prevents such analysis.
4.2.5. Token Life-cycle
Double Spending Protection
Two main approaches to double spending protection emerge: deposit-time checks and publicly verifiable ledgers. Centralized systems usually relying on Chaumian tokens do not provide a transparent ledger or the primitives for proactive verification of double spending for the payee. Instead, they check token validity at deposit time, i.e., when the recipient interact with the bank to deposit the token (e.g., (Chaum, 1983; Dold, 2019; Tomescu et al., 2022; Goodell et al., 2023)). The public visibility of the ledger in blockchains (e.g., (Nakamoto, 2008)) allows payee to autonomously check whether an received token has been already spent. This validity check is also possible in privacy preserving systems, which usually require the payer to provide a ZKP of token validity (e.g., (Miers et al., 2013; Poelstra, 2016; Noether et al., 2016; Li et al., 2021; Xue et al., 2023; Yu et al., 2024)).
Transferable Tokens
Only a few works (e.g., (Hayes, 1990; Goodell et al., 2023)) explore mechanisms for transferring Chaumian tokens, i.e., enabling safe ownership changes. Hayes (Hayes, 1990) proposes a system where each coin owner signs a chain of transactions to ensure traceability. For example, if user receives a token from the bank, he can pay another user by issuing a token for and attaching the token received from the bank. When the token chain becomes too long, the owner can exchange it for a fresh token via the bank. However, this approach is vulnerable to double spending, which is mitigated by imposing limited token lifetimes. Similarly, Goodell et al. (Goodell et al., 2023) define tokens that carry a proof of provenance: it verifies validity without requiring external interaction. This proof is a list of token commitments including the initial commitment and signature by the issuer.
Immediate Claiming
With Chaumian tokens, the payer-payee interaction takes place off-the-ledger; therefore, the payee should be online to deposit the token, after a validity check by the token issuer. This kind of interaction requires the payee to immediately claim the token to prevent double spending attacks (Chaum, 1983). One-time signatures can be used to reveal misbehaving senders’ identity, enabling revocation and blacklisting (Hayes, 1990). Similarly, in (Camenisch et al., 2006), spending a token multiple times reveals the sender’s identity. Conversely, Sander and Ta-Shma (Sander and Ta-Shma, 1999) avoids relying on immediate claiming of tokens by using Merkle hash trees to publicly authenticate coin validity through inclusion proofs, rather than secret signatures. Coins remain valid as long as users maintain updated hash chains to public tree roots, which are periodically broadcast. This enables offline, asynchronous payments where coins can be spent without real-time interaction with the bank.
4.3. Privacy Features
We delve into the details of the technical solutions used to obtain the different privacy goals. Privacy in digital currencies is often defined through different sub-properties focusing on anonymity of participants, value confidentiality, plausible deniability, and unlinkability (e.g., (Almashaqbeh and Solomon, 2022; Amarasinghe et al., 2019; Bernal Bernabe et al., 2019; Genkin et al., 2018)).
4.3.1. Sender Anonymity
Sender anonymity relates to the protection of the payer’s identity, i.e., the transaction sender. Some work implement asymmetric anonymity, where only one party remains anonymous (e.g., (Chaum, 1983; Sander and Ta-Shma, 1999; Miers et al., 2013; Dold, 2019; Zhang et al., 2019; BIS Innovation Hub, 2023; Goodell et al., 2023; Yu et al., 2024; Androulaki and Karame, 2014)). The main techniques for achieving sender anonymity are: blind signatures (e.g., (Chaum, 1983; Dold, 2019; Goodell et al., 2023; Liu et al., 2022)) and their variants—one-time (e.g., (Hayes, 1990)), threshold (e.g., (Sarencheh et al., 2025; Kiayias et al., 2022; Androulaki et al., 2020)), and randomizable (e.g., (Tomescu et al., 2022; Wüst et al., 2019))—, ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), group signatures (e.g., (Zhang et al., 2019)), and commitments with ZKPs (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)).
Chaum (Chaum, 1983) recurs to blind signature to unlink the sender request for signing a token and its later spending. To obtain revocable anonymity and reveal the sender on double spending, Hayes (Hayes, 1990) introduces a one-time blind signature schemes. However, since each transaction requires a new signing key, the storage and computation management is cumbersome; moreover, one-time signatures often require large signature size and expensive verification, making them impractical for high-throughput systems. UTT (Tomescu et al., 2022) realizes blind signatures using randomizable signatures by Pointcheval and Sanders (Pointcheval and Sanders, 2016), which allow signatures to be altered using fresh randomness without compromising their validity. All these schemes rely on a central issuer, who can represent a single point of failure. This can be mitigated, e.g., by using threshold blind signatures, where a quorum of authorities must cooperate to create a valid blind signature (e.g., (Androulaki et al., 2020; Kiayias et al., 2022; Sarencheh et al., 2022)).
Other works reuse the idea of mixing protocols through the concept of ring signatures, which conceals the sender’s public key—hence, his identity—in a group of other users’ public keys (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Li et al., 2021; Elfadul et al., 2024)); this group is referred to as anonymity set. For example, CryptoNote (Van Saberhagen, 2013) uses one-time (linkable) ring signatures for sender anonymity. For each transaction, the sender generates a one-time key and derives a key image, then selects an anonymity set and produces a ring signature including the key image and a verification challenge. The key image is stored on a public ledger managed by a centralized entity to detect double-spending. Verifying a signature requires the public keys of all users in the ring; thus, larger anonymity sets increase privacy but also transaction size. Different privacy-preserving cryptocurrency systems use ring signatures for obtaining sender’s anonymity, including Monero101010Möser et al. (Möser et al., 2018) showed weaknesses in Monero’s key sampling: of transactions do not include any other public key, sampling is not really random, and real inputs can be inferred with accuracy. (Noether et al., 2016), Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024). However, studies show that Monero’s anonymity can be compromised through transaction analysis (e.g., (Möser et al., 2018; Kappos et al., 2018; Kumar et al., 2017)).
Considering the requirement of revoking sender’s anonymity if needed, Zhang et al. (Zhang et al., 2019) use group signatures where a trusted centralized entity sets up the anonymity set and can, optionally, uncover the senders’ identity for inspection. Differently from ring signatures where the anonymity set is dynamic and can change for each transaction, group signatures have a less flexible management.
To be valid, unspent tokens require proving their ownership. To implement sender’s anonymity, existing approaches rely on two key tricks. First, the unspent token includes a commitment to its serial number; the serial number is revealed only to mark the token as spent (see Section 4.2.5). Second, the spending operation includes a ZKP that demonstrates knowledge of the opening to a commitment of an unspent coin. Sander and Ta-Shma (Sander and Ta-Shma, 1999) use multiple Merkle trees to track valid Chaumian tokens. To prove validity, the sender provides a NIZK proof showing that he knows a valid coin in the tree and that the coin’s serial number and a hash chain correctly lead to one of the live Merkle tree roots. Zerocoin (Miers et al., 2013), and its extensions (e.g., (Ben Sasson et al., 2014; Garman et al., 2017)), adopts ZKPs as well. In these systems, the payer publishes a transaction that includes the recipient’s address, an empty origin address, a ZKP, and the commitment opening. The ZKP prevents linking the sender to any specific commitment; the commitment itself hides the opening using a random value that is never revealed, preventing any association between the two.
Privacy-preserving account-based models mostly rely on ZKPs to obtain sender’s anonymity (e.g., (Gross et al., 2021; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024)). Specifically, a commitment to the account state is used to trace the balance (along with other information, as detailed later); account state updates include ZKPs to prove correctness. For example, Gross et al. (Gross et al., 2021) use the ideas of Zerocash by committing on the accounts state and by invalidating them on update. The authors suggest using Merkle trees to prove that a transaction proposal refers to an existing commitment in the ledger without pointing to (and thus revealing) it. Platypus (Wüst et al., 2022) uses a model where payer and payee exchange with the central bank commitments to their updated accounts, the transaction amount, a serial number, and two ZKPs proving the update’s validity and its link to a prior certified state. The payer sends his part to the payee, who completes it and interacts with the central bank. If the serial and the ZKPs are valid, the bank updates a publicly available log. Conversely, PEReDi (Kiayias et al., 2022) achieves sender anonymity through anonymous channels, an encrypted ledger, and a threshold signature scheme. Specifically, the sender and receiver each submit partial (specular) transactions that update only their respective account states. These transactions are encrypted using threshold encryption. Sender and receiver must both actively engage in updating their accounts. Solidus (Cecchetti et al., 2017) uses banks as proxies; therefore, transactions are not explicitly exchanged between users but among their respective banks.
4.3.2. Receiver Anonymity
The key techniques for obtaining receiver’s anonymity are one-time destination key or stealth address (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), encrypted address (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)), commitments as output (e.g., (Poelstra, 2016; Wüst et al., 2019)), and updatable public key (e.g., (Fauzi et al., 2019; Papadoulis et al., 2024)). Obtaining receiver anonymity with Chaumian tokens is challenging, as the payee must interact with the bank to deposit the received tokens. Hayes (Hayes, 1990) suggests using one-time aliases to obtain privacy, even though the idea is not detailed for practical usages. A similar idea can be found in CryptoNote (Van Saberhagen, 2013), where the sender generates a one-time destination key by combining a random number with the receiver’s public key—essentially performing a Diffie-Hellman exchange to derive a shared secret. Because this process requires two elliptic curve keys, a CryptoNote address is twice the size of a standard elliptic curve public key. The transaction is then sent to the one-time public key and includes a commitment to the random value. The receiver scans all the incoming transaction using his private key, reconstructs the one-time destination public key from the commitment, and checks if it matches; if it does, he recognizes the payment as intended for him. The sender and receiver never interact directly. The most obvious disadvantage of this approach is that every receiver has to read every transaction to identify those belonging to him. This approach has been later integrated in other privacy-preserving cryptocurrencies, including Monero (Noether et al., 2016), which calls the one-time receiver address as stealth address, Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024) (see Section 3.6.1).
In Zerocash (Ben Sasson et al., 2014) (unspent token-based ledger), the sender includes an encrypted version of the recipient’s address in the transaction (using asymmetric encryption). Therefore, periodically, each user has to scan the ledger, decrypt transactions using his private key, and identify those intended for him. Similarly, UTT (Tomescu et al., 2022) (Chaumian token-based system) ensures receiver’s anonymity using homomorphic commitments, randomizable signatures, and anonymous (Boneh-Franklin) IBE scheme (Boneh and Franklin, 2003). A coin is a commitment to a tuple containing the owner’s identifier, a serial number issued by the central bank, and its value. To be valid, each coin comes with a (randomizable) signature by the central bank. To make a payment, the sender creates a transaction for the payee, which includes the receiver’s identity commitment, the output value commitment, the encrypted coin details (using the recipient’s key), and ZKPs for correctness. The transaction is sent to the bank for validation. If valid, the bank blindly issues new coins for the recipient and records them on a public ledger—without learning identities or amounts, as they remain hidden via homomorphic commitments111111The homomorphic properties of the commitment scheme allows to directly combine the value commitment, identity commitment, and serial number commitment in a single coin commitment.. The recipient periodically scans the ledger for transactions encrypted for him, decrypts payments with his private key, verifies the amounts, and retrieves the coins’ details. Then, he can randomize the coins before adding them to his wallet.
The unspent token-based ledger MimbleWimble (Poelstra, 2016) (and PRCash (Wüst et al., 2019)) performs transactions between commitments rather than addresses; this guarantees receiver’s anonymity and unlinkability.
Account-based approaches typically achieve receiver anonymity using the same technique as for sender anonymity (e.g., (Wüst et al., 2022; Kiayias et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Lin et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Sarencheh et al., 2025; Papadoulis et al., 2024; Fauzi et al., 2019)). This is done by updating account state commitments and employing ZKPs to verify correctness without revealing the receiver’s identity. We mention Quisquis (Fauzi et al., 2019), where an account is a pair of public key and a homomorphic commitment to its balance. A transaction updates both the public key and the commitment using updatable public key (see Section 3.6.2). Senders select public keys—including the recipient’s and decoys—to form the transaction input, updating only the sender’s and recipient’s balances while leaving others unchanged. Since each address appears at most twice on the ledger, de-anonymization is harder. However, this approach requires users to periodically scan the ledger to update their keys, thus introducing some transaction serialization.
4.3.3. Transaction Value Confidentiality
Two main approaches emerge from the literature to conceal transaction details: commitments (e.g., (Ben Sasson et al., 2014; Androulaki and Karame, 2014; Noether et al., 2016; Poelstra, 2016; Garman et al., 2017; Fauzi et al., 2019; Wüst et al., 2019; Gross et al., 2021; Guan et al., 2022; Wüst et al., 2022; Beer et al., 2024; Elfadul et al., 2024; Tomescu et al., 2022; Androulaki et al., 2020; Narula et al., 2018)) and public key encryption (e.g., (Kiayias et al., 2022; Sarencheh et al., 2025; Bünz et al., 2020; Ma et al., 2021)). All the approaches using commitments and most of those adopting public key encryption (e.g., (Bünz et al., 2020; Ma et al., 2021)) use ZKPs to prove correctness.
Commitments
Camenisch et al. (Camenisch et al., 2006) use commitments with verifiable encryption to hide the commitment’s opening: the opening is a random value generated and encrypted during withdrawal, and decrypted and stored at deposit time to prevent double spending. In 2013, Maxwell proposed confidential transactions121212https://elementsproject.org/features/confidential-transactions/investigation, which use Pedersen commitments to hide transaction amounts while ensuring correctness (also using range proofs). Monero adopted and extended this technique by combining ring signatures for sender’s anonymity, obtaining RingCT (Noether et al., 2016). Zerocash (Ben Sasson et al., 2014) uses ZKPs to achieve full transaction privacy: a cryptographic commitment is made to a new coin, including its value, owner, and serial number; then, a ZKP ensures transaction validity without exposing sensitive data. Commitment-based schemes are the most widely adopted to obtain confidential transaction value, especially because their homomorphic versions allow to combine commitments without leaking sensitive data. To prevent users from creating negative or infinite amounts, range proofs prove that a committed value is greater than zero and within a valid range.
Public Key Encryption
In PEReDi (Kiayias et al., 2022), senders and receivers update their accounts by submitting encrypted transaction information to several maintainers who operate a distributed ledger. Transactions are encrypted using threshold ElGamal encryption (Elgamal, 1985), which requires a quorum of maintainer to cooperate to revoke confidentiality and decrypt transaction details. Authorities can trace both transactions and identities. PARScoin (Sarencheh et al., 2025) extends this approach to introduce an offline payment protocol that does not require interaction between sender and receiver. Androulaki et al. (Androulaki et al., 2020) extend Zerocoin to implement auditability: each user is assigned to an auditor at registration time, who can decrypt and inspect the user’s transactions when required. Using encryption instead of commitments allows to revoke anonymity and implement auditability and tracing. In Solidus (Cecchetti et al., 2017), Publicly Verifiable Oblivious RAMs (PVORMs) enable banks to update the accounts of their clients without revealing exactly which accounts are being updated. PVORM provides a ZKP demonstrating that the updates are correct with respect to the transaction triggering them.
A few solutions implement privacy-preserving currencies using smart contracts on public blockchains. To ensure confidentiality, transactions are encrypted with the sender’s public key and supplemented with a NIZK proof for public verifiability (e.g., (Ma et al., 2021; Bünz et al., 2020)). For example, Zether (Bünz et al., 2020) uses smart contracts to define a privacy-preserving tokens that can be exchanged between public keys associated with a confidential balance. They use ElGamal encryption, which has homomorphic properties, for hiding each account’s balance. Zether introduces the -Bullets (an enhancement of the Bulletproofs range ZKP) to efficiently prove statements over the encrypted transfer value and the new sender balance. However, Zether currently allows only one anonymous transfer per epoch (to prevent double spending), and execution costs (namely, gas costs on Ethereum) are very high (Almashaqbeh and Solomon, 2022). Similarly, DSC (Ma et al., 2021) optimizes privacy within Ethereum-style smart contract environments, but does not support anonymity: it uses ZKPs and homomorphic encryption to hide balances and transferred amounts. User balances are stored on the ledger in encrypted form. We briefly mention few other works that focus on privacy in smart contracts. For example, Hawk (Saleem et al., 2024) uses MPC to split a contract into private and public parts, allowing confidential state. Enigma (Zyskind et al., 2015) uses secure enclaves to run contract code on encrypted inputs. These solutions are beyond currency transfers, but they show that account privacy often requires splitting roles or off-chain computation. Recent research projects aim for private smart contracts via ZKP frameworks, often assuming a model with private channels or truster parties (e.g., (Bowe et al., 2020)).
4.3.4. Network Anonymity
Network anonymity concerns the protection of sender and receiver IP addresses, which could otherwise reveal real-world identities. Most of the existing approaches do not investigate (or even mention) the network anonymity feature. Few works mention it but consider this issue as out of scope of their investigation (e.g., (Cecchetti et al., 2017; Baudet et al., 2023; Tomescu et al., 2022; Wüst et al., 2022)). Monero (Noether et al., 2016) and Traceable Monero (Li et al., 2021) explicitly support network anonymity by using TOR (The Onion Router131313https://www.torproject.org/) and I2P (Invisible Internet Project141414https://geti2p.net/en/) to hide the real IP address (Shi et al., 2024). While TOR routes traffic through multiple relays to obscure the user’s IP address, I2P is a network optimized for anonymous peer-to-peer communication, using fully decentralized routing and encryption to enable different anonymous services. Moreover, Monero uses Dandelion++ (Fanti et al., 2018), an improvement of Dandelion (Bojja Venkatakrishnan et al., 2017), which prevents malicious attackers from linking transactions with their source IPs. Dandelion++ uses randomization mechanisms for message forwarding, which happens in two phases: stem (anonymization) and fluff (broadcast). In the stem phase, a transaction is first relayed privately across a small, randomly chosen subset of nodes; each node probabilistically decides whether to continue forwarding the transaction along the stem or to switch to the second phase. In the fluff phase, the node broadcasts to the entire network using traditional flood propagation protocol.
4.3.5. Linkability and Unlinkability
Unlinkability ensures that transactions cannot be correlated or traced back to the same user. For instance, it should be impossible to conclude that two transactions were delivered to or sent from the same user (other definitions can be found in (Chaum, 1983; Amarasinghe et al., 2019; Zhang, 2023)). Unlinkability can refer to the sender, receiver, or both. A few works break this property into two facets, internal and external. For instance, CoinJoin (Maxwell, 2013) does not preserve internal unlinkability as each participant in the group that mixes transactions can link entities within the group.
Obtaining unlinkability in public ledgers is challenging as linking between addresses in cryptocurrency systems stems from different patterns (e.g., (Ron and Shamir, 2013; Chan and Olmsted, 2017; Béres et al., 2021; Victor, 2020; Zhong and Mueen, 2024; Lu et al., 2022; Ermilov et al., 2017; Kalodner et al., 2020)). A non exhaustive list of patterns and heuristics to link addresses is the following.
Common input ownership heuristic: If multiple addresses appear as inputs in a single transaction, they are likely controlled by the same entity;
Change address: In a transaction with two outputs, one is the payment, the other one may be the change returned to the sender;
Address reuse: Using the same address for multiple transactions make it easy to link payments to the same user151515This works because address are pseudonymous but not private.;
Temporal analysis: If transactions occur in rapid succession or at predictable intervals, they may be linked to the same entity;
Clustering analysis: It deals with identifying wallet behavior such as round numbers, specific fee patterns, or common spending habits;
Dust attack analysis: Small amounts of cryptocurrency are sent to many addresses to track their movements (similar to a phishing attack); if dust is later spent with other transactions, it links addresses together.
Existing countermeasures primarily aim to obscure input-output relations within transactions through mixing protocols and address reuse prevention. Mixing protocols disrupt common ownership heuristics and help mitigate change address attacks. Address-changing techniques generate new sender or receiver addresses for each transaction to prevent linkability. Additionally, ensuring network anonymity helps further obscure user identities. To counter dust attacks, the simplest defense is to ignore dust transactions, preventing adversaries from establishing links between addresses. We review mixing protocols and approaches adopted in privacy-native currencies.
Mixing Protocols
Different mixing protocols have been proposed, and we summarize them in Appendix A, Table LABEL:tab:sota-mixing. The naïve approach for mixing transactions consists in a centralized service that collects transactions, and publishes on the ledger a single large transaction with multiple inputs and outputs. This hides the information regarding which inputs flows towards which output, thus limiting (if not preventing) linkability between sender and receiver. The Mixcoin centralized protocol considers an accountable mixer, which offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Blindcoin (Valenta and Rowan, 2015) extends this protocol by considering blind signatures for the warrants; this hides the link between input and output addresses of a user even from the mixer, achieving full unlinkability; however, the mixer can still thief coins. Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. In BlindHub (Qin et al., 2023), mixing is performed within payment channel hubs, where the hub acts as tumbler and uses blind signatures to hide the customers’ addresses from itself. Obscuro (Tran et al., 2018) is a mixing protocol that utilizes Trusted Execution Environments (TEEs) and enables participants to verify the mixer’s integrity via remote attestation. It is specifically designed to defend against two key threats: the participation rejection attack, where malicious users try to manipulating mixing by withdrawing their deposits before the TEE reads them, and the blockchain forking attack, where adversaries present stale blocks to mislead the mixer. Numerous papers also investigate decentralized protocols. CoinJoin (Maxwell, 2013) is a technique that allows multiple users to combine their transactions into a single joint transaction; the author does not propose a fully fledged protocol, but only describes the mixing idea. CoinShuffle (Ruffing et al., 2014) allow to combine multiple users’ inputs and outputs in a single transaction but may compromise anonymity if transaction values differ or the anonymity set is small. Later, Ruffing et al. (Ruffing et al., 2016) proposed CoinShuffle++, which uses a more efficient mechanism for anonymity, i.e., Dining Cryptographers Networks (DC-nets) instead of mix-nets: while a mix-net requires sequential processing, hence a number of communication rounds linear in the number of participants, DC-nets enable to process mixing in parallel, hence it requires only a constant number of communication rounds. Xim (Bissias et al., 2014) is a multi-round protocol for anonymously finding mix partners based on advertisements placed on the blockchain, with the property that no outside party can identify participants that pair up. SecureCoin (Ibrahim, 2017) creates an aggregated temporary deposit address and requires the collaboration of involved peers to generate novel destination addresses and collectively redistribute coins. Leveraging Confidential Transactions (which nonetheless is available only in a soft fork of Bitcoin), ValueShuffle (Ruffing and Moreno-Sanchez, 2017) provides full privacy besides parties anonymity. Confidentiality enables mixing funds of different values without compromising unlinkability. CoinParty (Ziegeldorf et al., 2018) is a decentralized mixing protocol that distributes the mixing process across a group of semi-trusted parties. They use an MPC protocol to jointly perform sensitive operations (i.e., encryption, decryption, and address shuffling), ensuring that no single party can link input and output addresses. Threshold signatures is used to enable claiming funds upon agreement of a majority of mixing nodes. It also includes mechanisms to detect and exclude misbehaving participants, increasing the overall protocol reliability.
Decentralized tumblers have also been designed using smart contracts on Ethereum. Möbius (Meiklejohn and Mercer, 2018) uses ring signatures and stealth addresses to obfuscate the senders and recipients associations (as Monero does (Noether et al., 2016)). First, senders furnish both their funds and stealth keys to the smart contract; then, each recipient creates a ring signature to withdraw their funds from the contract and transfer them to an ephemeral address. The size of the anonymity set is limited to the size of the ring, and the gas cost of the withdrawing transaction increases linearly with the size of the ring. MixEth (Seres et al., 2019) aims to minimize gas costs by utilizing Neff’s verifiable shuffle (Neff, 2001) and off-chain messages. Once the sender deposits coins, the MixEth contract enters rounds of shuffling and challenging, during which anyone can challenge the correctness of the preceding shuffle. When recipients believe that sufficient shuffling has been performed, they can withdraw their funds from the MixEth contract. We also mention AMR (Le and Gervais, 2021) which disrupts the traceability connection between coins deposited and withdrawn by the same user. Participants deposit a predetermined quantity of coins into a smart contract, which establishes a Merkle tree structure over the deposits. To withdraw coins, a zk-SNARK prove coins inclusion in the Merkle tree. AMR leverages lending platforms to provide users with the opportunity to accrue interest on their deposited assets.
Although improving unlinkability, mixing protocols are still vulnerable to transaction graph analysis to some extent. For example, if the mixing process does not involve adequate randomization or if there are patterns in the amounts or timing of transactions, likely connections can be identified, even after several rounds of mixing. Also, if a participant does not properly anonymize the change, it can be traced back to the sender. To maximize privacy, participants need to combine mixing protocols with other privacy-preserving techniques and follow best practices (e.g., distributing coins across multiple addresses) to further obscure their transaction history and reduce the risk of being traced through transaction graph analysis. Further details on mixing solutions for Bitcoin and Ethereum, e.g., in (Arbabi et al., 2023) and on transaction graph analysis, e.g., in (Ron and Shamir, 2013; Fleder et al., 2015; Chan and Olmsted, 2017).
Integrated Solutions
Privacy-preserving cryptocurrencies leverage a variety of strategies to improve unlinkability, focusing primarily on address-changing, transaction mixing, and use of ZKPs. Each of these techniques adds a layer of privacy, with trade-offs in terms of anonymity, computational complexity, and vulnerability to analysis. One-time destination addresses, stealth addresses (e.g., (Van Saberhagen, 2013; Noether et al., 2016)), or updatable public keys (e.g., (Papadoulis et al., 2024; Fauzi et al., 2019)) help obtain recipient unlinkability. While this ensures the recipient’s privacy, it does not address the issue of transaction amounts, which are still visible on the ledger. Moreover, existing protocols leveraging updatable public keys require the receiver to scan the ledger for incoming payments, which adds complexity for the user. Ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016)) build upon the idea of mixing, which redistributes funds in a manner that obscures the origins (and destinations) of the coins. With ring signatures, a sender creates a ring with multiple decoy inputs, making it impossible to determine which input belongs to the spender. However, the privacy protection is limited by the number of decoys in the ring, and sophisticated heuristics can sometimes reduce the anonymity set, thus compromising privacy. For the most robust privacy, ZKPs offer a powerful solution (e.g., (Ben Sasson et al., 2014; Poelstra, 2016)). They allow transactions to prove their validity without revealing any details about the sender, receiver, or transaction amount. With this approach, transactions are recorded as cryptographic proofs (instead of sender-receiver addresses), ensuring that outputs cannot be linked to inputs. Among ZK-based approaches, we mention the commitments and nullifiers strategy (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)): a coin contains a commitment to its serial number, and spending the coin involves publishing a nullifier, i.e., a pseudo-random value derived from the coin serial number. This ensures unlinkability to the original coin, while marking it as spent. However, designing a system with commitments and nullifiers is more complex than ring signatures or stealth addresses, and is also more computational demanding (see Section 4.5).
4.3.6. Plausible Deniability
With plausible deniability, we intend whether a user can deny intentional participation in a transaction, preventing others and adversaries from proving their involvement (e.g., (Almashaqbeh and Solomon, 2022; Ziegeldorf et al., 2018)). Existing work is roughly evenly divided between those that support this property (e.g., (Noether et al., 2016; Poelstra, 2016; Narula et al., 2018; Kiayias et al., 2022; Tomescu et al., 2022; Wüst et al., 2022; Goodell et al., 2023; Papadoulis et al., 2024)) and those that do not (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Garman et al., 2017; Bünz et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Yu et al., 2024)). Basically, protocols “on-demand”, such as mixing protocols, subsume an explicit user intention in using the additional service to obtain a privacy feature (such as unlinkability). Likewise, overlay systems implementing an anonymous cryptocurrency, such as Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), EZC (Androulaki and Karame, 2014), Zether (Bünz et al., 2020), require explicit intention to convert native currencies to a privacy-preserving one. Conversely, systems that force anonymity or unlinkability as default, readily provide the property of plausible deniability. Among these systems, we mention Monero (Noether et al., 2016), MimbleWimble (Poelstra, 2016), UTT (Tomescu et al., 2022), PEReDi (Kiayias et al., 2022), or Platypus (Wüst et al., 2022), that do not provide the functionality of combining anonymous and public transactions.
4.4. Auditability
Auditability refers to the ability to examine, track, and verify transactions or balances while preserving a certain degree of privacy. It is a key feature of the third generation of privacy-preserving cryptocurrencies and, in particular, CBDCs. The concept of auditability is closely related to that of policy, which refers to the rules enforced by the protocol or by users to govern behaviors (e.g., to restrict who can transact and under what conditions). While policies involve ex-ante enforcement (i.e., preventing undesired actions before they occur), auditability supports post-facto verification by auditors and regulators. We point out that accountability is also important, as it implies linking actions to individuals, e.g., through identity-binding mechanisms such as those provided by a registration authority to comply with AML/CFT regulations.
Full privacy often conflicts with compliance: if transactions are completely opaque, regulators have no oversight (Ballaschk and Paulick, 2021). Some privacy designs therefore incorporate policy-controlled privacy. We identify the following approaches that enable or support auditability: anonymity budget, operating limits, revocable anonymity and user tracing, coin tracing, and verification without disclosure.
4.4.1. Anonymity Budget
An anonymity budget limits the participant’s ability to perform anonymous payments. Each transaction consumes a portion of this budget, and once the participant runs out of it, he loses the ability to make further anonymous payments (e.g., (Camenisch et al., 2006; Wüst et al., 2019; Tomescu et al., 2022; ECB, 2019)). Camenisch et al. (Camenisch et al., 2006) introduce the bounded-anonymity business model, where there is a publicly-known limit to the number of coins a user may anonymously transfer to a specific merchant. PRCash (Wüst et al., 2019) restricts the total amount of anonymous payments a user can receive within a time window (or epoch). The anonymity budget is a special non-spendable output. By adjusting the budget size and duration, authorities can regulate the flow of anonymous funds. Once the budget is exhausted, or if the user opts for non-anonymous transactions, the payee creates an output that includes his identity encrypted with the regulator’s public key. UTT (Tomescu et al., 2022) employs the anonymity budget to balance privacy with accountability for spending operations. Beyond the allocated budget, transactions may either lose privacy or require auditing. An auditor periodically refreshes anonymity budgets and can approve transactions that exceed the budget, e.g., by requesting additional transaction details before signing. In all these systems, the anonymity budget is implemented as a special token that must be included in any anonymous but accountable transaction.
4.4.2. Operating Limits
Operating limits indicate constraints beyond which transactions cannot be carried out. Limits are usually imposed on the single transaction value, overall user’s balance, amount of sent or received funds, and number of performed transactions within a time window. Compliance with limits is generally proven in zero-knowledge. Garman et al. (Garman et al., 2017) extend Zerocash by introducing several building blocks to support auditability, including regulatory closure (i.e., restricting transactions to assets of the same type), spending limits, selective user inquiries, and coin tracing. A central idea is the enforcement of a per-transaction spending limit, whereby a transaction becomes invalid if it exceeds the defined limit, unless it is signed by an authorized party. Another proposal involves maintaining a per-user counter incremented with each outgoing transaction. The system can then reject transactions if the user’s counter is above a predefined threshold. Building on these ideas, Gross et al. (Gross et al., 2021) propose a two-layered account-based system for implementing a CBDC. To enforce the spending policy, each account maintains an epoch turnover, representing the total amount spent during the current epoch. Compliance with spending limits is proven using ZKPs as part of each account update transaction. Similarly, AQQUA (Papadoulis et al., 2024) uses multiple counters per account, which are publicly stored as three commitments: one to the account balance, and two to the cumulative amounts sent and received. AQQUA also introduces a registration authority, which links users’ real-world identities to their (possibly, multiple) public keys, and an audit authority, which verifies compliance with defined policies. During an audit, the authority selects a user by his public key and requests the opening of two historical account state snapshots (for every user’s accounts). These openings allow the authority to directly inspect whether the user complies with specific policies. Five types of policies exist: spending limits, receiving limits, transaction value limits, non-participation (i.e., inactivity over a time period), and transaction opening (i.e., selective disclosure of the value sent or received in a specific transaction). After the audit, the user randomizes his accounts (and provides a ZKP of valid update), thereby restoring his privacy. Xue et al. (Xue et al., 2023) use per-user counters, as well. They design an unspent token-based ledger, which stores counters in a Merkle tree managed by regulators, which are also in charge of user on-boarding. Users can define policies using ZKPs161616Specifically, zk-SNARKs are used for arithmetic statements and Sigma Protocols for algebraic statements. on counters, which are updated for each transaction through user-regulator interaction. Counters enable policies on total amount and frequency of fund transfers within a time period. When suspicious transactions are detected, the participants’ identities can be recovered by regulators. Similarly, Chen et al. (Chen et al., 2020) consider three policies—on transaction limits, rates, and value—and recur to ZKPs to attest compliance.
4.4.3. Revocable Anonymity and User Tracing
User tracing mechanisms enable authorities to discover the user’s identity or to identify all transactions associated with a specific user when needed. This concept is often related to revocable or conditional anonymity. Revocable anonymity refers to the privacy model where users remain anonymous by default, but their identity can be revealed under certain conditions or rules. Conversely, user tracing refers to technical mechanisms that allow authorized entities to identify users or link their transactions when required.
Some contributions use special digital signatures that reveal the sender’s identity if he double spends (e.g., (Hayes, 1990; Zhang et al., 2019; Beer et al., 2024)). Other approaches encrypt user-related information—such as identifiers, keys, or transaction metadata—within the transaction, in special tags, making them accessible only to tracing authorities with decryption capabilities (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Zhang et al., 2019; Liu et al., 2022; Lin et al., 2020; Li et al., 2021; Kiayias et al., 2022)). This is often referred to as the trapdoor approach, where decryption is enabled through a private key or other privileged information. Many systems also rely on registration authorities that link real-world identities to cryptographic keys, making traced identities resolvable (e.g., (Zhang et al., 2019; Liu et al., 2022)). For example, Cecchetti et al. (Cecchetti et al., 2017) and Liu et al. (Liu et al., 2022) consider a setting with a central bank. While Cecchetti et al. (Cecchetti et al., 2017) manage in oblivious RAM the transactions across users’ commercial banks, Liu et al. (Liu et al., 2022) require that each transaction includes the user’s identity encrypted with the central bank public key. In a more general setting, Zhang et al. (Zhang et al., 2019) introduce a supervision authority that can extract the payer’s identity from any transaction using its private key (and the registration authority). Garman et al. (Garman et al., 2017) propose a tracing mechanism where each user receives a unique encryption key: if the user is being traced, the key is a randomized version of the authority’s key; otherwise, it is derived from a null key. Users cannot tell which type of key they received. To enhance transparency, periodically the system spends tokens that publicly demonstrate how many users were traced without revealing their identities. Androulaki et al. (Androulaki et al., 2020) assign each user to an auditor at registration time; then, the system encrypts the transaction information under the public keys of the sender’s and the receivers’ auditors. Xue et al. (Xue et al., 2023) encode the payer’s identity in each transaction using the regulator’s key and the transaction’s epoch identifier. The epoch identifier enables temporal scoping, which limits the impact of key compromising and can enable granular tracing.
A few works assign long-term keys to users, from which they derive one-time keys for individual transactions. User tracing mechanisms aim to recover the long-term keys that uniquely identify a user. For example, Lin et al. (Lin et al., 2020) propose a system in which users perform transactions using anonymous keys derived from their long-term public addresses. To enable tracing, a designated manager can use his private key and the anonymous transaction key to recover the user’s long-term public address (and his identity). Traceable Monero (Li et al., 2021) introduces a method for generating one-time key pairs for payees, where each one-time address is associated with a tag. This tag can be decrypted by the tracing authority to recover the payee’s long-term public key. The tag is designed to support both backward and forward tracing: it can identify both the accounts that sent funds to the current account and the subsequent accounts that receive funds originating from it.
The ability of tracing authorities to decrypt transactions places a high trust burden on a centralized entity. To mitigate this, Yu et al. (Yu et al., 2024) require collaboration between two distinct authorities to detect correlations among a batch of transactions—specifically, those originating from the same user—and to trace user identities when necessary. To enable this functionality, the authors introduce the concept of a mark171717Although the concepts of tag and mark are closely related, we follow the terminology of (Yu et al., 2024), using mark to denote data that enables a two-party computation protocol in a privacy-preserving manner., which embeds the sender’s public key encrypted with random values. These values can be reconstructed only through the joint effort of both authorities. Conversely, PEReDi (Kiayias et al., 2022; Sarencheh et al., 2022) proposes distributing the decryption power across multiple parties, requiring them to cooperate and reach a quorum in order to access private information. This very idea is then also used by Sarencheh et al. (Sarencheh et al., 2025).
4.4.4. Coin Tracing
Garman et al. (Garman et al., 2017) propose an approach to trace certain coins, avoiding an extensive transaction auditing. In particular, the scheme enables tracing by marking specific coins as traceable, i.e., tainted coins181818A tainted coin is a traceable digital asset, which could be linked to illicit or suspicious activity, often flagged for enhanced scrutiny or restricted use.. When such a coin is spent, all resulting output coins also become traceable. To implement this, each coin contains a fresh encryption key, and tracing data for the outputs is encrypted under the input coin’s key and under the authority’s public key to prevent the sender from performing unauthorized tracing. To limit cascade tracing and collusion risk (i.e., when a user and the authority jointly trace non-tainted coins), exchanges are allowed to verifiably remove tracing using a public dummy key. Keller et al. (Keller et al., 2021) propose a method to de-anonymize users in CoinJoin and ring signature schemes. They model a scenario where a law enforcement agency collaborates with certain users—referred to as witnesses—who can confirm that their inputs were not used in a given mixing transaction. As the number of such witnesses increases, the anonymity set of the remaining users is reduced, making de-anonymization more feasible.
4.4.5. Verification without Disclosure
Auditability mechanisms that enable verification without disclosure allow a party to confirm compliance with a policy without learning any private data. These mechanisms rely on cryptographic proofs (usually NIZK) to show that hidden data satisfies specific constraints (e.g., (Narula et al., 2018; Wüst et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Ma et al., 2021)). Crucially, the verifier learns nothing beyond the validity of the claimed statement. Besides the approaches we already mentioned in Operating limits, here we review those supporting policies beyond basic balance or transfer checks. For example, zkLedger (Narula et al., 2018) uses Pedersen commitments enabling aggregate computations. zkLedger supports complex auditing via a map-reduce model: banks generate commitments over filtered and transformed data (e.g., checking for non-zero values or squaring for variance) along with NIZKs proofs of correctness, enabling auditors to compute metrics like means and variances. Platypus (Wüst et al., 2022) defines account states with opaque auxiliary data to support custom policy checks. Solidus (Cecchetti et al., 2017) uses PVORM and ZKPs to publicly verify a bank’s transaction log compliance. Zether (Bünz et al., 2020) implements confidential transactions in Ethereum, using NIZKs to verify properties like non-negative balances. DSC (Ma et al., 2021) enforces compliance of encrypted transactions with range and authorization proofs.
4.5. Summary and Takeaways
The analysis presented in this section highlights the diverse and evolving landscape of privacy goals in digital currencies. While early systems primarily targeted anonymity and unlinkability through address obfuscation and mixing techniques, more recent designs emphasize stronger notions of confidentiality and auditability. These are supported by advanced cryptographic tools such as ZKPs, homomorphic commitments, and threshold encryption.
Each privacy goal—anonymity, confidentiality, unlinkability, and auditability—requires distinct mechanisms and introduces specific trade-offs. For instance, unlinkability may conflict with auditability, and encrypted transactions that ensure confidentiality can complicate regulatory oversight. No single protocol fully satisfies all privacy and compliance objectives; rather, systems tend to prioritize different goals based on their threat models, trust assumptions, and intended deployment contexts. Table 2 provides a high-level summary of how these goals are addressed in practice. For each privacy property, we list relevant cryptographic primitives, system mechanisms, and example of representative protocols. This mapping serves as a reference point for researchers and practitioners navigating the trade-offs inherent in privacy-preserving digital currency design.
| Privacy Goal | Key Cryptographic Tools | System Mechanisms | Example Protocols |
| Sender Anonymity | Blind, ring, group, and randomizable signatures, commitments and ZKPs | One-time addresses, anonymity set, anonymous addresses | eCash, UTT, Zerocoin |
| Receiver Anonymity | One-time public keys, encrypted addresses, commitments | One-time addresses | MimbleWimble, Quisquis, Zerocash |
| Confidentiality | Commitments, encryption, ZKPs | Confidential values and balances | Zether, PEReDi, Monero |
| Network Anonymity | TOR, I2P, Dandelion++ | Obfuscate IP addresses and routing paths | Monero |
| Unlinkability | Mixing protocols, ring signatures, one-time addresses, commitments | Fresh key generation, decoy input, output set hiding | Platypus, BlockMaze, CoinShuffle++ |
| Auditability | Authenticated tokens, threshold encryption, anonymous keys, trapdoor, ZKPs | Anonymity budget, operating limits, tracing hooks, revocable anonymity | PEReDi, Solidus, zkLedger |
In general, achieving comprehensive privacy requires combining multiple techniques, each with specific trade-offs in terms of confidentiality, unlinkability, and computational demand. For example, blind signatures are lightweight and provide cash-like anonymity but lack unlinkability without centralized coordination. Ring signatures offer anonymity that scales with ring size, but increasing the anonymity set directly impacts computational cost and latency. Encryption, threshold encryption, and commitments facilitate selective auditability but do not provide sender anonymity on their own. ZKPs offer robust privacy guarantees, but they are computationally intensive, may require substantial RAM (e.g., (Almashaqbeh and Solomon, 2022; Liang et al., 2025)), and require careful engineering for practical deployment. One of the well-known ZKP-based systems is zk-SNARK, which makes transactions compact for verification but needs a trusted setup that, if mishandled, could compromise the security of the system. Importantly, as shown in Section 4.4, implementing auditability often requires careful design of ledger data structures and interaction protocols to avoid unintended information leaks while still enabling policy verification. This multidimensional design space underscores the need for a holistic, design-oriented approach: effective privacy-preserving digital currencies must integrate layered technical solutions that balance user privacy with institutional accountability.
5. Open Challenges and Research Directions
Despite significant advances in privacy-preserving digital currencies, numerous open challenges remain. These challenges stem from the unique intersection of cryptographic capabilities, monetary policy requirements, and societal expectations of privacy and auditability. We highlight a few directions that merit further investigation.
5.1. Formal Definitions of Privacy and Auditability
Existing systems use ad-hoc definitions for privacy and auditability. As pointed out in different previous works (e.g., (Genkin et al., 2018; Almashaqbeh and Solomon, 2022)), there exists several theoretical gaps related to the definition of unified privacy models and auditability. Most protocols informally claim anonymity, unlinkability, or confidentiality; however, formal and composable definitions that encompass transaction graph obfuscation, timing metadata, and value-hiding under active adversaries are lacking. We also point out that protocols combining multiple primitives often lack formal, compositional security proofs in frameworks like UC (Canetti, 2000) or IITM (Kuesters et al., 2013). Auditability mechanisms (e.g., view keys, trapdoor disclosures) are rarely formalized. There is a lack of frameworks to reason about selective disclosure, accountability guarantees, or conditional anonymity under adversarial inspectors.
5.2. Scalability and Efficiency
Many cryptographic protocols and, more broadly, PETs remain computationally demanding. This includes fully homomorphic encryption, ZKPs, and MPC. However, real-world cryptocurrencies and CBDCs must support high transaction throughput with low latency and high reliability. To bridge this gap, future research must focus on developing optimized, lightweight, and parallelizable privacy-preserving protocols that can operate efficiently in high-throughput environments. Promising directions include hardware acceleration (e.g., via TEEs) and the exploration of protocol variants that balance privacy guarantees with resource efficiency. ZKPs play a central role in many of the systems analyzed, yet several fundamental challenges remain. General-purpose ZKP schemes (e.g., zk-SNARKs, Bulletproofs) face issues related to trusted setup, proof aggregation, and post-quantum security. For instance, zk-SNARKs typically require a trusted setup, introducing centralization risks and limiting composability. While transparent SNARKs (e.g., Plonky2191919https://github.com/0xPolygonZero/plonky2 and Spartan (Setty, 2019)) and STARKs mitigate these concerns, they often incur higher computational costs. Achieving efficient, setup-free ZKPs without compromising performance remains an open research problem. Additionally, support for efficiently generating proofs that can verify other proofs—known as recursive proof composition—is still limited, even though this capability is crucial for scalable applications like off-chain rollups. Developing ZKPs that are not only succinct and recursive, but also universally verifiable (i.e., verifiable by anyone without special access), remains a technically challenging and active area of research. Finally, no current ZKP scheme offers both strong post-quantum guarantees202020ZKPs based on bilinear pairings (e.g., Plonk, Groth) or Pedersen vector commitments (i.e., Bulletproof) assume at least the discrete logarithm problem, hence are subject to quantum threats. Other SNARKs (e.g., Ligero, Orion) and STARKs using Merkle tree and other protocols to build a polynomial commitment are post quantum resistant. and performance comparable to elliptic curve-based constructions, highlighting the need for further advances in quantum-resistant proof systems.
5.3. Balancing Privacy with Regulatory Compliance
Total anonymity is generally incompatible with regulatory oversight, making auditable privacy and selective disclosure critical design goals. An open challenge is how to enable systems where honest users maintain strong privacy guarantees, while still allowing authorities to investigate illicit activities under well-defined conditions. Several preliminary approaches explore this trade-off, yet significant work remains to formalize disclosure policies, define the types of information subject to auditing, and clarify the roles and interactions of participating entities. Relevant research directions include ZKPs with selective opening, trapdoor, witness encryption, and anonymity budgeting—each enabling varying levels of conditional transparency. For instance, while fully underexplored, witness encryption holds promise for enabling advanced privacy-preserving policies, such as verifiable time locks or conditional disclosures without trusted third parties (see Appendix A.3.5). Additionally, emerging requirements such as time-bound anonymity, jurisdiction-aware confidentiality, and programmable disclosure call for more flexible and context-aware privacy controls. Designing PETs that can support such adaptive and policy-compliant behaviors, while remaining robust against misuse, poses a complex challenge involving technical, legal, and usability considerations.
5.4. Adaptive Privacy and Role Rotation
Privacy-preserving systems fundamentally rely on the concept of anonymity sets. However, managing these sets in a secure, scalable, and adaptive manner remains an open challenge, especially in decentralized or evolving environments. Many existing systems, such as those based on ring signatures, operate with fixed-size anonymity sets or employ heuristic-based decoy selection. These approaches struggle to maintain robustness as usage patterns shift, and they are vulnerable to poisoning attacks, where adversaries inject tainted or known outputs into the anonymity set to degrade privacy. Developing cryptographically sound methods for dynamic anonymity set construction that grows or shrinks with system usage, while resisting active attacks, remains an unsolved problem. Beyond the anonymity set, adaptive privacy must also account for the dynamic nature of system participants, particularly in regulatory or issuance roles. In systems like CBDCs or regulated cryptocurrencies, authorities responsible for auditing, compliance verification, or even token issuance may change over time due to policy shifts, jurisdictional realignments, or organizational restructuring. Supporting the evolution of designated entities, such as auditors or currency issuers, without compromising user anonymity or transaction confidentiality is a significant technical challenge. Current systems typically assume static trust anchors or fixed multi-signature sets. However, introducing cryptographic mechanisms to dynamically reconfigure these roles (e.g., via threshold cryptography, distributed key generation, or policy-based encryption) requires careful balance between adaptability and privacy. Specifically, open issues relate to securely update the set of auditing entities, policies evolving over time, or support role rotation or revocation while preserving unlinkability, correctness, and forward secrecy.
5.5. Network-Level Anonymity and Cross-layer Privacy Leakage
Network-layer anonymity in cryptocurrencies is typically outsourced to external systems (e.g., Tor, mix networks). However, native integration of network-level privacy mechanisms into cryptocurrency protocols remains largely unresolved. This delegation creates a critical cross-layer vulnerability, where even perfectly private on-chain transactions can be deanonymized through timing analysis, transaction ordering, volume inference, or network topology correlation. Joint modeling of privacy across the ledger layer and network layer is underexplored. Existing protocols often assume independence between transaction content and communication patterns, which adversaries can exploit. For instance, gossip-based mempool propagation leaks significant metadata about the origin and timing of transactions, enabling traffic correlation attacks. Emerging research directions include the design of privacy-preserving mempools that incorporate indistinguishable broadcasting, cover traffic, oblivious relaying, or dandelion-style (Fanti et al., 2018) diffusion to mask transaction origins. These approaches aim to reduce the information available to global or local adversaries observing network traffic (e.g., (Diaz et al., 2021; Bebel and Ojha, 2022)). However, these solutions often introduce latency, bandwidth overhead, or synchronization issues, leading to difficult trade-offs between efficiency, liveness, and anonymity guarantees. Addressing network anonymity in cryptocurrency systems requires cross-layer designs that align messaging protocols with the cryptographic guarantees of the ledger. This is a highly active area of research, particularly in adversarial environments where well-funded surveillance actors may target both layers simultaneously.
6. Conclusion
Designing digital currencies that balance user privacy with institutional auditability remains one of the most intricate and impactful challenges in modern cryptographic and systems engineering. Over the past decade, cryptocurrencies and CBDC prototypes have pushed innovation across cryptographic protocols, PETs, and system architectures. Yet, despite substantial progress, the design space remains fragmented, and no universally accepted blueprint has emerged.
This survey offers a structured and technically grounded overview of the evolving landscape of privacy-preserving digital currencies. We systematize the literature across decentralized cryptocurrencies and CBDCs, introducing a comprehensive taxonomy that connects privacy and auditability goals with underlying cryptographic primitives, protocol designs, and system architectures. In doing so, we bridge conceptual objectives—such as anonymity, confidentiality, and auditability—with concrete technical mechanisms and design models. Our analysis traces the evolution of privacy-preserving digital currencies across three generations, reflecting an increasing interest in addressing privacy-accountability trade-offs. We underscore that privacy is not a monolithic property but a multifaceted design goal shaped by system requirements and regulatory constraints. Finally, we outline several open research challenges that call for interdisciplinary collaboration across cryptography, distributed systems, economics, and policy. Key directions include the development of scalable and composable privacy primitives, architectures that balance auditability with confidentiality, and formal models that capture realistic adversarial and compliance scenarios. As digital currencies transition from academic prototypes to operational systems, ensuring robust privacy alongside regulatory compliance will require continued technical innovation and interdisciplinary coordination. By providing a unified taxonomy and analysis framework, this survey aims to support future research and guide the design of privacy-preserving digital currency systems.
Acknowledgements.
We thank Marco Benedetti, Giuseppe Galano, Sara Giammusso, Antonio Muci, and Fabiana Rossi for their insightful discussions in the early stages of this work and for their valuable feedback throughout its development.References
- (1)
- Almashaqbeh and Solomon (2022) Ghada Almashaqbeh and Ravital Solomon. 2022. SoK: Privacy-Preserving Computing in the Blockchain Era. In Proc. of IEEE EuroS&P’22. IEEE, New York City, US, 124–139.
- Alsalami and Zhang (2019) Nasser Alsalami and Bingsheng Zhang. 2019. SoK: A Systematic Study of Anonymity in Cryptocurrencies. In Proc. of IEEE DSC’19. 1–9.
- Amarasinghe et al. (2019) Niluka Amarasinghe, Xavier Boyen, and Matthew McKague. 2019. A Survey of Anonymity of Cryptocurrencies. In Proc. of ACSW ’19. ACM, Article 2, 10 pages.
- Androulaki et al. (2020) Elli Androulaki, Jan Camenisch, Angelo De Caro, Maria Dubovitskaya, Kaoutar Elkhiyaoui, and Björn Tackmann. 2020. Privacy-preserving auditable token payments in a permissioned blockchain system. In Proc. of AFT’20. ACM, 255–267.
- Androulaki and Karame (2014) Elli Androulaki and Ghassan O. Karame. 2014. Hiding Transaction Amounts and Balances in Bitcoin. In TRUST. Springer, 161–178.
- Androulaki et al. (2013) Elli Androulaki, Ghassan O. Karame, Marc Roeschlin, Tobias Scherer, and Srdjan Capkun. 2013. Evaluating User Privacy in Bitcoin. In Proc. of FC’13. Springer, 34–51.
- Arbabi et al. (2023) Alireza Arbabi, Ardeshir Shojaeinasab, Behnam Bahrak, and Homayoun Najjaran. 2023. Mixing Solutions in Bitcoin and Ethereum Ecosystems: A Review and Tutorial. arXiv:2310.04899
- Auer et al. (2022) Raphael Auer, Rainer Böhme, Jeremy Clark, and Didem Demirag. 2022. Mapping the Privacy Landscape for Central Bank Digital Currencies. Queue 20, 4 (2022), 16–38.
- Auer et al. (2025) Raphael Auer, Rainer Böhme, Jeremy Clark, and Didem Demirag. 2025. Privacy-enhancing technologies for digital payments: mapping the landscape. BIS: Working Papers (2025), 23.
- Baldimtsi et al. (2024) Foteini Baldimtsi, Kostas Kryptos Chalkias, Varun Madathil, and Arnab Roy. 2024. SoK: Privacy-Preserving Transactions in Blockchains. IACR ePrint 2024/1959.
- Baldimtsi et al. (2015) Foteini Baldimtsi, Melissa Chase, Georg Fuchsbauer, and Markulf Kohlweiss. 2015. Anonymous Transferable E-Cash. In Proc. of PKC’15. Springer, 101–124.
- Ballaschk and Paulick (2021) David Ballaschk and Jan Paulick. 2021. The public, the private and the secret: Thoughts on privacy in central bank digital currencies. J. Paym. Strategy Syst. 15, 3 (2021), 277–286.
- Baudet et al. (2023) Mathieu Baudet, Alberto Sonnino, Mahimna Kelkar, and George Danezis. 2023. Zef: Low-latency, Scalable, Private Payments. In Proc. of WPES’23. ACM, 1–16.
- Bauer et al. (2021) Balthazar Bauer, Georg Fuchsbauer, and Chen Qian. 2021. Transferable E-Cash: A Cleaner Model and the First Practical Instantiation. In Proc. of PKC’21. Springer, 559–590.
- Baum et al. (2023) Carsten Baum, James Hsin yu Chiang, Bernardo David, and Tore Kasper Frederiksen. 2023. SoK: Privacy-Enhancing Technologies in Finance. IACR ePrint 2023/122.
- Bebel and Ojha (2022) Joseph Bebel and Dev Ojha. 2022. Ferveo: Threshold Decryption for Mempool Privacy in BFT networks. IACR ePrint 2022/898.
- Bech and Garratt (2017) Morten L Bech and Rodney Garratt. 2017. Central bank cryptocurrencies. BIS Quarterly Review September (2017).
- Beer et al. (2024) Carolin Beer, Sheila Zingg, Kari Kostiainen, Karl Wüst, Vedran Capkun, and Srdjan Capkun. 2024. PayOff: A Regulated Central Bank Digital Currency with Private Offline Payments. arXiv:2408.06956
- Ben-Sasson et al. (2019) Eli Ben-Sasson, Iddo Bentov, Yinon Horesh, and Michael Riabzev. 2019. Scalable Zero Knowledge with No Trusted Setup. In CRYPTO’19. Springer, 701–732.
- Ben Sasson et al. (2014) Eli Ben Sasson, Alessandro Chiesa, Christina Garman, Matthew Green, Ian Miers, Eran Tromer, and Madars Virza. 2014. Zerocash: Decentralized Anonymous Payments from Bitcoin. In Proc. of IEEE S&P’14. 459–474.
- Benarroch et al. (2024) Daniel Benarroch, Bryan Gillespie, Ying Tong Lai, and Andrew Miller. 2024. SoK: Programmable Privacy in Distributed Systems. IACR ePrint 2024/982.
- Bernal Bernabe et al. (2019) Jorge Bernal Bernabe, Jose Luis Canovas, Jose L. Hernandez-Ramos, Rafael Torres Moreno, and Antonio Skarmeta. 2019. Privacy-Preserving Solutions for Blockchain: Review and Challenges. IEEE Access 7 (2019), 164908–164940.
- BIS Innovation Hub (2023) BIS Innovation Hub. 2023. Project Tourbillon: Exploring privacy, security and scalability for CBDCs. Technical Report. BIS.
- Bissias et al. (2014) George Bissias, A. Pinar Ozisik, Brian N. Levine, and Marc Liberatore. 2014. Sybil-Resistant Mixing for Bitcoin. In Proc. of WPES’14. ACM, 149–158.
- Bitansky et al. (2012) Nir Bitansky, Ran Canetti, Alessandro Chiesa, and Eran Tromer. 2012. From extractable collision resistance to succinct non-interactive arguments of knowledge, and back again. In Proc. of ITCS’12. ACM, 326–349.
- Bojja Venkatakrishnan et al. (2017) Shaileshh Bojja Venkatakrishnan, Giulia Fanti, and Pramod Viswanath. 2017. Dandelion: Redesigning the Bitcoin Network for Anonymity. Proc. ACM Meas. Anal. Comput. Syst. 1, 1, Article 22 (2017), 34 pages.
- Boneh et al. (2020) Dan Boneh, Justin Drake, Ben Fisch, and Ariel Gabizon. 2020. Halo Infinite: Recursive zk-SNARKs from any Additive Polynomial Commitment Scheme. IACR ePrint 2020/1536.
- Boneh and Franklin (2003) Dan Boneh and Matthew Franklin. 2003. Identity-Based Encryption from the Weil Pairing. SIAM J. Comput. 32, 3 (2003), 586–615.
- Bonneau et al. (2015) Joseph Bonneau, Andrew Miller, Jeremy Clark, Arvind Narayanan, Joshua A. Kroll, and Edward W. Felten. 2015. SoK: Research Perspectives and Challenges for Bitcoin and Cryptocurrencies. In Proc. of IEEE S&P’15. 104–121.
- Bonneau et al. (2014) Joseph Bonneau, Arvind Narayanan, Andrew Miller, Jeremy Clark, Joshua A. Kroll, and Edward W. Felten. 2014. Mixcoin: Anonymity for Bitcoin with Accountable Mixes. In Proc. of FC’14. Springer, 486–504.
- Bowe et al. (2020) Sean Bowe, Alessandro Chiesa, Matthew Green, Ian Miers, Pratyush Mishra, and Howard Wu. 2020. ZEXE: Enabling Decentralized Private Computation. In Proc. of IEEE S&P’20. 947–964.
- Bünz et al. (2020) Benedikt Bünz, Shashank Agrawal, Mahdi Zamani, and Dan Boneh. 2020. Zether: Towards Privacy in a Smart Contract World. In Proc. of FC’20. Springer, 423–443.
- Bünz et al. (2018) Benedikt Bünz, Jonathan Bootle, Dan Boneh, Andrew Poelstra, Pieter Wuille, and Greg Maxwell. 2018. Bulletproofs: Short Proofs for Confidential Transactions and More. In Proc. of IEEE S&P’18. 315–334.
- Buterin et al. (2013) Vitalik Buterin et al. 2013. Ethereum white paper. https://ethereum.org/it/whitepaper/.
- Béres et al. (2021) Ferenc Béres, István A. Seres, András A. Benczúr, and Mikerah Quintyne-Collins. 2021. Blockchain is Watching You: Profiling and Deanonymizing Ethereum Users. In Proc. of IEEE DAPPS’21. 69–78.
- Camenisch et al. (2006) Jan Camenisch, Susan Hohenberger, and Anna Lysyanskaya. 2006. Balancing Accountability and Privacy Using E-Cash (Ext. Abstr.). In Proc. of SCN’06. Springer, 141–155.
- Camenisch and Lysyanskaya (2003) Jan Camenisch and Anna Lysyanskaya. 2003. A Signature Scheme with Efficient Protocols. In Secur. Commun. Netw. Springer, 268–289.
- Camenisch and Lysyanskaya (2004) Jan Camenisch and Anna Lysyanskaya. 2004. Signature Schemes and Anonymous Credentials from Bilinear Maps. In Proc. of CRYPTO’04. Springer, 56–72.
- Canetti (2000) Ran Canetti. 2000. Universally Composable Security: A New Paradigm for Cryptographic Protocols. IACR ePrint 2000/067.
- Cecchetti et al. (2017) Ethan Cecchetti, Fan Zhang, Yan Ji, Ahmed Kosba, Ari Juels, and Elaine Shi. 2017. Solidus: Confidential Distributed Ledger Transactions via PVORM. In Proc. of ACM SIGSAC CCS’17. ACM, 701–717.
- Chan and Olmsted (2017) Wren Chan and Aspen Olmsted. 2017. Ethereum transaction graph analysis. In Proc. of ICITST’17. 498–500.
- Chase et al. (2014) Melissa Chase, Markulf Kohlweiss, Anna Lysyanskaya, and Sarah Meiklejohn. 2014. Malleable Signatures: New Definitions and Delegatable Anonymous Credentials. In Proc. of IEEE CSF’14. 199–213.
- Chatzigiannis et al. (2021) Panagiotis Chatzigiannis, Foteini Baldimtsi, and Konstantinos Chalkias. 2021. SoK: Auditability and Accountability in Distributed Payment Systems. In Proc. of ACNS’21. Springer, 311–337.
- Chaum (1983) David Chaum. 1983. Blind Signatures for Untraceable Payments. In CRYPTO. Springer, 199–203.
- Chaum et al. (1990) David Chaum, Amos Fiat, and Moni Naor. 1990. Untraceable Electronic Cash. In Proc. of CRYPTO’88. Springer, 319–327.
- Chaum et al. (2021) David Chaum, Christian Grothoff, and Thomas Moser. 2021. How to Issue a Central Bank Digital Currency. arXiv:2103.00254
- Chaum and van Heyst (1991) David Chaum and Eugène van Heyst. 1991. Group Signatures. In Proc. of EUROCRYPT’91. Springer, 257–265.
- Chen et al. (2023) Thomas Chen, Hui Lu, Teeramet Kunpittaya, and Alan Luo. 2023. A Review of zk-SNARKs. arXiv:2202.06877
- Chen et al. (2020) Yu Chen, Xuecheng Ma, Cong Tang, and Man Ho Au. 2020. PGC: Decentralized Confidential Payment System with Auditability. In Proc. of ESORICS’20. Springer, 591–610.
- Christ et al. (2024) Miranda Christ, Foteini Baldimtsi, Konstantinos Kryptos Chalkias, Deepak Maram, Arnab Roy, and Joy Wang. 2024. SoK: Zero-Knowledge Range Proofs. Crypt. IACR 2024/430.
- Diaz et al. (2021) Claudia Diaz, Harry Halpin, and Aggelos Kiayias. 2021. The nym network.
- Dogan and Bicakci (2023) Ali Dogan and Kemal Bicakci. 2023. KAIME : Central Bank Digital Currency with Realistic and Modular Privacy. IACR ePrint 2023/713.
- Dold (2019) Florian Dold. 2019. The GNU Taler system: practical and provably secure electronic payments. Ph. D. Dissertation. Université de Rennes.
- ECB (2019) ECB. 2019. Exploring anonymity in central bank digital currencies. Focus, Dec. (4) 1, 11 (2019).
- El-Hajj and Oude Roelink (2024) Mohammed El-Hajj and Bjorn Oude Roelink. 2024. Evaluating the Efficiency of zk-SNARK, zk-STARK, and Bulletproof in Real-World Scenarios: A Benchmark Study. Information (Switzerland) 15, 8 (2024), 463.
- Elfadul et al. (2024) Issameldeen Elfadul, Lijun Wu, Rashad Elhabob, and Ahmed Elkhalil. 2024. A privacy and compliance in regulated anonymous payment system based on blockchain. J. Ambient Intell. Humaniz. Comput. 15, 8 (2024), 3141–3157.
- Elgamal (1985) T. Elgamal. 1985. A public key cryptosystem and a signature scheme based on discrete logarithms. IEEE Trans. Inf. Theory. 31, 4 (1985), 469–472.
- Ermilov et al. (2017) Dmitry Ermilov, Maxim Panov, and Yury Yanovich. 2017. Automatic Bitcoin Address Clustering. In Proc. of IEEE ICMLA’17. 461–466.
- Fanti et al. (2018) Giulia Fanti, Shaileshh Bojja Venkatakrishnan, Surya Bakshi, Bradley Denby, Shruti Bhargava, Andrew Miller, and Pramod Viswanath. 2018. Dandelion++: Lightweight Cryptocurrency Networking with Formal Anonymity Guarantees. Proc. ACM Meas. Anal. Comput. Syst. 2, 2, Article 29 (2018), 35 pages.
- Fauzi et al. (2019) Prastudy Fauzi, Sarah Meiklejohn, Rebekah Mercer, and Claudio Orlandi. 2019. Quisquis: A New Design for Anonymous Cryptocurrencies. In Proc. of ASIACRYPT’19. Springer, 649–678.
- Feng et al. (2019) Qi Feng, Debiao He, Sherali Zeadally, Muhammad Khurram Khan, and Neeraj Kumar. 2019. A survey on privacy protection in blockchain system. J. Netw. Comput. App. 126 (2019), 45–58.
- Fiat and Shamir (1987) Amos Fiat and Adi Shamir. 1987. How To Prove Yourself: Practical Solutions to Identification and Signature Problems. In Proc. of CRYPTO’86. Springer, 186–194.
- Fleder et al. (2015) Michael Fleder, Michael S. Kester, and Sudeep Pillai. 2015. Bitcoin Transaction Graph Analysis. arXiv 1502.01657.
- Friolo et al. (2024) Daniele Friolo, Geoffrey Goodell, Dann Toliver, and Hazem Danny Nakib. 2024. Private Electronic Payments with Self-Custody and Zero-Knowledge Verified Reissuance. arXiv:2409.01958
- Fujisaki (2011) Eiichiro Fujisaki. 2011. Sub-linear Size Traceable Ring Signatures without Random Oracles. In Proc. of CT-RSA’11. Springer, 393–415.
- Fujisaki and Suzuki (2007) Eiichiro Fujisaki and Koutarou Suzuki. 2007. Traceable Ring Signature. In Proc. of PKC’07. Springer, 181–200.
- Garman et al. (2017) Christina Garman, Matthew Green, and Ian Miers. 2017. Accountable Privacy for Decentralized Anonymous Payments. In Proc. of FC’17. Springer, 81–98.
- Genkin et al. (2018) Daniel Genkin, Dimitrios Papadopoulos, and Charalampos Papamanthou. 2018. Privacy in decentralized cryptocurrencies. Commun. ACM 61, 6 (may 2018), 78–88.
- Ghesmati et al. (2022) Simin Ghesmati, Walid Fdhila, and Edgar Weippl. 2022. SoK: How private is Bitcoin? Classification and Evaluation of Bitcoin Privacy Techniques. In Proc. of ARES’22. ACM, Article 5, 14 pages.
- Goldreich (2004) Oded Goldreich. 2004. Foundations of Cryptography, Volume 2. Cambridge university press Cambridge.
- Goodell et al. (2023) Geoff Goodell, D. R. Toliver, and Hazem Danny Nakib. 2023. A Scalable Architecture for Electronic Payments. In Proc. of FC’22 Workshops. Springer, 645–678.
- Gross et al. (2021) Jonas Gross, Johannes Sedlmeir, Matthias Babel, Alexander Bechtel, and Benjamin Schellinger. 2021. Designing a central bank digital currency with support for cash-like privacy. Available at SSRN 3891121 (2021), 44 pages.
- Groth (2016) Jens Groth. 2016. On the Size of Pairing-Based Non-interactive Arguments. In Proc. of EUROCRYPT’16. Springer, 305–326.
- Guan et al. (2022) Zhangshuang Guan, Zhiguo Wan, Yang Yang, Yan Zhou, and Butian Huang. 2022. BlockMaze: An Efficient Privacy-Preserving Account-Model Blockchain Based on zk-SNARKs. IEEE Trans. Dependable Secure Comput.. 19, 3 (2022), 1446–1463.
- Hayes (1990) Barry Hayes. 1990. Anonymous one-time signatures and flexible untraceable electronic cash. In AUSCRYPT ’90. Springer, 294–305.
- Heilman et al. (2016) Ethan Heilman et al. 2016. TumbleBit: An Untrusted Bitcoin-Compatible Anonymous Payment Hub. Crypt. IACR 2016/575.
- Herrera-Joancomartí (2015) Jordi Herrera-Joancomartí. 2015. Research and Challenges on Bitcoin Anonymity. In Proc. of DPM’15. Springer, 3–16.
- Herskind et al. (2020) Lasse Herskind, Panagiota Katsikouli, and Nicola Dragoni. 2020. Privacy and Cryptocurrencies-A Systematic Literature Review. IEEE Access 8 (2020), 54044–54059.
- Ibrahim (2017) Maged Hamada Ibrahim. 2017. Securecoin: a robust secure and efficient protocol for anonymous bitcoin ecosystem. Int. J. Netw. Secur. 19, 2 (2017), 295–312.
- Kalodner et al. (2020) Harry Kalodner, Malte Möser, Kevin Lee, Steven Goldfeder, Martin Plattner, Alishah Chator, and Arvind Narayanan. 2020. BlockSci: Design and applications of a blockchain analysis platform. In Proc. of Security’20. USENIX, 2721–2738.
- Kappos et al. (2018) George Kappos, Haaroon Yousaf, Mary Maller, and Sarah Meiklejohn. 2018. An Empirical Analysis of Anonymity in Zcash. In Proc. of Security’18. USENIX, 463–477.
- Keller et al. (2021) Patrik Keller, Martin Florian, and Rainer Böhme. 2021. Collaborative Deanonymization. In Proc. of FC’21 Workshops. Springer, 39–46.
- Kiayias et al. (2022) Aggelos Kiayias, Markulf Kohlweiss, and Amirreza Sarencheh. 2022. PEReDi: Privacy-Enhanced, Regulated and Distributed Central Bank Digital Currencies. In Proc. of ACM SIGSAC CCS’22. ACM, 1739–1752.
- Koerhuis et al. (2020) Wiebe Koerhuis, Tahar Kechadi, and Nhien-An Le-Khac. 2020. Forensic analysis of privacy-oriented cryptocurrencies. Forensic Sci. Int.: Digit. Investig. 33 (2020), 200891.
- Kuesters et al. (2013) Ralf Kuesters, Max Tuengerthal, and Daniel Rausch. 2013. The IITM Model: a Simple and Expressive Model for Universal Composability. IACR ePrint 2013/025.
- Kumar et al. (2017) Amrit Kumar, Clément Fischer, Shruti Tople, and Prateek Saxena. 2017. A Traceability Analysis of Monero’s Blockchain. In Proc. of ESORICS’17. Springer, 153–173.
- Le and Gervais (2021) Duc V. Le and Arthur Gervais. 2021. AMR: autonomous coin mixer with privacy preserving reward distribution. In Proc. of ACM AFT’21. ACM, 142–155.
- Li et al. (2022) Rujia Li, Qin Wang, Qi Wang, David Galindo, and Mark Ryan. 2022. SoK: TEE-assisted Confidential Smart Contract. arXiv:2203.08548
- Li et al. (2021) Yannan Li, Guomin Yang, Willy Susilo, Yong Yu, Man Ho Au, and Dongxi Liu. 2021. Traceable Monero: Anonymous Cryptocurrency with Enhanced Accountability. IEEE Trans. Dependable Secure Comput.. 18, 2 (2021), 679–691.
- Liang et al. (2025) Junkai Liang, Daqi Hu, Pengfei Wu, Yunbo Yang, Qingni Shen, and Zhonghai Wu. 2025. SoK: Understanding zk-SNARKs: The Gap Between Research and Practice. arXiv:2502.02387
- Lin et al. (2020) Chao Lin, Debiao He, Xinyi Huang, Muhammad Khurram Khan, and Kim-Kwang Raymond Choo. 2020. DCAP: A Secure and Efficient Decentralized Conditional Anonymous Payment System Based on Blockchain. IEEE Trans. Inf. Forensics Secur. 15 (2020), 2440–2452.
- Liu and Wong (2005) Joseph K. Liu and Duncan S. Wong. 2005. Linkable Ring Signatures: Security Models and New Schemes. In Proc. of ICCSA’05. Springer, 614–623.
- Liu et al. (2018) Yi Liu, Xingtong Liu, Chaojing Tang, Jian Wang, and Lei Zhang. 2018. Unlinkable Coin Mixing Scheme for Transaction Privacy Enhancement of Bitcoin. IEEE Access 6 (2018), 23261–23270.
- Liu et al. (2022) Yunke Liu, Jianbing Ni, and Mohammad Zulkernine. 2022. AT-CBDC: Achieving Anonymity and Traceability in Central Bank Digital Currency. In Proc. of IEEE ICC’22. 4402–4407.
- Lovejoy et al. (2023) James Lovejoy, Madars Virza, Cory Fields, Kevin Karwaski, Anders Brownworth, and Neha Narula. 2023. Hamilton: A High-Performance Transaction Processor for Central Bank Digital Currencies. In Proc. of NSDI’23. USENIX, 901–915.
- Lu et al. (2022) Qinghua Lu, Xiwei Xu, H.M.N. Dilum Bandara, Shiping Chen, and Liming Zhu. 2022. Patterns for Blockchain-Based Payment Applications. In Proc. of EuroPLoP’21. ACM, Article 28, 17 pages.
- Ma et al. (2021) Shunli Ma, Yi Deng, Debiao He, Jiang Zhang, and Xiang Xie. 2021. An Efficient NIZK Scheme for Privacy-Preserving Transactions Over Account-Model Blockchain. IEEE Trans. Dependable Secure Comput.. 18, 2 (2021), 641–651.
- Maxwell (2013) G. Maxwell. 2013. CoinJoin: Bitcoin privacy for the real world. https://bitcointalk.org/index.php?topic=279249.0.
- Meiklejohn and Mercer (2018) Sarah Meiklejohn and Rebekah Mercer. 2018. Möbius: Trustless Tumbling for Transaction Privacy. Proc. Priv. Enhancing Technol. 2018, 2 (2018), 105–121.
- Meiklejohn et al. (2013) Sarah Meiklejohn, Marjori Pomarole, Grant Jordan, Kirill Levchenko, Damon McCoy, Geoffrey M. Voelker, and Stefan Savage. 2013. A fistful of bitcoins: characterizing payments among men with no names. In Proc. IMC’13. ACM, 127–140.
- Meiklejohn et al. (2016) Sarah Meiklejohn, Marjori Pomarole, Grant Jordan, Kirill Levchenko, Damon McCoy, Geoffrey M. Voelker, and Stefan Savage. 2016. A fistful of Bitcoins: characterizing payments among men with no names. Commun. ACM 59, 4 (2016), 86–93.
- Miers et al. (2013) Ian Miers, Christina Garman, Matthew Green, and Aviel D. Rubin. 2013. Zerocoin: Anonymous Distributed E-Cash from Bitcoin. In Proc. of IEEE S&P’13. 397–411.
- Möser et al. (2018) Malte Möser, Kyle Soska, Ethan Heilman, Kevin Lee, Henry Heffan, Shashvat Srivastava, Kyle Hogan, Jason Hennessey, Andrew Miller, Arvind Narayanan, and Nicolas Christin. 2018. An Empirical Analysis of Traceability in the Monero Blockchain. arXiv:1704.04299
- Nakamoto (2008) Satoshi Nakamoto. 2008. Bitcoin: A peer-to-peer electronic cash system. Satoshi Nakamoto (2008).
- Narula et al. (2018) Neha Narula, Willy Vasquez, and Madars Virza. 2018. zkLedger: Privacy-Preserving Auditing for Distributed Ledgers. In Proc. of NSDI’18. USENIX, 65–80.
- Neff (2001) C. Andrew Neff. 2001. A verifiable secret shuffle and its application to e-voting. In Proc. of the ACM CCS’01. ACM, 116–125.
- Noether et al. (2016) Shen Noether, Adam Mackenzie, et al. 2016. Ring confidential transactions. Ledger 1 (2016), 1–18.
- Nosouhi et al. (2023) Mohammad Reza Nosouhi, Shui Yu, Keshav Sood, Marthie Grobler, Raja Jurdak, Ali Dorri, and Shigen Shen. 2023. UCoin: An Efficient Privacy Preserving Scheme for Cryptocurrencies. IEEE Trans. Dependable Secure Comput.. 20, 1 (2023), 242–255.
- Papadoulis et al. (2024) George Papadoulis, Danai Balla, Panagiotis Grontas, and Aris Pagourtzis. 2024. AQQUA: Augmenting Quisquis with Auditability. IACR ePrint 2024/1181.
- Pedersen (1992) Torben Pryds Pedersen. 1992. Non-Interactive and Information-Theoretic Secure Verifiable Secret Sharing. In CRYPTO ’91, Joan Feigenbaum (Ed.). Springer, 129–140.
- Peng et al. (2021) Li Peng, Wei Feng, Zheng Yan, Yafeng Li, Xiaokang Zhou, and Shohei Shimizu. 2021. Privacy preservation in permissionless blockchain: A survey. Digit. Commun. Netw. 7, 3 (2021), 295–307.
- Perera et al. (2022) Maharage Nisansala Sevwandi Perera, Toru Nakamura, Masayuki Hashimoto, Hiroyuki Yokoyama, Chen-Mou Cheng, and Kouichi Sakurai. 2022. A Survey on Group Signatures and Ring Signatures: Traceability vs. Anonymity. Cryptography 6, 1 (2022).
- Pocher and Veneris (2022) Nadia Pocher and Andreas Veneris. 2022. Privacy and Transparency in CBDCs: A Regulation-by-Design AML/CFT Scheme. EEE Trans. Netw. Serv. Manag. 19, 2 (2022), 1776–1788.
- Poelstra (2016) Andrew Poelstra. 2016. Mimblewimble. Technical Report. The University of Rostock.
- Pointcheval and Sanders (2016) David Pointcheval and Olivier Sanders. 2016. Short Randomizable Signatures. In Proc. of CT-RSA’16. Springer, 111–126.
- Qin et al. (2023) Xianrui Qin, Shimin Pan, Arash Mirzaei, Zhimei Sui, Oğuzhan Ersoy, Amin Sakzad, Muhammed F. Esgin, Joseph K. Liu, Jiangshan Yu, and Tsz Hon Yuen. 2023. BlindHub: Bitcoin-Compatible Privacy-Preserving Payment Channel Hubs Supporting Variable Amounts. In Proc. of IEEE S&P’23. 2462–2480.
- Rahalkar and Virgaonkar (2024) Chaitanya Rahalkar and Anushka Virgaonkar. 2024. Summarizing and Analyzing the Privacy-Preserving Techniques in Bitcoin and other Cryptocurrencies. arXiv:2109.07634
- Reserve Bank of Australia (2023) Reserve Bank of Australia. 2023. Australian CBDC Pilot for Digital Finance Innovation. Technical Report. RBA.
- Rivest et al. (1978) Ronald L. Rivest, Leonard Adleman, and Michael L. Dertouzos. 1978. On Data Banks and Privacy Homomorphisms. In Foundations of Secure Computation. Academic Press, 169–179.
- Rivest et al. (2001) Ronald L. Rivest, Adi Shamir, and Yael Tauman. 2001. How to Leak a Secret. In Proc. of ASIACRYPT’01. Springer, 552–565.
- Ron and Shamir (2013) Dorit Ron and Adi Shamir. 2013. Quantitative Analysis of the Full Bitcoin Transaction Graph. In Proc. of FC’13. Springer, 6–24.
- Ruffing and Moreno-Sanchez (2017) Tim Ruffing and Pedro Moreno-Sanchez. 2017. ValueShuffle: Mixing Confidential Transactions for Comprehensive Transaction Privacy in Bitcoin. In Proc. of FC’17. Springer, 133–154.
- Ruffing et al. (2014) Tim Ruffing, Pedro Moreno-Sanchez, and Aniket Kate. 2014. CoinShuffle: Practical Decentralized Coin Mixing for Bitcoin. In Proc. of ESORICS’14. Springer, 345–364.
- Ruffing et al. (2016) Tim Ruffing, Pedro Moreno-Sanchez, and Aniket Kate. 2016. P2P Mixing and Unlinkable Bitcoin Transactions. IACR ePrint 2016/824. https://eprint.iacr.org/2016/824
- Sai Anand and Madhavan (2000) R. Sai Anand and C.E.Veni Madhavan. 2000. An Online, Transferable E-Cash Payment System. In Proc. of INDOCRYPT’00. Springer, 93–103.
- Saleem et al. (2024) Hamza Saleem, Amir Ziashahabi, Muhammad Naveed, and Salman Avestimehr. 2024. Hawk: Accurate and Fast Privacy-Preserving Machine Learning Using Secure Lookup Table Computation. arXiv:2403.17296
- Sander and Ta-Shma (1999) Tomas Sander and Amnon Ta-Shma. 1999. Auditable, Anonymous Electronic Cash. In CRYPTO’99. Springer, 555–572.
- Sarencheh et al. (2022) Amirreza Sarencheh, Aggelos Kiayias, and Markulf Kohlweiss. 2022. PEReDi: Privacy-Enhanced, Regulated and Distributed Central Bank Digital Currencies. IACR ePrint 2022/974.
- Sarencheh et al. (2025) Amirreza Sarencheh, Aggelos Kiayias, and Markulf Kohlweiss. 2025. PARScoin: A Privacy-preserving, Auditable, and Regulation-friendly Stablecoin. In Proc. of CNS’25. Springer Nature Singapore, 289–313.
- Seres et al. (2019) István András Seres, Dániel A. Nagy, Chris Buckland, and Péter Burcsi. 2019. MixEth: efficient, trustless coin mixing service for Ethereum. IACR ePrint 2019/341.
- Setty (2019) Srinath Setty. 2019. Spartan: Efficient and general-purpose zkSNARKs without trusted setup. IACR ePrint 2019/550.
- Shamir (1985) Adi Shamir. 1985. Identity-Based Cryptosystems and Signature Schemes. In CRYPTO’84. Springer, 47–53.
- Shi et al. (2024) Ruisheng Shi, Yulian Ge, Lina Lan, Zhiyuan Peng, Shenwen Lin, and Lin Li. 2024. Deanonymizing Transactions Originating from Monero Tor Hidden Service Nodes. In Companion Proc. of WWW’24. ACM, 678–681.
- Sonnino et al. (2020) Alberto Sonnino, Mustafa Al-Bassam, Shehar Bano, Sarah Meiklejohn, and George Danezis. 2020. Coconut: Threshold Issuance Selective Disclosure Credentials with Applications to Distributed Ledgers. arXiv:1802.07344
- Sun et al. (2021) Xiaoqiang Sun, F. Richard Yu, Peng Zhang, Zhiwei Sun, Weixin Xie, and Xiang Peng. 2021. A Survey on Zero-Knowledge Proof in Blockchain. IEEE Network 35, 4 (2021), 198–205.
- Tomescu et al. (2022) Alin Tomescu, Adithya Bhat, Benny Applebaum, Ittai Abraham, Guy Gueta, Benny Pinkas, and Avishay Yanai. 2022. UTT: Decentralized Ecash with Accountable Privacy. IACR ePrint 2022/452.
- Tran et al. (2018) Muoi Tran, Loi Luu, Min Suk Kang, Iddo Bentov, and Prateek Saxena. 2018. Obscuro: A Bitcoin Mixer using Trusted Execution Environments. In Proc. of the ACSAC’18. ACM, 692–701.
- Valenta and Rowan (2015) Luke Valenta and Brendan Rowan. 2015. Blindcoin: Blinded, Accountable Mixes for Bitcoin. In Proc. of FC’15. Springer, 112–126.
- Van Saberhagen (2013) Nicolas Van Saberhagen. 2013. CryptoNote v 2.0. https://cryptonote.org/whitepaper.pdf.
- Victor (2020) Friedhelm Victor. 2020. Address Clustering Heuristics for Ethereum. In Proc. of FC’20. Springer, 617–633.
- Wang et al. (2020) Dan Wang, Jindong Zhao, and Yingjie Wang. 2020. A Survey on Privacy Protection of Blockchain: The Technology and Application. IEEE Access 8 (2020), 108766–108781.
- Wang et al. (2023) Zhipeng Wang, Marko Cirkovic, Duc V. Le, William Knottenbelt, and Christian Cachin. 2023. Pay Less for Your Privacy: Towards Cost-Effective On-Chain Mixers. IACR ePrint 2023/1222.
- Wüst et al. (2019) Karl Wüst, Kari Kostiainen, Vedran Čapkun, and Srdjan Čapkun. 2019. PRCash: Fast, Private and Regulated Transactions for Digital Currencies. In Proc. of FC’19. Springer, 158–178.
- Wüst et al. (2022) Karl Wüst, Kari Kostiainen, Noah Delius, and Srdjan Capkun. 2022. Platypus: A Central Bank Digital Currency with Unlinkable Transactions and Privacy-Preserving Regulation. In Proc. of ACM SIGSAC CCS’22. ACM, 2947–2960.
- Xue et al. (2023) Liang Xue, Dongxiao Liu, Jianbing Ni, Xiaodong Lin, and Xuemin Sherman Shen. 2023. Enabling Regulatory Compliance and Enforcement in Decentralized Anonymous Payment. IEEE Trans. Dependable Secure Comput.. 20, 2 (2023), 931–943.
- Yu et al. (2024) Qiming Yu, Shilei Liao, Lianhai Wang, Yong Yu, Lingyue Zhang, and Yanqi Zhao. 2024. A regulated anonymous cryptocurrency with batch linkability. Comput. Stand. Interfaces 87 (2024), 103770.
- Zhang et al. (2019) Lingyue Zhang, Huilin Li, Yannan Li, Yong Yu, Man Ho Au, and Baocang Wang. 2019. An efficient linkable group signature for payer tracing in anonymous cryptocurrencies. Future Gener. Comput. Syst. 101 (2019), 29–38.
- Zhang (2023) Tao Zhang. 2023. Privacy Evaluation of Blockchain Based Privacy Cryptocurrencies: A Comparative Analysis of Dash, Monero, Verge, Zcash, and Grin. IEEE Trans. Sustain. Comput. 8, 4 (2023), 574–582.
- Zhong and Mueen (2024) Sheng Zhong and Abdullah Mueen. 2024. BitLINK: Temporal Linkage of Address Clusters in Bitcoin Blockchain. In Proc. of ACM SIGKDD KDD’24. ACM, 4583–4594.
- Ziegeldorf et al. (2018) Jan Henrik Ziegeldorf, Roman Matzutt, Martin Henze, Fred Grossmann, and Klaus Wehrle. 2018. Secure and anonymous decentralized Bitcoin mixing. Future Gener. Comput. Syst. 80 (2018), 448–466.
- Zyskind et al. (2015) Guy Zyskind, Oz Nathan, and Alex Pentland. 2015. Enigma: Decentralized Computation Platform with Guaranteed Privacy. arXiv:1506.03471
Supplementary Material to:
A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability
MATTEO NARDELLI, FRANCESCO DE SCLAVIS, MICHELA IEZZI, Bank of Italy, Italy
Appendix A Appendix
This appendix provides supplementary material for A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability, which complements the main analysis but is excluded from the core sections to preserve narrative flow. Specifically, it includes: (i) a breakdown of publication venues for the works reviewed in this survey, and (ii) a summary of mixing protocols referenced in Section 4.3.5. These materials are presented as tables and offer additional context and technical detail for readers interested in the broader research landscape and protocol-level design trade-offs. Furthermore, in Section A.3, we briefly present interesting yet underexplored cryptographic tools that can be used to design privacy-preserving digital currencies.
A.1. Most Popular Publication Venues
Table 3 summarizes the most common publication venues among the reviewed works. We report only the most frequent venues (with more than works), including journals (e.g., IEEE Transactions on Dependable and Secure Computing, Future Generation Computer Systems), international conferences (e.g., Financial Cryptography and Data Security (FC), ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P)), and other (non-peer reviewed) sources, such as preprints and whitepapers.
| Venue Type | # of Publications | Frequent Venues ( 2 works) | |||||||
| Journal | 13 |
| Conference | 30 |
| Other | 16 |
A.2. Reviewed Mixing Protocols
Similar to Table LABEL:tab:sota-cryptocurrencies, which surveys key privacy-preserving digital currencies, Table LABEL:tab:sota-mixing focuses specifically on mixing protocols. The corresponding legend in Table LABEL:tab:sota_mixing_legend extends the notation introduced in Table 1. These mixing protocols are discussed in Section 4.3.5, and the table serves as a reference to facilitate comparison across existing approaches in the literature.
| Group | Legend | |||
| Protocol | ||||
| No-Trust Assumption | ● | Accountable | ||
| Other | ||||
| Cryptographic | DC-net | Dining cryptographers network | HDC-net | Harmonized DC-net |
| Primitives | MPC | Multi-party computation | TEE | Trusted Execution Environment |
| VS | Verifiable shuffling | |||
| Protocol | Privacy | Other | |||||||||
| Paper |
Year |
Decentralized |
Blockchain Compatibility |
No-Trust Assumptions |
Implemented |
Different Denominations |
Sender Anonymity |
Receiver Anonymity |
Transaction Value Confidentiality |
Unlinkability |
Cryptographic Primitives |
| CoinJoins (Maxwell, 2013) | 2013 | ✓ | ✓ | ✓ | ◐ | DS | |||||
| Xim (Bissias et al., 2014) | 2014 | ✓ | ✓ (B) | ✓ | ✓ | ✓ | PE | ||||
| Mixcoin (Bonneau et al., 2014) | 2014 | ✓ (B) | ● | ✓ | DS | ||||||
| CoinShuffle (Ruffing et al., 2014) | 2014 | ✓ | ✓ (B) | ✓ | ✓ | DS, PE | |||||
| Blindcoin (Valenta and Rowan, 2015) | 2015 | ✓ (B) | ● | ✓ | BS | ||||||
| TumbleBit (Heilman et al., 2016) | 2016 | ✓ (B) | ✓ | ✓ | ✓ | ✓ | ✓ | MPC, ZK | |||
| CoinShuffle++ (Ruffing et al., 2016) | 2016 | ✓ | ✓ (B) | ✓ | ✓212121https://github.com/decred/cspp | ✓ | ✓ | ✓ | DC-net | ||
| SecureCoin (Ibrahim, 2017) | 2017 | ✓ | ✓ (B) | ✓ | ✓ | ✓ | SA | ||||
| ValueShuffle (Ruffing and Moreno-Sanchez, 2017) | 2017 | ✓ | ✓ (B) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | PE, SA |
| Liu et al. (Liu et al., 2018) | 2018 | ✓ | ✓ | ✓ | RS | ||||||
| Möbius (Meiklejohn and Mercer, 2018) | 2018 | ✓ | ✓ (E) | ✓ | ✓ | ✓ | ✓ | ✓ | RS,SA | ||
| Obscuro (Tran et al., 2018) | 2018 | ✓ (B) | ✓ | ✓222222https://github.com/BitObscuro/Obscuro | ✓ | ✓ | TEE | ||||
| CoinParty (Ziegeldorf et al., 2018) | 2018 | ✓ | ✓ (B) | ✓ | ✓ | DSt, Mixnets, MPC | |||||
| MixEth (Seres et al., 2019) | 2019 | ✓ | ✓ (E) | ✓ | ✓232323https://github.com/seresistvanandras/MixEth | ✓ | VS | ||||
| AMR (Le and Gervais, 2021) | 2021 | ✓ | ✓ (E) | ✓ | ✓ | ZK | |||||
| UCoin (Nosouhi et al., 2023) | 2023 | ✓ | ✓ (B) | ✓ | ✓ | HDC-net | |||||
| BlindHub (Qin et al., 2023) | 2023 | ✓ (B) | ✓ | ✓242424https://github.com/blind-channel/blind-hub | ✓ | ✓ | BS* | ||||
| Wang et al. (Wang et al., 2023) | 2023 | ✓ | ✓ | ✓ | ✓ | ✓ | MT | ||||
A.3. Emerging Cryptographic Tools for Future Digital Currencies
In addition to the cryptographic tools discussed in the main survey, several advanced primitives offer promising tools for future research in privacy-preserving digital currencies. Although not yet widely adopted in practical systems, these techniques may enable more flexible or robust privacy features in regulated or decentralized contexts.
A.3.1. Anonymous Credentials
[…]
A.3.2. Cryptographic Accumulators
Accumulators provide compact representations of sets along with efficient membership proofs. They are useful for constructing privacy-preserving mechanisms such as spent coin tracking, revocation, and anonymous group membership verification. Although some systems already use Merkle trees to prove coin or user inclusion in a list of valid or authorized entities (e.g., (Sander and Ta-Shma, 1999; Gross et al., 2021; Xue et al., 2023; Wang et al., 2023; Elfadul et al., 2024)), we believe that accumulators remain underutilized—particularly in the design of revocation mechanisms, where they could offer efficient and privacy-preserving solutions.
A.3.3. Functional Encryption
[…]
A.3.4. Time-Lock Encryption
[…]
A.3.5. Witness Encryption
[…]
These emerging primitives highlight exciting directions for foundational research in the design of privacy-preserving digital currencies. While most remain experimental, their potential to address regulatory, architectural, or usability gaps makes them worth further exploration.
[…]