### v3

# A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability

Abstract.
Crypto-assets and central bank digital currencies (CBDCs) are reshaping how value is exchanged in distributed computing environments. These systems combine cryptographic primitives, protocol design, and system architectures to provide transparency and efficiency while raising critical challenges around privacy and regulatory compliance. This survey offers a comprehensive overview of privacy-preserving digital payment systems, covering both decentralized ledger systems and CBDCs. We present a taxonomy of privacy goals—including anonymity, confidentiality, unlinkability, and auditability—and map them to the cryptographic primitives, protocols, and system architectures that implement them. Our work adopts a design-oriented perspective, linking high-level privacy objectives to concrete implementations. We also trace the evolution of privacy-preserving digital payment systems through three generations, highlighting shifts from basic anonymity guarantees toward more nuanced privacy-accountability trade-offs. Finally, we identify open challenges, motivating further research into architectures and solutions that balance strong privacy with real-world auditability needs.
1. Introduction
The rapid proliferation of digital payment systems has sparked renewed interest in designing distributed architectures that balance user privacy with institutional accountability. Spanning decentralized crypto-assets [^1] The term cryptocurrency is widely used to describe decentralized systems such as Bitcoin, Ethereum, or Monero, that enable peer-to-peer asset transfer without a central authority. However, this usage differs from the terminology adopted by central banks and regulators (ECB, 2018), which reserve the term “currency” for instruments with legal tender status. To reflect this distinction and bridge the terminology gap between these communities, we adopt the more nuanced terms crypto-assets or digital payment systems throughout the paper. and centrally managed infrastructures such as central bank digital currencies (CBDCs), these systems leverage cryptographic techniques and distributed consensus to enable secure, efficient, and programmable value transfer. However, the same transparency and auditability that make these systems attractive also raise fundamental questions about user privacy, institutional control, and regulatory compliance (e.g., (Auer et al., 2025; Ghesmati et al., 2022; Herskind et al., 2020)). For example, public blockchains typically offer pseudonymity. [^2] Anonymity denotes indistinguishability within a set, while pseudonymity uses identifiers that are not directly linked to real-world identities. This enables transaction tracing, address clustering, and ultimately user de-anonymization, thus exposing financial behavior (e.g., (Meiklejohn et al., 2013; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020; Karame et al., 2015)). As a result, privacy-preserving digital payments have emerged as a critical research field, attracting attention from cryptographers, engineers, economists, and policymakers.
Researchers have proposed a variety of privacy-preserving solutions leveraging privacy-enhancing technologies (PETs), including mixing protocols (e.g., (Maxwell, 2013; Ziegeldorf et al., 2018; Ruffing and Moreno-Sanchez, 2017)), confidential transactions (e.g., (Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022)), and privacy-native crypto-assets (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Tomescu et al., 2022; Noether et al., 2016)). These approaches provide different privacy guarantees, such as anonymity, confidentiality, and unlinkability. However, privacy alone is not sufficient: regulatory frameworks such as Anti-Money Laundering (AML) and Combating the Financing of Terrorism (CFT) require mechanisms for auditability and selective disclosure. To reconcile these conflicting goals, recent work has explored cryptographic tools such as zero-knowledge proofs (ZKPs), threshold encryption, and homomorphic commitments.
This survey proposes a unifying, design-oriented perspective on privacy-preserving digital payment systems, covering both decentralized ledgers and centrally managed systems with privacy guarantees—including CBDCs. It organizes the literature around conceptual dimensions that highlight recurring design choices and trade-offs. Unlike previous surveys that focus on broad overviews (e.g., (Genkin et al., 2018; Auer et al., 2025; Herskind et al., 2020)), specific technologies (e.g., (Almashaqbeh and Solomon, 2022; Ghesmati et al., 2022; Sun et al., 2021)), or do not consider auditability (e.g., (Peng et al., 2021; Herskind et al., 2020; Feng et al., 2019)), our work aims to bridge these dimensions. Our contributions are threefold. First, we introduce a taxonomy that classifies privacy-preserving digital payments systems along multiple dimensions, including system architecture, trust model, value representation, privacy guarantees, and auditability mechanisms. Second, we systematize how privacy properties are formalized and realized, mapping high-level goals to technical mechanisms. We trace the development of privacy-preserving digital payment systems, highlighting the role of cryptographic techniques across three generations. Third, we identify open research challenges stemming from the design of privacy-preserving digital payment systems, which lie at the intersection of cryptography, distributed systems, and privacy engineering. Supplementary material presents the methodology used to collect and select the surveyed literature, and further contextualizes key publication venues and technical details of specific privacy-enhancing mechanisms.
This paper is organized as follows. Section 2 reviews related surveys and outlines how our work complements and differs from them. Section 3 introduces the key cryptographic primitives used in privacy-preserving digital payments, which are then analyzed along key architectural and cryptographic dimensions in Section 4. Then, Section 5 outlines open challenges and future research directions, and Section 6 concludes the paper.
2. Related Surveys
A large body of literature has explored privacy-preserving digital payment infrastructures, driven by both the technical challenges and societal concerns about financial surveillance.
Privacy in Crypto-assets
The rapid growth of crypto-asset systems has spurred multiple efforts to systematize knowledge (SoK) across diverse technical communities. Bonneau et al. (Bonneau et al., 2015) reviewed extensions of Bitcoin and alternative protocols (sometimes referred to as altcoins), including early privacy-native protocols. Genkin et al. (Genkin et al., 2018) offered a broad and accessible overview of techniques for anonymity in decentralized crypto solutions, while emphasizing the absence of a unified definition of privacy. More recent surveys attempted to formalize anonymity notions. For instance, Amarasinghe et al. (Amarasinghe et al., 2019) developed a taxonomy of anonymity features (including anonymity, confidentiality, unlinkability, untraceability, and deniability), and evaluated Bitcoin-derived solutions against it. Our work draws from and extends this analysis by further considering the trade-offs between anonymity, confidentiality, and auditability across a broader range of systems, including CBDCs. Feng et al. (Feng et al., 2019) offered a layered classification of privacy-preserving mechanisms from network-level obfuscation to cryptographic primitives such as ring signatures, Pedersen commitments, homomorphic encryption, and NIZKs. While technically detailed, their coverage of payment protocols and privacy-accountability trade-offs is limited. Similarly, a systematic literature review by Herskind et al. (Herskind et al., 2020) catalogs techniques such as stealth addresses, confidential transactions, and network anonymity. However, it does not analyze core system features such as value representation models or system architectures, which are key components of our taxonomy. Peng et al. (Peng et al., 2021) identify PETs and illustrate how various protocols integrate them, but do not investigate protocol evolution and design trade-offs.
Other related surveys take a narrower or orthogonal perspective. Almashaqbeh and Solomon (Almashaqbeh and Solomon, 2022) focus primarily on privacy in blockchain computation, i.e., smart contracts; Zhang (Zhang, 2023) and Wang et al. (Wang et al., 2020) overview privacy features in deployed public ledger-based systems; Alsalami and Zhang (Alsalami and Zhang, 2019) present an analysis of PETs, not considering accountability.
Privacy in CBDCs
A distinct body of work addresses privacy in CBDCs, driven by the need to balance individual privacy with institutional oversight. Auer et al. (Auer et al., 2022, 2025) distinguish between soft privacy, based on access controls and trust in intermediaries, and hard privacy, based on cryptographic techniques. While they offer a high-level classification and conceptual framework, their discussion does not consider low-level protocol mechanisms. Other regulatory-focused contributions include Darbha and Arora (Bank of Canada) [^3] https://www.bankofcanada.ca/2020/06/staff-analytical-note-2020-9/ and Pocher and Veneris (Pocher and Veneris, 2022) explore PETs from a compliance and governance perspective, mapping privacy mechanisms to AML/CFT auditability requirements in retail CBDC designs. These works offer blueprints but lack detailed cryptographic analysis. Conversely, our work presents a comparative review of technical design choices including primitives, protocols, and architectures.
Specialized Surveys
Several surveys focus on specific technologies (e.g., Bitcoin (Ghesmati et al., 2022), smart contracts (Benarroch et al., 2024; Li et al., 2022), DeFi (Baum et al., 2023), mixers (Arbabi et al., 2023), accountability mechanisms (Chatzigiannis et al., 2021), and ZKPs (Sun et al., 2021; Almashaqbeh and Solomon, 2022; Liang et al., 2025)). These contributions do not consider the full design space of privacy-preserving payment systems as we do, nor do they explore how different privacy features interact—particularly in settings requiring auditability. Koerhuis et al. (Koerhuis et al., 2020) present a forensic analysis of Monero and Verge, and highlight how implementation flaws threaten privacy.
Contribution of This Survey
We propose a unified taxonomy of privacy-preserving payment systems that integrates both decentralized ledger-based solutions and CBDCs. Specifically, we present a systematic analysis of existing solutions to identify key value representation models, privacy features, i.e., anonymity, confidentiality, accountability, and auditability guarantees. We focus on how such features are realized through cryptographic protocols and specific design choices, resulting in a classification of contributions that connects design goals to technical implementations.
3. Background
We first introduce key cryptographic primitives (Sections 3.13.3), and then we briefly review key cryptographic protocols used in privacy-preserving digital payment systems (Section 3.43.6).
3.1. Commitments and Homomorphic Commitments
A commitment scheme is a cryptographic primitive that allows a party to commit to a value in such a way that it cannot be changed later. The scheme consists of two phases: (i) commit phase, where a commitment is made to a value $v$, but the value is kept secret from third parties; (ii) opening phase, where the committed value $v$ is revealed. A commitment is hiding if an adversary cannot distinguish which of two values corresponds to the commit; it is binding if the committer cannot generate the same commit from two different values. These properties can be either computational, if the adversary is computationally bounded, or perfect, if not. A commitment scheme cannot be both perfectly hiding and perfectly binding, due to fundamental cryptographic limitations (Goldreich, 2004).
Homomorphic commitments preserve the addition or multiplication of corresponding values. For example, let $g$ and $h$ be elements of a prime $q$ group $\mathbb{G}_{q}$ such that $\log_{g}h$ is not feasible to compute, let $r_{i}$ be a random value over $\mathbb{Z}_{q}$ and let $v_{i}\in\mathbb{Z}_{q}$ be the committer’s value. Perdersen commitments are defined as $c_{v_{i}}=g^{v_{i}}h^{r_{i}}$. The homomorphic properties of Pedersen commitments guarantee that $g^{v_{1}}h^{r_{1}}\cdot g^{v_{2}}h^{r_{2}}=g^{v_{1}+v_{2}}h^{r_{1}+r_{2}}$ namely, the sum of the commitments of $v_{1}$ and $v_{2}$ is equal to the commitment of the sum of the two values under the sum of the two randomness $r_{1}$ and $r_{2}$, allowing multiple operations on the committed messages. In particular, Pedersen (Pedersen, 1992) commitments are perfectly hiding, computationally binding, and additively homomorphic; ElGamal (Elgamal, 1985) commitments are computationally hiding, perfectly binding, and multiplicatively homomorphic.
3.2. Encryption Schemes
Identity-based encryption (IBE)
This scheme (Shamir, 1985) is a public-key encryption scheme in which the public key is a string that uniquely identifies a user (e.g., an email). The corresponding private keys for decryption are derived from a master private key, by a trusted central authority. IBE can be used to simplify key management and recover, and access control within an organization. Anonymous variants make the public key indistinguishable from the ciphertext, effectively hiding the identity of the recipient of the message.
Threshold encryption
It allows data to be encrypted in such a way that it can only be decrypted when a predefined quorum of participants agrees to reveal it. Threshold encryption can help maintain privacy while removing single point of failures. A popular implementation uses the ElGamal encryption scheme (Elgamal, 1985). A threshold-issuance variant of Boneh-Franklin IBE is used in UTT (Tomescu et al., 2022), to distribute the master secret key among multiple authorities.
Homomorphic encryption
This technique (Rivest et al., 1978) allows computing additions, multiplications or both, without revealing the plaintext. Operations can be performed directly on ciphertexts such that the result corresponds to the encryption of the same operation applied to the plaintexts, i.e., $\mathrm{Enc}(m_{1})\star\mathrm{Enc}(m_{2})=\mathrm{Enc}(m_{1}\star m_{2})$, where $\mathrm{Enc}(\cdot)$ represent an encryption function and $\star$ denotes either addition or multiplication. This allows to delegate computation to a third party while maintaining confidentiality. It enables confidential transactions when used as a building block for ZKPs, like to homomorphic commitments (e.g., (Ma et al., 2021; Bünz et al., 2020)).
3.3. Zero-knowledge Proofs
A ZKP is a protocol enabling a prover to convince a verifier that a statement is true, without disclosing any information beyond the validity of the statement itself. A dishonest prover can convince the verifier of a false statement, only with negligible probability (soundness). If the prover is computationally bounded (computational soundness), the proof is known as argument (Bitansky et al., 2012). ZKPs can be either interactive, requiring the parties to exchange messages, or non-interactive (NIZK), when the verifier is convinced through a single message. ZKPs can also be proofs of knowledge when the prover can convince the verifier that he knows a certain value, e.g., a secret key (or arguments of knowledge, if the prover is computationally bounded). If the ZKP is short and fast to verify, it is said to be succinct. If it requires no trusted setup, it is transparent.
A conceptual framework for ZK protocols is the $\Sigma$-protocol, which is a type of 3-move interactive proof system: the prover commits a value, the verifier provides a challenge, and the prover provides a proof based on the commit and the challenge, so that they cannot choose a “malicious” proof. Using the Fiat-Shamir transform, $\Sigma$-protocols can be made non-interactive (Fiat and Shamir, 1987). Building on this, various NIZK protocols have been designed, including zk-SNARK (Bitansky et al., 2012), Bulletproofs (Bünz et al., 2018), and—to some extent—zk-STARK (Ben-Sasson et al., 2019). Zk-SNARKs are succinct non-interactive arguments of knowledge. Many zk-SNARK constructions have been proposed in past years (e.g., (Bitansky et al., 2012; Boneh et al., 2020; Groth, 2016; Chen et al., 2023)), with different approaches also for setup. Among them, the most popular approach is Groth16 (Groth, 2016), an instance optimized for performance but with some trade-offs like requiring a trusted setup. Bulletproofs are optimized for range proof and does not require trusted setup. Zk-STARK is a recent approach offering efficient proving times and no need for trusted setup. STARKs are succinct transparent arguments of knowledge. However, in the current state, zk-STARK generates large proof size (a few hundred kilobytes), so they are not as succinct as SNARKs. A recent survey on ZKPs can be found, e.g., in (Christ et al., 2024; El-Hajj and Oude Roelink, 2024).
3.4. Anonymity-enhanced Signatures
Blind signatures
Blind signatures (Chaum, 1983) are a form of digital signature in which the signer produces a valid signature on a message without learning its content: before signing, the message is blinded, i.e., transformed through a random element that makes it unrecognizable. Once the contents of the signed message are unblinded, they can be publicly verified as usual.
Randomizable signatures
They provide a way to randomize a previously generated signature, transforming it into a new valid signature, but making it impossible to link the two instances together. This enables a level of anonymity by unlinking user identities from their transaction history, as each transaction appears distinct even if performed by the same user. The original idea was proposed by Camenisch and Lysyanskaya (Camenisch and Lysyanskaya, 2003, 2004) but suffered from a linear size in the number of messages to be signed. Later, Pointcheval-Sanders proposed more efficient signature and verification algorithms (Pointcheval and Sanders, 2016). Existing works (e.g., (Baudet et al., 2023; Wüst et al., 2019; Sarencheh et al., 2025; Tomescu et al., 2022)) adopt either Pointcheval-Sanders or Coconut (Sonnino et al., 2020), a threshold scheme derived from it. The Pointcheval-Sanders scheme can produce signatures of a message directly from its commitment; this effectively makes it a blind signature, thanks to the hiding property of commitments.
Group signatures
This scheme (Chaum and van Heyst, 1991) enables any member of a designated group to generate a valid signature without revealing who signed (ensuring anonymity). A verifier uses a group public key to confirm that a valid signature was produced by some group member. This property also implies unlinkability: it is not possible to determine whether two different signatures were produced by the same user (Perera et al., 2022). However, group signatures can be made linkable to prevent double spending in transactions (e.g., (Zhang et al., 2019)). Group signatures also support user traceability: a designated group manager holds a group secret key that enables him to revoke a signer’s anonymity and reveal his identity when necessary.
Ring signatures
Ring signatures (Rivest et al., 2001) are similar to group signatures, as they hide the signer’s identity behind a group. Differently from group signatures, user anonymity is permanent, because there is no revoke mechanism. Indeed, the signer simply joins their public key with a set of other public keys during the signing phase. A verifier needs the whole list of public keys to verify the signature validity, and cannot discern which of the $n$ public keys is of the signer, but can only guess with probability $1/n$. There are some variant ring signatures, where anonymity is somewhat limited and the signer can be traced under special circumstances, e.g., after double spending. These are called traceable ring signatures  (Fujisaki and Suzuki, 2007; Fujisaki, 2011). Ring signatures can also be linkable (Liu and Wong, 2005), if they allow anyone to determine if two signatures were produced by the same signer (e.g., (Van Saberhagen, 2013; Li et al., 2021)).
Threshold signatures
They are a type of signature that can be jointly produced by a committee of signers. A $t$-of-$n$ signature scheme requires at least $t$ signers out of a total of $n$ participants to produce a valid signature. Works that use threshold signatures (Androulaki et al., 2020; Sarencheh et al., 2025; Kiayias et al., 2022; Baudet et al., 2023) use threshold blind signatures, based on Coconut (Sonnino et al., 2020).
3.5. Secure Multi-party Computation
Secure Multi-Party Computation (MPC) allows multiple parties to jointly compute a result over their private inputs without any party having to reveal its own input to the others. The key idea is that the parties collaborate to perform some computation and, at the end, they all learn the correct output of the function, but nothing else. Beyond applications such as CoinParty (Ziegeldorf et al., 2018), MPC is also used to decentralize trusted setup procedures in zk-SNARKs, enabling security to hold as long as at least one participant behaves honestly.
3.6. One-time Public Keys
Stealth addresses
They hide the recipient’s identity making each payment unlinkable. They are one-time public keys generated by the transaction sender with public key provided by the receiver. Recipient does not update his public key, but the sender can derive fresh one-time keys for each transaction. Let $\mathrm{G}$ be a generator of an elliptic curve group $\mathbb{G}$ with scalar field $\mathbb{Z}_{q}$. Each receiver $R$ has a public view key $V=vG$ and a public spend key $S=sG$, where $v,s\in\mathbb{Z}_{q}$ are secret scalars. A sender chooses a random scalar $r\in\mathbb{Z}_{q}$ and computes the shared secret $H=\mathcal{H}_{p}(rV)\in\mathbb{Z}_{q}$, with $\mathcal{H}_{p}(\cdot)$ a hash-to-scalar function, and the one-time public key $P=HG+S$. The sender includes $P$ and $R=rG$ in the transaction. The recipient can scan the ledger and for each $P$ and $R$ computes $P^{\prime}=\mathcal{H}_{p}(vR)G+S$. If $P^{\prime}=P$ then the output belongs to the recipient.
Updatable public keys
They allow to update the recipient’s public key in a decentralized manner. Quisquis (Fauzi et al., 2019) introduces this technique to update both sender and recipient public keys. Given a prime-order group $\mathbb{G}_{p}$ with $p$ the prime order of the group, and $G$ its generator, each user has a secret key $s_{i}$ and a public key $P_{i}=(G^{r},G^{r\cdot s_{i}})=(U,V)$, for some random value $r\in\mathbb{Z}_{q}$. To perform a transaction, the sender updates the public key by choosing a random value $z\in\mathbb{Z}_{q}$ and computing $P^{\prime}_{i}=(U^{z},V^{z})$. The updated key is indistinguishable from a fresh key, thanks to the Decisional Diffie-Hellman assumption. The recipient verifies whether a given updated public key $P^{\prime}_{i}$ corresponds to his secret key $s_{i}$. This works because updated keys still verify correctly with the original secret key.
Figure 1. Dimension of analysis along which we review privacy-preserving digital payment systems.
Dimension of analysis along which we review privacy-preserving digital payment systems.
Table 1. Labels and notation used in Table LABEL:tab:sota-cryptocurrencies.
Dimension Legend Protocol Type CD CBDC D Digital payment system Blockchain Compatibility B Bitcoin E Ethereum Value Representation Data Model A Account-based CT Chaumian token T2B Transaction per bank table UT Unspent token Other Auditability AB Anonymity budget L Operating limits RA Revocable anonymity TR Coin tracing ZD Verification without disclosure Cryptographic Primitives BS Blind signature C Commitment DS Digital signature GS Group signatures HC Homomorphic commitment IBE Identity-based encryption MS Multi signature MT Merkle tree O-RAM Oblivious RAM PE Public key encryption RS Ring signature SA Stealth address UP Updatable public keys ZK Zero-knowledge proof $X^{t}$ Threshold-based version of $X$ $X^{r}$ Randomizable version of $X$ XX* Custom variant of $X$ Notation Yes $\times$ No Partial $\Leftrightarrow$ Alternative Abstracted
4. Design Dimensions of Privacy-Preserving Digital Payment Systems
The research community has explored various solutions to ensure anonymity, confidentiality, and auditability in crypto-assets and digital payment systems. We identify key dimensions of analysis—protocol, value representation, privacy features, and auditability—and highlight the most significant design choices. [^4] To avoid terminological ambiguity, we align with De Roode and Everts (de Roode and Everts, 2025), who distinguish privacy (a user right) from privacy-preservation (a system property) and define anonymity, unlinkability, and auditability in a unified DLT terminology. These dimensions are interconnected and jointly embed privacy features into the system. Figure 1 outlines the dimensions that structure this section. Table 1 describes the labels used in Table LABEL:tab:sota-cryptocurrencies, which summarizes our literature review. [^5] Some properties shown in Figure 1 and Table LABEL:tab:sota-cryptocurrencies (e.g., double-spending protection, ledger state updates) are treated as supporting mechanisms underpinning correctness, privacy, or auditability, and are discussed within those dimensions rather than as independent classification axes.
4.1. Protocol
The protocol dimension outlines the main system design choices through which existing solutions can be compared.
4.1.1. Generations of Privacy-preserving Digital Payment Systems
We introduce the notion of generations of privacy-preserving digital payment systems to group works that share similar assumptions, deployment contexts, and design goals. Under this lens, first-generation systems typically rely on centralized issuers and focus on payer anonymity; second-generation systems emphasize decentralized value transfer, unlinkability, and untraceability; while third-generation systems balance privacy with auditability and regulatory compliance. [^6] These groupings are not strictly chronological, and several systems naturally exhibit hybrid characteristics across generations. The “generations” serve as conceptual lens to interpret recurring design goals and assumptions across the dimensions discussed in this section.
First Generation
The first generation of privacy-preserving digital payment systems focuses on centralized value transfer architectures, where a trusted issuing authority (typically a bank) is responsible for minting and redeeming digital coins, while cryptographic mechanisms are used to protect the payer’s anonymity during spending. The central challenge in this setting is reconciling the need for user anonymity with the requirement that the issuer validate, redeem, and potentially audit issued tokens. Chaum (Chaum, 1983) introduces the seminal untraceable payment system, known as ecash [^7] Later, Chaum et al. (Chaum et al., 2021) show how these mechanisms can be embedded into a two-tier, regulator-compliant retail CBDC architecture that preserves cash-like payer privacy while enforcing income transparency and AML/CFT through commercial banks., which employs blind signatures to prevent the issuing bank from linking withdrawn coins to subsequent payments (see Section 4.2 for details). Building on this foundation, a large body of work explored extensions of ecash (Andola et al., 2021), including different coin denominations (e.g., (Dold, 2019)), partial spending (e.g., (Dold, 2019; Okamoto and Ohta, 1992)), token transferability (e.g., (Hayes, 1990; Sai Anand and Madhavan, 2000; Baldimtsi et al., 2015; Bauer et al., 2021; Chase et al., 2014)), offline spending (e.g., (Chaum et al., 1990)), and controlled forms of auditability (e.g., (Camenisch et al., 2006; Schoenmakers, 1998; Brands, 1994; Camenisch and Lysyanskaya, 2001; Batten and Yi, 2019)). [^8] In contrast to Chaum’s ecash, Netcoin (Tewari and Nuallain, 2015) proposes a fiat-backed mint-based system in which coins are bound to user identities and fully traceable (no anonymity). This highlights how privacy was not a universal requirement in early centralized ecash designs.
Second Generation
The second generation emerged with the introduction of Bitcoin (Nakamoto, 2008). Bitcoin and, later, Ethereum (Buterin et al., 2013) propose public and permissionless blockchains, meaning that the ledger contains transactions visible to everyone and anyone can participate in the process of updating the public ledger. Being public, these blockchains use addresses to exchange assets in a pseudo-anonymous manner, as addresses are not directly tied to user identities. [^9] Bitcoin addresses are designed to be ephemeral (though often reused), whereas Ethereum relies on persistent pseudonymous accounts. Public blockchains cannot easily and fully satisfy two privacy-related features: untraceability and unlinkability (Rahalkar and Virgaonkar, 2024). Untraceability ensures that, for any given incoming transaction, all potential senders appear equally likely. Unlinkability means that, for any two outgoing transactions, there is no way to prove they were sent to the same recipient. Untraceability and unlinkability can arise at different layers: at the ledger level, they refer to preventing correlations between on-chain records; at the network level, they refer to preventing correlations from communication patterns [^10] Unless otherwise stated, we primarily focus on ledger-level privacy, and discuss network-layer aspects in Section 4.3.4.. While generating a new address for each transaction can enhance these properties, various studies indicate that achieving true anonymity and unlinkability on public blockchains remains challenging. An analysis of the transaction graph can reveal clusters of pseudonyms or patterns that link with high probability two or more addresses (e.g., (Meiklejohn et al., 2013; Möser et al., 2018; Kappos et al., 2018; Herskind et al., 2020)). The second generation of solutions mainly focuses on improving unlinkability, anonymity, and confidentiality, often considering decentralized architectures (e.g., (Miers et al., 2013; Van Saberhagen, 2013; Fauzi et al., 2019; Baudet et al., 2023; Friolo et al., 2026; Sarencheh et al., 2025; Androulaki et al., 2013; Herrera-Joancomartí, 2015; Meiklejohn et al., 2013)).
We classify this second generation of approaches into three main categories: privacy-native protocols, mixing protocols, and off-ledger solutions. Changing the protocol of existing blockchains is hard, as it involves forks. Therefore, privacy-native protocols (mostly) defines novel solutions with improved confidentiality, anonymity, and unlinkability features (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Ben Sasson et al., 2014; Tomescu et al., 2022)). Conversely, mixing protocols propose supplementary services for permissionless blockchains aimed to collect a group of transactions, scramble senders and receivers, and publish them as a single group transaction (e.g., (Maxwell, 2013; Bissias et al., 2014; Bonneau et al., 2014; Valenta and Rowan, 2015; Ruffing and Moreno-Sanchez, 2017; Ziegeldorf et al., 2018; Qin et al., 2023; Wang et al., 2023)). Mixing protocols break the link between sender and receiver, thus reducing the possibility of linking together addresses of the blockchain (an analysis of protocols in Appendix A.3.2). Off-ledger solutions are a specific approach related to blockchains (that serve as Layer 1) and the ability to build decentralized peer-to-peer Layer 2 systems on top of them. Layer 2 solutions process transactions off the main chain, reducing the need to record every transaction on the shared ledger (e.g., (Qin et al., 2023; Wang et al., 2023; Baudet et al., 2023; Seres et al., 2019)). Focusing on on-ledger privacy, we omit the analysis of Layer 2 privacy solutions.
Third Generation
While protecting the privacy and anonymity of payers and payees remains essential, the third generation of solutions introduces mechanisms that balance these guarantees with auditability and regulatory compliance (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Narula et al., 2018; Li et al., 2021; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Yu et al., 2024; Xue et al., 2023)). These approaches are motivated by concerns around AML and CFT. Auditability refers to the system’s ability to track, verify, or selectively disclose information about transactions, without fully compromising user privacy. This enables the enforcement of specific policies that apply to users, coins, or transactions behaviors over time. To this end, modern privacy-preserving systems integrate mechanisms such as revocable anonymity, trapdoors, and ZKPs for correctness and compliance.
4.1.2. System Architecture
We examine the architecture of privacy-preserving payment systems and protocols.
Centralized Systems
The first generation of privacy preserving solutions typically rely on centralized architectures involving up to four trusted authorities, each responsible for distinct functions. They include: (i) a central bank, which issues digital coins and maintains monetary control; (ii) a tracing authority, responsible for auditing and tracing under specific conditions; (iii) a registration authority, in charge of user on-boarding and identity verification; and, in some cases, (iv) a technical support entity, which collaborates with the tracing authority to enable conditional de-anonymization. Most early proposals assume a single central bank (e.g., (Chaum, 1983; Dold, 2019; Cecchetti et al., 2017; Zhang et al., 2019; Beer et al., 2024)). For privacy, special cryptographic primitives are used, such as blind signatures (Chaum, 1983), which allow the bank to issue valid coins without learning user identities. To support auditability and regulatory compliance, systems often include a separate tracing authority that can decrypt transactions or enforce specific policies (e.g., check or bypass an anonymity budget). This entity operates independently from the central bank, thereby preserving some separation of duties (e.g., (Sander and Ta-Shma, 1999; Wüst et al., 2022)). Yu et al. (Yu et al., 2024) introduce a technical support authority that must cooperate with the tracing authority to extract sender identities. In the context of CBDCs, system architectures usually include a registration authority, responsible for enrolling user registration and managing identity records (e.g., (Zhang et al., 2019))—an essential requirement considering regulatory frameworks (e.g., AML and CFT).
Decentralized Systems
Blockchains introduced a fundamental shift in system design by leveraging decentralization, which serves two main purposes. First, it enhances resilience against malicious actors or censorship by distributing trust across multiple entities. Second, it enables intermediary-free value exchange, using commitments and digital signatures to ensure integrity, authenticity, and non-repudiation. Many privacy-native solutions adopt decentralized architectures in line with these principles (e.g., (Noether et al., 2016; Van Saberhagen, 2013; Poelstra, 2016; Androulaki et al., 2020; Narula et al., 2018; Ma et al., 2021; Xue et al., 2023)). Furthermore, the programmability of blockchains has enabled privacy-preserving overlays (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)) and smart contracts [^11] Although logically centralized, smart contracts are executed on each node; therefore, we consider them as a decentralized solution. implementing private tokens “pegged” to the coin of the primary chain (e.g., (Bünz et al., 2020; Baudet et al., 2023; Sarencheh et al., 2025)).
4.1.3. Protocol Integration and Compatibility with Existing Blockchains
Second-generation privacy-preserving payment systems vary in their integration with existing (blockchain) infrastructures: some remain compatible, others introduce non-compatible extensions, and some adopt entirely new designs.
Compatible Solutions
Mixing protocols propose supplementary services for existing blockchains (mostly for Bitcoin), and are specifically designed to be fully compatible (e.g., (Ruffing et al., 2016; Valenta and Rowan, 2015; Qin et al., 2023; Ruffing and Moreno-Sanchez, 2017)). Another category of compatible solutions leverages blockchain programmability to introduce privacy-preserving overlays (e.g., (Bünz et al., 2020; Baudet et al., 2023)). For instance, Zether (Bünz et al., 2020) uses a smart contract on Ethereum to store and manage private accounts; Zef (Baudet et al., 2023) introduces opaque coins in the form of off-chain certificates exchanged between Ethereum user accounts.
Non-compatible Extensions
Most privacy-enhancing proposals are non-compatible with the blockchain protocol they extend. For example, Zerocoin (Miers et al., 2013) is an overlay over Bitcoin that allows user to convert bitcoins in zerocoins. A zerocoin can be proved (in ZK) to originate from a valid and unspent bitcoin, without explicitly revealing such link. Zerocoin does not provide full anonymity: it obscures coin traceability, but it does not hide the number of transactions, recipient addresses, and balances. Building on Zerocoin, Zerocash (Ben Sasson et al., 2014) enhances privacy by concealing both sender and recipient addresses, supporting arbitrary transaction values, and enabling change handling. Both Zerocoin and Zerocash require new transaction types and payment semantics, and break compatibility with Bitcoin. While, in theory, they could be integrated if a super-majority of miners and nodes adopt the new protocol, the authors do not expect such integration and instead recommend deploying Zerocash as a separate system. Garman et al. (Garman et al., 2017) extends Zerocash with no backward compatibility, by introducing policy enforcement and tracing of users and tainted coins—features that further diverge from Bitcoin’s existing design. Yu et al. (Yu et al., 2024) propose an overlay on Bitcoin for anonymous yet regulated transactions using a coin backed by bitcoins. Their approach relies on three centralized entities (i.e., bank, tracing authority, and technical support) that cooperate to enable batch linkability and payer tracing while preserving privacy. These changes are not backward-compatible with Bitcoin.
Novel Approaches
Most approaches present greenfield solutions for privacy-preserving payments or asset exchanges (e.g., (Van Saberhagen, 2013; Poelstra, 2016; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Wüst et al., 2019; Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020)), often extending open-source solutions (e.g., (Li et al., 2021; Papadoulis et al., 2024)). Both centralized (e.g., (Cecchetti et al., 2017; Zhang et al., 2019)) and decentralized architectures (e.g., (Narula et al., 2018; Androulaki et al., 2020; Lin et al., 2020; Ma et al., 2021; Goodell et al., 2023; Elfadul et al., 2024)) exist, with a sensible preference for the latter. Early approaches like CryptoNote (Van Saberhagen, 2013), MimbleWimble (Poelstra, 2016), and Monero (Noether et al., 2016) address Bitcoin’s privacy limitations by still proposing fully decentralized alternatives. CryptoNote (Van Saberhagen, 2013) enhances sender and receiver anonymity through one-time ring signatures and NIZK proofs, inspiring solutions like ByteCoin, DigitalNote, and Aeon (Amarasinghe et al., 2019). MimbleWimble (Poelstra, 2016) improves privacy and scalability by aggregating confidential transactions. However, this approach does not guarantee unlinkability among participants who undertake the aggregation. Grin [^12] https://grin.mw/ and Beam [^13] https://www.beam.mw/ build on MimbleWimble ideas. Monero (Noether et al., 2016) strengthens privacy with sender, receiver, and network anonymity as well as confidential transactions. Despite its strong anonymity measures, research has identified vulnerabilities in Monero’s mixing selection strategy that could lead to transaction de-anonymization (e.g., (Möser et al., 2018; Kappos et al., 2018)).
Other approaches propose partially decentralized solutions (e.g., (Wüst et al., 2019)), combine concepts from centralized mixers (e.g., (Fauzi et al., 2019)), or consider hierarchical bank-intermediated settings (e.g., (Cecchetti et al., 2017; Narula et al., 2018)). Wüst et al. (Wüst et al., 2019) propose PRCash to strike a trade-off between fast payments, user privacy, and regulatory oversight: it includes two centralized entities for issuance and oversight and a permissioned blockchain, with MimbleWimble-like transactions (i.e., commitment-based; see Section 4.2.2). Quisquis (Fauzi et al., 2019) builds on the concept of mixing to redistribute wealth: each transaction updates a bunch of other accounts on chain, thus preventing addresses from appearing multiple times on the ledger, reducing the risk of de-anonymization and linkability. Quisquis ensures integrity through ZKPs while maintaining privacy using commitments and updatable public keys.
4.1.4. Trust Assumptions
We now focus on the degree of trust required to facilitate secure transactions. Most solutions for digital payments do not rely on trusted parties for confidentiality (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Noether et al., 2016; Narula et al., 2018; Fauzi et al., 2019; Li et al., 2021; Ma et al., 2021; Guan et al., 2022; Tomescu et al., 2022; Baudet et al., 2023; Nosouhi et al., 2023; Sarencheh et al., 2025; Wang et al., 2023)), even when they propose a centralized system (e.g., (Chaum, 1983; Chaum et al., 2021; Wüst et al., 2022)).
The solutions that introduce trust assumptions do it for different purposes. The approaches using zk-SNARK as ZKP require a trusted setup to initialize and correctly share some cryptographic parameters (e.g., (Miers et al., 2013; Garman et al., 2017; Ben Sasson et al., 2014; Xue et al., 2023)). A few works require a trusted central bank to issue coins that can be later exchanged anonymously (e.g., (Goodell et al., 2023; Friolo et al., 2026)). Elfadul et al. (Elfadul et al., 2024) assume the presence of a regulatory supervision that grants permissions to users to perform transaction; this centralized authority could asymmetrically revoke such authorization. Contributions that explore special forms of auditability related to revocable anonymity (e.g. (Wüst et al., 2019; Lin et al., 2020; Xue et al., 2023; Yu et al., 2024)), trapdoor (e.g., (Zhang et al., 2019; Androulaki et al., 2020; Chen et al., 2020; Elfadul et al., 2024)), or user tracing (e.g., (Garman et al., 2017)) require trust in the tracing authority that uncovers transaction details only when necessary. PEReDi (Kiayias et al., 2022) reduces trust requirements by relying on quorum of centralized entities that must agree to perform specific operations (such as decrypting transactions).
4.1.5. Interaction Models
Privacy-preserving digital payment systems differ in whether transactions require direct interaction between payer and payee. Interactive payment protocols need active cooperation during the transaction process. This is common in account-based models where recipients help in constructing or approving the updated account state, via ZKPs or digital signatures. For example, PEReDi (Kiayias et al., 2022) and Platypus (Wüst et al., 2022) enforce synchronized interaction: both parties participate in updating the ledger with proofs and commitments, ensuring mutual accountability and preventing attacks like dusting or replay. These systems may offer enhanced control and auditability but at the cost of real-time coordination and increased communication overhead. Similarly, Chaumian token systems often require immediate token claiming to prevent double spending (e.g., (Dold, 2019)). In contrast, non-interactive payment protocols let the payer unilaterally create and broadcast transactions to the network. Many token-based systems, especially unspent token models, follow this approach (e.g., (Ben Sasson et al., 2014; Van Saberhagen, 2013; Noether et al., 2016; Poelstra, 2016)). This improves user convenience and enables asynchronous payments, although receivers may need periodic ledger scanning or local computation. [^14] Off-chain communication can alleviate the burden of scanning the ledger, but current protocols lack this feature. Some account-based protocols approximate non-interactive behavior via delayed claiming. For instance, Sarencheh et al. (Sarencheh et al., 2025) propose a two stage and non-interactive variant of PEReDi: first, the sender interacts with issuers to update his account (transfer stage); then, the receiver claims funds using enclosed ZKP and credits his account (claim stage). Overall, interactive designs enable rich policy enforcement and conditional privacy, whereas non-interactive models better suit high-latency or asynchronous networks.
4.2. Value Representation
The value representation dimension examines how balances and transfers are modeled, focusing in particular on their privacy implications.
4.2.1. Ledger Data Models
Digital payment systems and decentralized ledgers follow three main approaches for representing and tracking ownership of digital assets: accounts, tokens, and tabular ledger.
Account-based
Account-based ledgers associate a balance with each account address, which is debited or credited by transactions. This simplifies state management and enables simple definitions of policies based on cumulative balances. Nonetheless, it requires safeguards against replay and double-spending attacks, often implemented by storing a sequence number (or nonce) in each account and updating it with every transaction. Since the balance of an account is explicitly represented, achieving (pseudo-)anonymity and unlinkability is challenging, especially in systems where users cannot freely generate multiple addresses (e.g., (Guan et al., 2022)). Privacy-preserving account-based solutions replace plain-text balances with commitments, achieving confidentiality of balances and transactions (e.g., (Fauzi et al., 2019; Bünz et al., 2020; Baudet et al., 2023; Guan et al., 2022; Sarencheh et al., 2025)) [^15] While in the first generation of systems accounts are kept centralized and private, second and third generation systems are often public, enabling inspection and verification.. Liu et al. (Liu et al., 2022) propose an account-based CBDC with sender and receiver anonymity but with no unlinkability: The system follows a two-tier model where the central bank can identify users via encrypted identifiers, while commercial banks cannot: users use blind signatures by the central bank for account generation and withdrawals and prove validity of their operations using ZKPs. To guarantee anonymity while ensuring unlinkability, most of the approaches change users’ account address at each transaction and use an authenticator to prove correctness. The authenticator can be implemented, e.g., using digital signatures by trusted or centralized authorities (e.g., (Kiayias et al., 2022)) or ZKPs (e.g., (Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)). Platypus (Wüst et al., 2022) rely on ZKPs; conversely, in PEReDi (Kiayias et al., 2022), account states must be (blindly) signed by a quorum of maintainers, and since each transaction reveals prior account snapshots to construct new authenticators, randomized signatures are used to prevent linkability. Quisquis (Fauzi et al., 2019) and AQQUA (Papadoulis et al., 2024) use updatable public keys to update accounts in a decentralized manner. The account-based model lends itself well to auditability implemented in third generation systems, as it allows simple definition of policies on holding limits (e.g., (Gross et al., 2021; Wüst et al., 2022; Papadoulis et al., 2024)) or spending and receiving limits (e.g., (Papadoulis et al., 2024)). Gross et al. (Gross et al., 2021) includes, in the account state, an accumulator of all spending in the current epoch, enabling policies on this information. AQQUA (Papadoulis et al., 2024) includes commitments to three different counters (i.e., balance, total amounts sent and received). Conversely, Platypus (Wüst et al., 2022) leaves the specification of the additional information to be application dependent.
Token-based
Token-based ledgers do not maintain explicit balances per address, but instead track individual tokens. The user’s balance can be derived from the sum of all owned tokens. This model enables parallel validation of user transactions and stronger privacy guarantees (as balances are not explicitly materialized) but introduces management complexity, particularly when users distribute their coins across multiple addresses for added privacy. There are two main types of token structures: Chaumian tokens and unspent tokens.
Chaumian tokens (Chaum, 1983) use blind signatures and were first deployed in centralized, bank-issued electronic cash schemes (i.e., first generation systems). When issuing a token, the bank blindly signs a random value provided by the payer, allowing the payer to later unblind it for spending (ensuring sender anonymity). These tokens rely on a trusted authority to issue and verify their validity; to this end, the authority must hold an ever increasing list of spent tokens that might introduce scalability challenges. Some implementations introduce expiration to limit circulation (e.g., (Dold, 2019; Lovejoy et al., 2023)). Chaumian-inspired decentralized privacy-preserving systems of the second generation (e.g., Monero (Noether et al., 2016), Zerocoin (Miers et al., 2013), and Zerocash (Ben Sasson et al., 2014)) eliminate the need of a central issuer using a combination of different cryptographic techniques. For example, Zerocash supports transparent (like Bitcoin) and shielded transactions, with the latter hiding sender, recipient, and amount. Shielded payments encrypt these values with one-time keys and use zk-SNARK proofs to validate inputs-outputs equality, concealing all on-chain details.
Unspent tokens, introduced in Bitcoin (Nakamoto, 2008), represent transaction outputs that can be used as inputs for future transactions, which in turn create new unspent tokens. As such, they are inherently traceable unless PETs are used (Baldimtsi et al., 2024). Unspent tokens are issued in a decentralized manner, usually as result of the mining process. Their core attributes include an identifier (e.g., transaction hash and output index), value (or its commitment), and ownership information. Their recording in a public, transparent ledger prevents double-spending but complicates privacy. Privacy-focused crypto-asset systems employ stealth addresses, ring signatures, confidential transactions, and ZKPs to obscure transaction details and unlink payments (e.g., (Noether et al., 2016; Poelstra, 2016; Miers et al., 2013; Ben Sasson et al., 2014))—these properties are discussed in more detail later. Tokens can also include auxiliary data to support the evaluation of specific policies (e.g., (Xue et al., 2023; Yu et al., 2024)) for third generation systems. For example, Xue et al. (Xue et al., 2023) include data to prove in ZK that the total value transferred and the transaction frequency in a time period is limited. Section 4.4 further discusses the auditability of token-based systems.
Tabular Ledger
Narula et al. (Narula et al., 2018) propose zkLedger, a third approach tailored to a specific use case: a bank-intermediated CBDC. It uses a tabular ledger where rows represent transactions and columns represent banks. Each cell contains a Pedersen commitment to hide values, while enabling cryptographic verification. Transactions include ZKPs to prove balance, asset ownership, and consistency with audit records. This structure supports rich privacy-preserving auditing: this third generation system includes an auditor who can issue a rich set of auditing queries—using primitives like sums, moving averages, variance, standard deviation, ratios—and receive answers that are provably consistent with the ledger. As a downside, the tabular structure scales poorly with the number of banks and transactions, since the ledger grows monotonically. MiniLedger (Chatzigiannis and Baldimtsi, 2021) adopts a similar tabular abstraction but addresses scalability by replacing commitments with (twisted ElGamal) encryptions and introducing cryptographically verifiable ledger pruning. MiniLedger allows old transaction rows to be compacted into digests without losing auditability: banks can later selectively disclose individual transactions, aggregates, or balances together with ZKPs, even after pruning. This makes MiniLedger better suited for long-running inter-bank settlement scenarios.
4.2.2. Ledger State Update
Updating the ledger may reveal sensitive information when it exposes which state elements are created, updated, or consumed to non-involved parties (e.g., globally on public ledgers or to intermediaries), enabling linkability through update patterns (even if values are hidden). Privacy-preserving systems rely on commitments, cryptographic accumulators (e.g., Merkle trees), or tags such as nullifiers, to validate state transitions and prevent double spending while minimizing information leakage.
Account-based
A generic state update transaction for account-based ledger includes the sender and receiver addresses, the exchanged amount, and the sender digital signature. For privacy-enhanced account-based, a naïve state update would reveal the updated accounts. Therefore, privacy-preserving state updates in account-based ledger usually adopt three strategies (e.g., (Kiayias et al., 2022; Wüst et al., 2022; Fauzi et al., 2019; Bünz et al., 2020; Papadoulis et al., 2024)): to obtain confidentiality, they update a commitment to the new account state; to obtain unlinkability, they always update the public key associated with the (recipient) account; and to obtain anonymity, they update a group of accounts at the same time.
Token-based
Token-based ledgers store the list of valid (unspent) tokens or burnt (spent) tokens. A spending operation must invalidate an existing unspent token, to prevent double-spending. With Chaumian tokens (e.g., (Chaum, 1983; Dold, 2019; Lovejoy et al., 2023)), payment systems include a withdrawal operation (or mint transaction), which creates a new valid token that the payer can spend later in time, and a spend operation (or pour transaction). The state of the system is represented by the list of spent tokens, which is updated at spending time; unspent tokens are usually not explicitly stored (if not by payers). Privacy-preserving systems based on Chaumian tokens do not usually guarantee transaction confidentiality (e.g., (Chaum, 1983; Hayes, 1990; Sander and Ta-Shma, 1999; Dold, 2019; Goodell et al., 2023; Friolo et al., 2026)), therefore they still maintain the list of spent tokens. Unlike other approaches, Sander and Ta-Shma (Sander and Ta-Shma, 1999) represent valid tokens in a Merkle tree. The bank does not blind sign tokens; instead, valid tokens come with a hash chain proving membership in the Merkle tree. The Merkle root is publicly available to ensure transparency and auditability. Withdrawal coins can be invalidated if needed, addressing certain attack scenarios (e.g., blackmail). To hide amounts, Monero wraps each amount in a Pedersen commitment, and proves via a Bulletproof range ZKP that each value lies in a valid range (no negative or overflow). The additive properties of Pedersen commitments allows proving input-output equality without revealing any value. Friolo et al. (Friolo et al., 2026) employ a ZKP-based protocol to achieve unlinkability: Transactions are represented as commitments, and when spent, their opening is revealed along with a ZKP demonstrating that it is linked to a valid, unspent token. In contrast, UTT (Tomescu et al., 2022) ensures transaction confidentiality, and it recurs to nullifiers to track spent coins. Nullifiers enable the bank to detect double-spending without knowing transaction specifics. In (Tomescu et al., 2022), a nullifier is a pseudo-random value computed from the coin’s serial number (and a secret). The system checks their validity through ZKPs and, if valid, stores them in a list for future checks.
With unspent tokens, the payment operation simultaneously invalidates previous tokens and generate new ones. The state of the system is represented by the list of unspent tokens [^16] Since blockchains store transactions in an append-only log, they trace the derivation history of each unspent token.. Unspent tokens cannot be easily marked as spent without compromising sender anonymity. Most of the approaches, including CryptoNote (Van Saberhagen, 2013), Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), and their extensions (e.g., (Garman et al., 2017; Androulaki and Karame, 2014; Xue et al., 2023)), let the payer spend a token by using a serial number and a ZKP that indicates that the number is a valid opening to some commitment stored in an unspent token. Monero (Noether et al., 2016) (and its extension Traceable Monero (Li et al., 2021)) uses one-time signatures for unspent tokens, which can be used only once to spend the token. At spending time, the ledger registers a salted hash of the public key. MimbleWimble (and PRCash that reuses its transaction format (Wüst et al., 2019)) uses a binding and hiding commitment scheme (Pedersen, 1992), together with Bulletproof range proofs (Bünz et al., 2018), to define the recipient. By using commitments as outputs, coins act as having their own private key, the knowledge of which, along with the knowledge of their value, enable spending the token. This guarantees anonymity and unlinkability.
4.2.3. Handling Change
Account-based ledgers use transactions to transfer exact value between payer and payee, therefore these kind of systems do not need to support change. Instead, token-based ledgers propose different approaches for handling change. Chaumian tokens do not easily support change or partial spending (e.g., (Dold, 2019)). Tomescu et al. (Tomescu et al., 2022) allow multiple inputs and outputs for a single transaction, but requires the payer to prove value preservation in ZK. The proof consists in two parts: the first proves that the sum of inputs equals the sum of outputs; the second is a range proof on the output value (needed because ZKPs works with modulo arithmetic and the equality proof is not enough to guarantee value preservation). System based on unspent tokens inherit the change implementation strategy by Bitcoin, where the change is another unspent token owned by a (new) address managed by the payer (e.g., (Miers et al., 2013; Androulaki et al., 2020; Elfadul et al., 2024)) or a commitment controlled by the payer (e.g., (Noether et al., 2016; Wüst et al., 2019)).
4.2.4. Support for Different Denominations
Chaum (Chaum, 1983) considered a simple system where tokens have implicit value. A few extensions consider the possibility to emit tokens with different denominations (e.g., (Dold, 2019; Tomescu et al., 2022)). For example, Dold (Dold, 2019) in GNU Taler uses multiple denomination key pairs for blind-signing coins of different monetary values. Conversely, Tomescu et al. (Tomescu et al., 2022) explicitly represent the token value by concealing it within a commitment. As regards unspent token-based ledgers, Zerocoin does not support arbitrary denominations (Miers et al., 2013). This design simplifies the ZKP used for anonymity but introduces usability limitations, as users must split or combine coins when making payments. Zerocash (Ben Sasson et al., 2014) supports arbitrary denominations by using homomorphic commitments, allowing users to transact any amount while preserving privacy. The remaining approaches allow tokens of different denominations, as each unspent token internally holds its value (or a commitment to it). Among them, CryptoNote (Van Saberhagen, 2013) does not hide transaction amounts, making it vulnerable to statistical analysis—e.g., identifying change as the smaller output. Using multiple change addresses can obscure amounts but leads to proliferation of dust transactions (Noether et al., 2016). Hiding values with commitments prevents such analysis.
4.2.5. Token Life-cycle
Double Spending Protection
Two approaches protect from double spending: deposit-time checks and publicly verifiable ledgers. Centralized systems with Chaumian tokens do not usually rely on a transparent ledger; instead, they check token validity at deposit time (e.g., (Chaum, 1983; Dold, 2019; Tomescu et al., 2022; Goodell et al., 2023)). The public visibility of the ledger in blockchains (e.g., (Nakamoto, 2008)) allows payee to autonomously check token validity. This is also possible in privacy preserving systems, which usually require the payer to provide a ZKP of token validity (e.g., (Miers et al., 2013; Poelstra, 2016; Noether et al., 2016; Li et al., 2021; Xue et al., 2023; Yu et al., 2024)).
Transferable Tokens
Few works (e.g.,(Hayes, 1990; Goodell et al., 2023)) propose mechanisms for transferring Chaumian tokens. Hayes (Hayes, 1990) proposed chaining transactions, where each owner signs over the previous link; long chains can be refreshed by the bank, though the scheme risks double spending, mitigated by token lifetimes. Goodell et al. (Goodell et al., 2023) instead attach a proof of provenance—a list of commitments with the issuer’s signature—allowing validity checks without external interaction.
Immediate Claiming
With Chaumian tokens, transactions occur off-ledger, so payees must deposit tokens promptly to prevent double spending (Chaum, 1983). Misbehavior can be deterred by one-time signatures that reveal the sender’s identity (Hayes, 1990; Camenisch et al., 2006). Sander and Ta-Shma (Sander and Ta-Shma, 1999) instead use Merkle trees: coin validity is proven via inclusion proofs to public roots, enabling offline asynchronous payments without real-time bank interaction.
4.3. Privacy Features
Privacy in digital payment systems is often defined through several sub-properties, including participant anonymity, value confidentiality, plausible deniability, and unlinkability (e.g., (Almashaqbeh and Solomon, 2022; Amarasinghe et al., 2019; Bernal Bernabe et al., 2019; Genkin et al., 2018)). The privacy mechanisms discussed in this section are largely shared across generations; differences arise primarily in how these mechanisms are combined and integrated with system-level objectives, with techniques such as encryption, ZKPs, and credentials [^17] By credentials we refer to cryptographic constructions—typically based on digital signatures combined with ZKPs—that bind attributes or rights to a user and enable selective disclosure without revealing unnecessary information. enabling selective disclosure and auditability.
4.3.1. Sender Anonymity
Asymmetric anonymity indicates that only one of the participants’ identities (sender or receiver) is protected (e.g., (Chaum, 1983; Sander and Ta-Shma, 1999; Miers et al., 2013; Dold, 2019; Zhang et al., 2019; BIS Innovation Hub, 2023; Goodell et al., 2023; Yu et al., 2024; Androulaki and Karame, 2014)). Sender anonymity focuses only on hiding the payer’s identity, usually using the following techniques: blind signatures (e.g., (Chaum, 1983; Dold, 2019; Goodell et al., 2023; Liu et al., 2022)) and their variants—one-time (e.g., (Hayes, 1990)), threshold (e.g., (Sarencheh et al., 2025; Kiayias et al., 2022; Androulaki et al., 2020)), and randomizable (e.g., (Tomescu et al., 2022; Wüst et al., 2019))—, ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024)), group signatures (e.g., (Zhang et al., 2019)), and commitments with ZKPs (e.g., (Miers et al., 2013; Ben Sasson et al., 2014)).
Chaum (Chaum, 1983) proposes blind signatures to unlink the sender request of token issuance and its later spending. To reveal the sender on double spending, Hayes (Hayes, 1990) introduces one-time blind signatures. However, since each transaction requires a new signing key, storage and computation management become cumbersome; moreover, one-time signatures often require large signature size and expensive verification, making them impractical for high-throughput systems. UTT (Tomescu et al., 2022) implements blind signatures using randomizable signatures (Pointcheval and Sanders, 2016), which allow signatures to be re-randomized using fresh randomness without compromising their validity. In particular, each coin consists of a commitment signed by the central bank, where the commitment binds the payer’s public identifier, the coin value, and a random nonce (to ensure hiding). All these schemes rely on a central issuer, which represents a potential single point of failure. This can be mitigated, e.g., by using threshold blind signatures, where a quorum of authorities must cooperate to create a valid signature (e.g., (Androulaki et al., 2020; Kiayias et al., 2022; Sarencheh et al., 2022)).
Other works reuse the idea of mixing through ring signatures, which conceals the sender’s public key within a group of other public keys—the anonymity set (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Li et al., 2021; Elfadul et al., 2024)). For example, CryptoNote (Van Saberhagen, 2013) employs one-time ring signatures: each transaction uses a fresh key and produces a key image, enabling detection of double spends. Different privacy-preserving crypto-asset systems use ring signatures for obtaining sender anonymity, including Monero [^18] Möser et al. (Möser et al., 2018) showed weaknesses in Monero’s key sampling: $66.09\%$ of transactions do not include any other public key, sampling is not really random, and real inputs can be inferred with $80\%$ accuracy. (Noether et al., 2016), Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024). However, studies show that Monero’s anonymity can be compromised through transaction analysis (e.g., (Möser et al., 2018; Kappos et al., 2018; Kumar et al., 2017)).
To revoke sender anonymity if needed, Zhang et al. (Zhang et al., 2019) use group signatures where a trusted centralized entity sets up the anonymity set and can, optionally, uncover the senders’ identity for inspection.
To be valid, unspent tokens require proving their ownership. To implement sender anonymity, existing approaches rely on two key techniques. First, the unspent token includes a commitment to its serial number; the serial number is revealed only to mark the token as spent (see Section 4.2.5). Second, the spending operation includes a ZKP that demonstrates knowledge of the opening to a commitment of an unspent coin. Sander and Ta-Shma (Sander and Ta-Shma, 1999) use multiple Merkle trees to track valid Chaumian tokens. To prove validity, the sender provides a ZKP attesting knowledge of a valid coin in the tree and of the coin’s serial number and a hash chain leading to one of Merkle tree roots. Zerocoin (Miers et al., 2013), and its extensions (e.g., (Ben Sasson et al., 2014; Garman et al., 2017)), adopts ZKPs as well: the payer publishes a transaction that includes the recipient’s address, an empty origin address, a ZKP, and the commitment opening. The ZKP prevents linking the sender to any specific commitment; the commitment itself hides the opening using a random value that is never revealed, preventing any association between the two.
Privacy-preserving account-based models mostly rely on ZKPs to obtain sender anonymity (e.g., (Gross et al., 2021; Kiayias et al., 2022; Wüst et al., 2022; Beer et al., 2024)). Specifically, a commitment to the account state is used, with updates including ZKPs to prove correctness. For example, Gross et al. (Gross et al., 2021) use the ideas of Zerocash by committing on the accounts state and by invalidating them on update. The authors suggest using Merkle trees to prove that a transaction proposal refers to an existing commitment in the ledger without pointing to (and thus revealing) it. [^19] For a broader discussion of set accumulators—including Merkle trees as a key example—and their role in enhancing anonymity, see (Loporchio et al., 2023). Platypus (Wüst et al., 2022) uses a model where payer and payee collaborate to exchange with the central bank commitments to their updated accounts, the transaction (amount and serial number), and two ZKPs proving the update’s validity and its link to a prior certified state. If valid the bank updates a public ledger. Conversely, PEReDi (Kiayias et al., 2022) achieves sender anonymity through anonymous channels, an encrypted ledger, and a threshold signature scheme. Sender and receiver submit partial (specular) transactions that update only their respective account states, encrypted using threshold encryption. Solidus (Cecchetti et al., 2017) is a hierarchical system where banks act as proxies, managing on-chain transactions on behalf of users and ensuring transaction-graph confidentiality. [^20] While banks can identify their respective clients, other entities only learn the identities of the banks.
4.3.2. Receiver Anonymity
The key techniques for receiver anonymity are one-time destination keys or stealth address (e.g., (Van Saberhagen, 2013; Noether et al., 2016; Elfadul et al., 2024; Zhang et al., 2024)), encrypted address (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)), commitments as output (e.g., (Poelstra, 2016; Wüst et al., 2019)), and updatable public key (e.g., (Fauzi et al., 2019; Papadoulis et al., 2024)). Obtaining receiver anonymity with Chaumian tokens is challenging, as the payee must interact with the bank to deposit the received tokens. Hayes (Hayes, 1990) suggests using one-time aliases to obtain privacy, even though the idea is not detailed for practical usages. A similar idea can be found in CryptoNote (Van Saberhagen, 2013), where the sender generates a one-time destination key by combining a random number with the receiver’s public key—essentially performing a Diffie-Hellman exchange to derive a shared secret. [^21] Because this process requires two elliptic curve keys, a CryptoNote address is twice the size of a standard elliptic curve public key. The sender posts the transaction to a one-time public key with a commitment to a random value. The receiver scans all transactions with his private key, reconstructs the destination key, and accepts the payment if it matches. Sender and receiver never interact directly, but each receiver must scan the entire ledger to find his payments. This approach has been later integrated in other systems, including Monero (Noether et al., 2016), which calls the one-time receiver address as stealth address, Traceable Monero (Li et al., 2021), and PCRAP (Elfadul et al., 2024) (see Section 3.6).
In Zerocash (Ben Sasson et al., 2014) (unspent token-based ledger), the sender includes an encrypted version of the recipient’s address in the transaction (using asymmetric encryption). Therefore, periodically, each user has to scan the ledger, decrypt transactions using his private key, and identify those intended for him. This prevents external observers from linking payments to specific recipients while allowing recipients to identify incoming transactions without revealing their addresses publicly. UTT (Tomescu et al., 2022) (Chaumian token-based system) ensures receiver anonymity using homomorphic commitments and anonymous IBE scheme (Boneh and Franklin, 2003). A coin is a commitment to a tuple containing the owner’s identifier, a serial number issued by the central bank, and its value. To be valid, each coin comes with a (randomizable) signature by the central bank. To make a payment, the sender creates a transaction for the payee, which includes the receiver’s identity commitment, the output value commitment, the encrypted coin details (with the recipient’s key), and ZKPs. The transaction is sent to the bank for validation; if valid, the bank blindly issues new coins for the recipient and records them on a public ledger [^22] The homomorphic properties of the commitment scheme ensures that the bank does not learn identities or amounts, and allows to directly combine the value commitment, identity commitment, and serial number commitment in a single coin commitment.. Recipients periodically scan the ledger to get their coins, to be randomized before adding to the wallet.
The unspent token-based ledger MimbleWimble (Poelstra, 2016) (and PRCash (Wüst et al., 2019)) performs transactions between commitments rather than addresses; this guarantees both receiver anonymity and unlinkability.
Account-based approaches typically achieve receiver anonymity using the same technique as for sender anonymity (e.g., (Wüst et al., 2022; Kiayias et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Lin et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Sarencheh et al., 2025; Papadoulis et al., 2024; Fauzi et al., 2019)). This is done by updating account state commitments and employing ZKPs to verify correctness. We mention Quisquis (Fauzi et al., 2019), where an account is a pair of public key and a homomorphic commitment to its balance. A transaction updates both the public key and the commitment using updatable public key (see Section 3.6). Senders select public keys—including the recipient’s and decoys—to form the transaction input, updating only the sender’s and recipient’s balances while leaving others unchanged. Since each address appears at most twice on the ledger, de-anonymization is harder. However, this approach requires users to periodically scan the ledger to update their keys, thus introducing some transaction serialization.
4.3.3. Transaction Value Confidentiality
Two main approaches conceal transaction details: commitments (e.g., (Ben Sasson et al., 2014; Androulaki and Karame, 2014; Noether et al., 2016; Poelstra, 2016; Garman et al., 2017; Fauzi et al., 2019; Wüst et al., 2019; Gross et al., 2021; Guan et al., 2022; Wüst et al., 2022; Beer et al., 2024; Elfadul et al., 2024; Tomescu et al., 2022; Androulaki et al., 2020; Narula et al., 2018)) and public key encryption (e.g., (Kiayias et al., 2022; Sarencheh et al., 2025; Bünz et al., 2020; Ma et al., 2021)).
Commitments
Camenisch et al. (Camenisch et al., 2006) use commitments with verifiable encryption to hide the commitment’s opening. In 2013, Maxwell proposed Confidential Transactions [^23] https://elementsproject.org/features/confidential-transactions/investigation, which use Pedersen commitments to hide transaction amounts while ensuring correctness via range proofs (e.g., Bulletproofs). Monero combines this technique with ring signatures obtaining RingCT (Noether et al., 2016). Zerocash (Ben Sasson et al., 2014) uses ZKPs to achieve full transaction privacy: a cryptographic commitment is made to a new coin, including its value, owner, and serial number; then, a ZKP ensures transaction validity without exposing sensitive data. Commitment-based schemes are the most widely adopted to obtain confidential transaction value, especially because their homomorphic versions allow to combine commitments without leaking sensitive data. To prevent users from creating negative or infinite amounts, range proofs prove that a committed value is greater than zero and within a valid range.
Public Key Encryption
In PEReDi (Kiayias et al., 2022), senders and receivers update their accounts by submitting encrypted transactions to several maintainers who operate a distributed ledger. Threshold ElGamal encryption (Elgamal, 1985) requires a quorum of maintainers to cooperate to decrypt transaction details. Authorities can trace both transactions and identities. PARScoin (Sarencheh et al., 2025) extends this approach to introduce an offline payment protocol that does not require interaction between sender and receiver. In Solidus (Cecchetti et al., 2017), Publicly Verifiable Oblivious RAMs (PVORMs) enable banks to update the accounts of their clients without revealing exactly which accounts are being updated. PVORM provides a ZKP demonstrating that the updates are correct with respect to the transaction triggering them.
A few solutions implement privacy-preserving digital payments using smart contracts on public blockchains. To ensure confidentiality, transactions are encrypted with the sender’s public key and supplemented with a NIZK proof for public verifiability (e.g., (Ma et al., 2021; Bünz et al., 2020)). For example, Zether (Bünz et al., 2020) uses smart contracts to define privacy-preserving tokens that can be exchanged between public keys associated with a confidential balance. They use ElGamal encryption, which has homomorphic properties, to hide each account’s balance. Zether introduces the $\Sigma$-Bullets (an enhancement of the Bulletproofs range ZKP) to efficiently prove statements over the encrypted transfer value and the new sender balance. However, Zether currently allows only one anonymous transfer per epoch (to prevent double spending), and execution costs (namely, gas costs on Ethereum) are very high (Almashaqbeh and Solomon, 2022). Similarly, DSC (Ma et al., 2021) optimizes privacy within Ethereum-style smart contract environments, but does not support anonymity: it uses ZKPs and homomorphic encryption to hide balances and transferred amounts. User balances are stored on the ledger in encrypted form. The homomorphic properties allows validators to update the balance in ciphertext without the need of decryption.
4.3.4. Network Anonymity
Network anonymity concerns the protection of sender and receiver IP addresses, which could otherwise reveal real-world identities. Most of the existing approaches do not investigate (or even mention) the network anonymity feature. Few works mention it but consider this issue as out of scope of their investigation (e.g., (Cecchetti et al., 2017; Baudet et al., 2023; Tomescu et al., 2022; Wüst et al., 2022)). They consider this feature as orthogonal hence that can be achieved by implementing the digital payment system on top of a private overlay network. Monero (Noether et al., 2016) and Traceable Monero (Li et al., 2021) explicitly support network anonymity by using TOR (The Onion Router [^24] https://www.torproject.org/) and I2P (Invisible Internet Project [^25] https://geti2p.net/en/) to hide the real IP address (Shi et al., 2024). While TOR routes traffic through multiple relays to obscure the user’s IP address, I2P is a network optimized for anonymous peer-to-peer communication, using fully decentralized routing and encryption to enable different anonymous services. Moreover, Monero uses Dandelion++ (Fanti et al., 2018), an improvement of Dandelion (Bojja Venkatakrishnan et al., 2017), which prevents attackers from linking transactions to their source IP addresses. Dandelion++ uses randomization mechanisms for message forwarding, which happens in two phases: stem (anonymization) and fluff (broadcast). In the stem phase, a transaction is first relayed privately across a small, randomly chosen subset of nodes; each node probabilistically decides whether to continue forwarding the transaction along the stem or to switch to the second phase. In the fluff phase, the node broadcasts to the entire network using traditional flood propagation protocol.
4.3.5. Unlinkability
Unlinkability ensures that transactions cannot be correlated or traced back to the same user (other definitions in (Chaum, 1983; Amarasinghe et al., 2019; Zhang, 2023)). It can refer to the sender, receiver, or both. Obtaining unlinkability in public ledgers is challenging as links between addresses in crypto-asset systems stems from different patterns (e.g., (Ron and Shamir, 2013; Chan and Olmsted, 2017; Béres et al., 2021; Victor, 2020; Zhong and Mueen, 2024; Lu et al., 2022; Ermilov et al., 2017; Kalodner et al., 2020)); a non-exhaustive list follows.
  • Common input ownership heuristic: If multiple addresses appear as inputs in a single transaction, they are likely controlled by the same entity;
  • Change address: In a transaction with two outputs, one is the payment, the other one may be the change returned to the sender;
  • Address reuse: Using the same address for multiple transactions makes it easy to link payments to the same user [^26] This works because address are pseudonymous but not private.;
  • Temporal analysis: If transactions occur in rapid succession or at predictable intervals, they may be linked to the same entity, e.g., due to automated systems that process transactions in batches or periodically;
  • Clustering analysis: It deals with identifying wallet behavior such as round numbers, specific fee patterns, or common spending habits;
  • Dust attack analysis: Small amounts of crypto-assets are sent to many addresses to track their movements (similar to a phishing attack); if dust is later spent with other transactions, it links addresses together.
Existing countermeasures primarily aim to obscure input-output relations within transactions by using mixing protocols and preventing address reuse. Mixing protocols disrupt common-ownership heuristics and help mitigate change-address attacks (an analysis of mixing protocols in  Appendix A.3.2). Address-changing mechanisms, such as generating a fresh sender or receiver address for each transaction, further reduce linkability. In addition, using network-level anonymity tools helps conceal user identities. Finally, the simplest defense against dust attacks is to ignore dust transactions, thereby preventing adversaries from linking addresses.
Integrated Solutions
Privacy-native solutions with public ledgers leverage a variety of strategies to improve unlinkability, focusing primarily on address-changing, transaction mixing, and use of ZKPs. Each of these techniques adds a layer of privacy, with trade-offs in terms of anonymity, computational complexity, and vulnerability to analysis. One-time destination addresses, stealth addresses (e.g., (Van Saberhagen, 2013; Noether et al., 2016)), or updatable public keys (e.g., (Papadoulis et al., 2024; Fauzi et al., 2019)) help obtain recipient unlinkability. While this ensures the recipient’s privacy, it does not hide transaction amounts. With ring signatures (e.g., (Noether et al., 2016; Van Saberhagen, 2013)), the privacy protection is limited by the number of decoys in the ring, and sophisticated heuristics can sometimes reduce the anonymity set. For the most robust privacy, ZKPs offer a powerful solution (e.g., (Ben Sasson et al., 2014; Poelstra, 2016)). They allow transactions to prove their validity without revealing any details about the sender, receiver, or transaction amount. With this approach, transactions are recorded as cryptographic proofs (instead of sender-receiver addresses), ensuring that outputs cannot be linked to inputs. Among ZK-based approaches, we mention the commitments and nullifiers strategy (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)): a coin contains a commitment to its serial number, and spending the coin involves publishing a nullifier, i.e., a pseudo-random value derived from the coin serial number. This ensures unlinkability to the original coin, while marking it as spent. However, designing a system with commitments and nullifiers is more complex than ring signatures or stealth addresses, and is also more computationally demanding (see Section 4.5).
4.3.6. Plausible Deniability
Plausible deniability indicates whether users can deny intentional participation in a transaction, preventing others from proving their involvement (e.g., (Almashaqbeh and Solomon, 2022; Ziegeldorf et al., 2018)). Existing works are roughly evenly divided between those supporting this property (e.g., (Noether et al., 2016; Poelstra, 2016; Narula et al., 2018; Kiayias et al., 2022; Tomescu et al., 2022; Wüst et al., 2022; Goodell et al., 2023; Papadoulis et al., 2024)) and those not (e.g., (Maxwell, 2013; Van Saberhagen, 2013; Garman et al., 2017; Bünz et al., 2020; Gross et al., 2021; Baudet et al., 2023; Guan et al., 2022; Yu et al., 2024)). Basically, protocols “on-demand”, such as mixing protocols, subsume an explicit user intention in using the additional service to obtain a privacy feature (such as unlinkability). Likewise, overlay systems implementing anonymous crypto-asset systems, such as Zerocoin (Miers et al., 2013), Zerocash (Ben Sasson et al., 2014), EZC (Androulaki and Karame, 2014), Zether (Bünz et al., 2020), require explicit intention to convert a native asset to a privacy-preserving one. Conversely, systems that force anonymity or unlinkability as default, readily provide the property of plausible deniability. Among these systems, we mention Monero (Noether et al., 2016), MimbleWimble (Poelstra, 2016), UTT (Tomescu et al., 2022), PEReDi (Kiayias et al., 2022), or Platypus (Wüst et al., 2022), that do not provide the functionality of combining anonymous and public transactions.
4.4. Auditability
While privacy is a core requirement of digital payment systems, complete opacity is often incompatible with regulatory and operational constraints. In regulated environments (e.g., CBDCs), some form of oversight is required to support AML/CFT controls, institutional accountability, and post-incident analysis (e.g., fraud, theft, protocol failures). Auditability should not be conflated with full transparency or continuous surveillance. Prior work has highlighted a broad design space between cash-like anonymity and full transaction observability, in which selective oversight can be enabled without pervasive disclosure (e.g., (Goodell and Aste, 2019; Pocher and Veneris, 2022; Ballaschk and Paulick, 2021; Michalopoulos et al., 2025)). In this sense, auditability enables authorized parties to verify compliance properties or selectively access transaction information, rather than making transactions universally visible. From a design perspective, auditability introduces a controlled relaxation of privacy guarantees. Increasing audit power typically strengthens trust assumptions or governance requirements and may reduce anonymity. Therefore, existing systems differ in what can be audited, when auditing may occur, and who is trusted to perform it. Following a regulation-by-design approach, auditability mechanisms can be embedded directly into the protocol to reconcile compliance with privacy. As discussed in Sections 4.1 and 4.2, the feasibility and scope of auditability mechanisms depend strongly on protocol architecture and on the underlying value representation. As such, auditability represents an explicit design concern of third generation systems, rather than an isolated feature.
Auditability, Policy, and Accountability
Auditability refers to the ability to examine, track, and verify transactions and balances while preserving a certain degree of privacy. It is a key feature of the third generation of privacy-preserving digital payment systems and, in particular, CBDCs. Auditability operates between these notions by supporting post-facto verification: it allows authorized parties to assess compliance or investigate suspicious behavior after transactions have occurred, and only within a prescribed scope. Two related concepts clarify its role: policy and accountability. A policy specifies admissible behavior and is typically enforced ex-ante by preventing invalid transactions. Accountability concerns the ability to associate actions with responsible entities, often via identity-binding mechanisms established at registration time.
What is audited, in practice?
Usually, auditability does not grant auditors unrestricted access to transaction data. Instead, it provides structured access to specific protocol state. This typically includes persistent state associated with users or tokens (e.g., counters or commitments), encrypted auxiliary metadata, and ZKPs attesting that policy constraints were respected. Importantly, auditability is rarely tied to a single account or address. To prevent trivial circumvention (e.g., by generating fresh accounts), audit-relevant state is anchored to identity credentials, token histories, or cryptographically bound counters that persist across transactions. As a result, address rotation does not reset spending history or policy enforcement. Auditability mechanisms represent constraints on system evolution rather than direct inspection tools: they regulate how hidden state may evolve over time and enable authorized parties to verify compliance without pervasive disclosure.
Mechanisms
We identify the following approaches that enable or support auditability: anonymity budget, operating limits, revocable anonymity and user tracing, coin tracing, and verification without disclosure.
4.4.1. Anonymity Budget
An anonymity budget restricts how many anonymous payments a participant can perform, with each transaction consuming part of it (e.g., (Camenisch et al., 2006; Wüst et al., 2019; Tomescu et al., 2022; ECB, 2019)). Camenisch et al. (Camenisch et al., 2006) propose a bounded-anonymity business model, where a public limit defines how many coins a user may anonymously transfer to a merchant. PRCash (Wüst et al., 2019) instead caps the total anonymous funds a user can receive within a time window. The anonymity budget is represented as a non-spendable output, whose size and duration regulate the flow of anonymous funds. Once exhausted—or if the user chooses non-anonymous transactions—the payee attaches his identity, encrypted under the regulator’s key. UTT (Tomescu et al., 2022) applies the budget to spending, requiring either loss of privacy or auditor intervention once the allocation is exceeded. Auditors can refresh budgets or approve extra transactions after requesting more details. In all these systems, the budget is realized as a special token included in every anonymous yet accountable transaction.
4.4.2. Operating Limits
Operating limits indicate constraints on transactions and accounts, such as maximum value, user balance, total sent/received funds, or transaction count within a time window. Compliance is generally proven in zero-knowledge. Garman et al. (Garman et al., 2017) extend Zerocash with auditability features, including regulatory closure (i.e., restricting transactions to assets of the same type), spending limits, selective user inquiries, and coin tracing. A central idea is the per-transaction spending limit: transactions above the limit are invalid unless signed by an authority. Another approach uses a per-user counter incremented with each outgoing transaction. The system can then reject transactions beyond a threshold. Gross et al. (Gross et al., 2021) build on this with a two-layer account-based system for implementing a CBDC, where each account tracks epoch turnover (total spent per epoch), proven with ZKPs. Similarly, AQQUA (Papadoulis et al., 2024) maintains multiple counters per account (balance, sent, received), publicly stored as three commitments. A registration authority links users identities to their (possibly, multiple) public keys, while an audit authority check compliance with five policy types: spending/receiving limits, transaction value limits, inactivity over a time period, selective transaction disclosure. During an audit, the authority selects a user by his public key and requests the opening of two historical account state snapshots (for every user’s accounts). After audits, the user randomizes his accounts (and provides a ZKP of valid update), thereby restoring his privacy. Other systems also rely on counters. Xue et al. (Xue et al., 2023) use per-user counters stored in a Merkle tree managed by regulators, which are also in charge of user on-boarding. ZKPs [^27] Specifically, zk-SNARKs are used for arithmetic statements and Sigma Protocols for algebraic statements. on counters enable policies on total amount and frequency of fund transfers within a time period. When suspicious transactions are detected, the participants’ identities can be recovered by regulators. Chen et al. (Chen et al., 2020) define policies on transaction limits, rates, and value, also verified with ZKPs. Across these proposals, auditability is achieved by maintaining cumulative state (e.g., counters or commitments) whose evolution is constrained by policy.
4.4.3. Revocable Anonymity and User Tracing
User tracing mechanisms enable authorities to discover a user’s identity or link his transactions when required, implementing revocable or conditional anonymity models. [^28] Revocable anonymity refers to a model where users remain anonymous by default, but their identity can be revealed under certain conditions. User tracing refers to technical mechanisms enabling authorized entities to identify users or link their transactions when required.
Some schemes rely on special digital signatures that expose the sender on double spending (e.g., (Hayes, 1990; Zhang et al., 2019; Beer et al., 2024)). For example, Chaum-Fiat-Naor (Chaum et al., 1990) embed identifying information that is revealed upon double spending, while Schoenmakers (Schoenmakers, 1998) improves efficiency using “once-concealed, twice-revealed” proofs. Brands (Brands, 1994) replaces blind signatures with credential-based encodings to support selective disclosure, and Camenisch-Lysyanskaya (Camenisch and Lysyanskaya, 2001) signatures further generalize these ideas to multi-attribute tokens with flexible revocation policies.
Other approaches embed encrypted identifiers, keys, or metadata into tags, accessible only to tracing authorities (e.g., (Cecchetti et al., 2017; Garman et al., 2017; Zhang et al., 2019; Liu et al., 2022; Lin et al., 2020; Li et al., 2021; Kiayias et al., 2022)). This is often referred to as the trapdoor approach. Many systems also introduce a registration authority that links cryptographic keys to real-world identities (e.g., (Zhang et al., 2019; Liu et al., 2022)). For instance, Barki et al. (Barki and Gouget, 2020) adopt this model in a permissioned setting, embedding encrypted identity information into transactions and enabling a designated revocation authority to deanonymize users when required. Cecchetti et al. (Cecchetti et al., 2017) use oblivious RAM to manage transactions across banks, while Liu et al. (Liu et al., 2022) encrypt user identities under the central bank’s key. Zhang et al. (Zhang et al., 2019) rely on a supervision authority with decryption power. Li et al. (Li et al., 2019) propose an identity tag encrypted under a tracing authority’s key. Garman et al. (Garman et al., 2017) propose a tracing mechanism where each user receives a unique encryption key: if the user is being traced, the key is a randomized version of the authority’s key; otherwise, it is derived from a null key. Androulaki et al. (Androulaki et al., 2020) encrypt transaction data under the auditors of both sender and receiver, while Xue et al. (Xue et al., 2023) scope tracing by encoding user identities with epoch identifiers. The epoch identifier enables temporal scoping, which limits the impact of key compromising and can enable granular tracing (e.g., to investigate suspicious behavior in a particular time window). Other systems derive one-time keys from long-term keys, enabling tracing by recovering the latter. Lin et al. (Lin et al., 2020) let users perform transactions using anonymous keys derived from their public addresses. A manager can use his key and the anonymous key to recover the user’s public address. Traceable Monero (Li et al., 2021) tags payee one-time keys: an authority can decrypt the tag and perform backward and forward tracing (i.e., identify the accounts that sent/receive funds to/from the current one).
A common concern is the trust concentration in a single tracing authority. To address this, Yu et al. (Yu et al., 2024) split tracing between two authorities via encrypted “marks” requiring joint computation, while PEReDi (Kiayias et al., 2022) (and Sarencheh et al. (Sarencheh et al., 2025)) distribute decryption among multiple parties, requiring quorum approval to identify users and access transaction metadata. Similarly, PICTURE (Zhang et al., 2024) adopts a threshold-based regulation model, where unlinkable transactions embed encrypted trace records that can be opened only by a quorum of authorities to recover sender and receiver identities.
4.4.4. Coin Tracing
Garman et al. (Garman et al., 2017) enable tracing by marking specific tainted coins, where all derived outputs remain traceable. Each coin carries a fresh encryption key; tracing data for outputs is encrypted under both this key and the authority’s public key, preventing unauthorized tracing by users. To avoid cascade tracing and collusion, exchanges can (verifiably) remove traceability with a public dummy key. Keller et al. (Keller et al., 2021) study de-anonymization in CoinJoin and ring signatures through collaboration with witnesses—i.e., users who confirm their inputs were not part of a mix. As the number of witnesses grows, the anonymity set shrinks, enabling more effective deanonymization.
4.4.5. Verification without Disclosure
Verification without disclosure lets verifiers confirm policy compliance without accessing private data. These mechanisms rely on cryptographic proofs (usually NIZK) to attest that hidden data satisfies constraints (e.g., (Narula et al., 2018; Wüst et al., 2022; Cecchetti et al., 2017; Bünz et al., 2020; Ma et al., 2021)). The verifier learns only the truth of the statement. Besides the approaches in Operating limits, here we mention those supporting policies beyond basic balance or transfer correctness check (e.g., (Bünz et al., 2020; Miers et al., 2013; Fauzi et al., 2019; Ma et al., 2021; Baudet et al., 2023; Guan et al., 2022; Beer et al., 2024; Papadoulis et al., 2024; Friolo et al., 2026)). zkLedger (Narula et al., 2018) uses Pedersen commitments enabling aggregate computations. It employs generalized Schnorr-based NIZKs to prove the correctness of computations without revealing underlying data. zkLedger supports complex auditing via a map-reduce model: banks generate commitments over filtered and transformed data (e.g., checking for non-zero values or squaring for variance) along with NIZKs proofs of correctness, enabling auditors to compute metrics like means and variances. [^29] Recently, Yang et al. (Yang et al., 2025) extend the zkLedger lineage to permissioned enterprise settings with auditable confidential asset transfers. Platypus (Wüst et al., 2022) defines account states with opaque auxiliary data to support custom policy checks. Solidus (Cecchetti et al., 2017) uses PVORM and ZKPs to publicly verify a bank’s transaction log compliance. DSC (Ma et al., 2021) enforces compliance of encrypted transactions with range and authorization proofs, without revealing transaction details. These approaches primarily support compliance verification rather than attribution: they enable auditors to assess whether rules were followed, without necessarily identifying the transacting parties.
4.4.6. Discussion
Choosing an auditability mechanism depends on the application context and threat model. Lightweight approaches such as anonymity budgets and operating limits are suitable for retail payments where limiting large-scale abuse is sufficient, but they offer limited forensic capabilities and are vulnerable to Sybil strategies unless combined with identity binding. Revocable anonymity provides stronger deterrence by enabling authorized identity recovery, at the cost of increased trust assumptions and governance complexity. Coin-tracing mechanisms support post-incident analysis of fund flows and recovery of tainted assets, but may weaken unlinkability and expose transaction graph structure. Verification without disclosure enables compliance checks (e.g., balance or limit enforcement) without revealing identities or values, but typically does not support attribution. Consequently, no single approach dominates the design space: compliance-oriented systems favor revocable or traceable designs, privacy-first systems rely on limited or aggregate auditability, and hybrid architectures increasingly combine multiple techniques to balance privacy, trust, and enforcement.
4.5. Summary and Takeaways
This section highlights the evolving landscape of privacy goals in digital payment systems. Early approaches focused primarily on anonymity and unlinkability via address obfuscation and mixing, while recent designs emphasize confidentiality and auditability, enabled by advanced cryptographic tools such as ZKPs, homomorphic commitments, and threshold encryption.
Each privacy goal—anonymity, confidentiality, unlinkability, and auditability—relies on distinct mechanisms and entails trade-offs. For instance, unlinkability may hinder auditability, while encrypted transactions that preserve confidentiality can complicate oversight. No single protocol satisfies all privacy and compliance objectives; systems tend to prioritize different goals based on their threat models, trust assumptions, and intended deployment contexts. Table 2 summarizes how privacy goals are addressed in practice, mapping privacy properties to cryptographic primitives, system mechanisms, and representative protocols. This provides a reference for researchers and practitioners navigating the trade-offs in privacy-preserving payment system design.
Table 2. Summary of privacy goals, main enabling technologies, and representative protocols
Privacy Goal Key Cryptographic Tools System Mechanisms Example Protocols
Sender Anonymity Blind, ring, group, and randomizable signatures, commitments and ZKPs One-time addresses, anonymity set, anonymous addresses eCash, UTT, Zerocoin
Receiver Anonymity One-time public keys, encrypted addresses, commitments One-time addresses MimbleWimble, Quisquis, Zerocash
Confidentiality Commitments, encryption, ZKPs Confidential values and balances Zether, PEReDi, Monero
Network Anonymity TOR, I2P, Dandelion++ Obfuscate IP addresses and routing paths Monero
Unlinkability Mixing protocols, ring signatures, one-time addresses, commitments Fresh key generation, decoy input, output set hiding Platypus, BlockMaze, CoinShuffle++
Auditability Authenticated tokens, threshold encryption, anonymous keys, trapdoor, ZKPs Anonymity budget, operating limits, tracing hooks, revocable anonymity PEReDi, Solidus, zkLedger
Privacy-preserving payment systems typically combine multiple cryptographic tools, each with distinct trade-offs. Blind signatures are lightweight but centralized; ring signatures provide scalable anonymity at latency cost; threshold encryption and commitments support auditability but not anonymity; and ZKPs (e.g., zkSNARKs, zkSTARKs) offer strong guarantees, but they are computationally intensive, may require substantial RAM (e.g., (Almashaqbeh and Solomon, 2022; Liang et al., 2025)), and require careful engineering to achieve practical deployment. One of the well-known ZKP-based systems is zk-SNARK, which makes transactions compact for verification but needs a trusted setup that, if mishandled, could compromise the security of the system. [^30] In practice, parameters are generated via publicly auditable multi-party computation ceremonies, and security holds as long as at least one participant is honest and destroys the corresponding secret randomness. Importantly, as shown in Section 4.4, implementing auditability often requires careful design of ledger data structures and interaction protocols to avoid unintended information leaks while still enabling policy verification. This multidimensional design space highlights the need for a holistic, design-oriented perspective, where privacy-preserving digital payment systems combine complementary mechanisms to balance user privacy with institutional accountability.
5. Open Challenges and Research Directions
Despite significant advances, numerous open challenges remain at the intersection of business requirements, system design, and expectations of privacy and auditability. We identify a few interesting research directions.
5.1. Formal Definitions of Privacy and Auditability
Existing systems use ad-hoc definitions for privacy and auditability, and previous works (e.g., (Genkin et al., 2018; Almashaqbeh and Solomon, 2022)) highlight theoretical gaps in unified privacy and auditability models. Most protocols informally claim anonymity, unlinkability, or confidentiality, but lack formal, composable definitions covering graph obfuscation, timing metadata, value-hiding, selective disclosure, and conditional anonymity. Systems combining multiple primitives rarely have UC (Canetti, 2000) or IITM (Küsters et al., 2020) proofs, and auditability features (e.g., view keys, trapdoors) are seldom formalized.
5.2. Scalability and Efficiency
Many cryptographic protocols and PETs remain computationally demanding (e.g., homomorphic encryption, ZKPs, and MPC). Future research must focus on developing optimized, lightweight, and parallelizable privacy-preserving protocols that can operate efficiently in high-throughput environments. Promising directions include hardware acceleration (e.g., via TEEs) and the exploration of protocol variants that balance privacy guarantees with resource efficiency. ZKPs play a central role in many systems, yet general-purpose schemes (e.g., zk-SNARKs, zk-STARKs) still face challenges related to trusted setup, proof aggregation, and post-quantum security. For instance, zk-SNARKs typically require a trusted setup, introducing centralization risks and limiting composability. While transparent SNARKs (e.g., Plonky2313131https://github.com/0xPolygonZero/plonky2 and Spartan (Setty, 2019)) and STARKs mitigate these concerns, they often incur higher computational costs. Achieving efficient, setup-free ZKPs without compromising performance remains an open research problem. Additionally, support for efficiently generating proofs that can verify other proofs—known as recursive proof composition—is still limited, even though this capability is crucial for scalable applications like off-chain rollups. Developing ZKPs that are not only succinct and recursive, but also universally verifiable (i.e., verifiable by anyone without special access), remains a technically challenging and active area of research. No current ZKP scheme offers simultaneously strong post-quantum guarantees [^32] ZKPs based on bilinear pairings (e.g., Plonk, Groth) or Pedersen vector commitments (i.e., Bulletproof) assume at least the discrete logarithm problem, hence are subject to quantum threats. Other SNARKs (e.g., Ligero, Orion) and STARKs using Merkle tree and other protocols to build a polynomial commitment are post quantum resistant. and performance comparable to elliptic curve-based constructions.
5.3. Balancing Privacy with Regulatory Compliance
Total anonymity is incompatible with regulatory oversight, making auditable privacy and selective disclosure critical design goals. An open challenge is how to enable systems where honest users maintain strong privacy guarantees, while still allowing authorities to investigate illicit activities under well-defined conditions. While preliminary approaches exist, significant work remains to formalize disclosure policies, define the types of information subject to auditing, and clarify the roles and interactions of participating entities. Relevant research directions include ZKPs with selective opening, trapdoor, witness encryption, and anonymity budgeting—each enabling varying levels of conditional transparency. Less explored primitives, such as witness encryption, may enable advanced privacy-preserving policies (see Appendix A.4.5). Additionally, emerging requirements, such as time-bound anonymity, jurisdiction-aware confidentiality, and programmable disclosure, call for flexible and context-aware privacy controls. Designing PETs that can support such adaptive behaviors, while remaining robust against misuse, poses a complex challenge involving technical, legal, and usability considerations.
5.4. Adaptive Privacy and Role Rotation
Privacy-preserving systems rely on anonymity sets, yet managing cryptographically sound methods for dynamic and attack-resistant management remain lacking, especially in decentralized or evolving environments. Current schemes, such as fixed-size ring signatures or heuristic decoy selection, degrade as usage patterns shift and are vulnerable to poisoning attacks, where adversaries insert tainted outputs to weaken privacy. Beyond anonymity sets, adaptive privacy must address the evolving nature of system participants, especially in regulatory or issuance roles. In CBDCs and regulated payment systems, auditors, compliance entities, or issuers may change due to policy or organizational shifts. Supporting this evolution without compromising anonymity or confidentiality is difficult: most systems assume static trust anchors or fixed multi-signature sets. Cryptographic tools such as threshold cryptography, distributed key generation, or policy-based encryption could enable dynamic reconfiguration, but balancing adaptability with unlinkability, correctness, and forward secrecy remains unresolved.
5.5. Network-Level Anonymity and Cross-layer Privacy Leakage
Network-layer anonymity is typically outsourced to external systems (e.g., Tor, mix networks). However, native integration of network-level privacy mechanisms into privacy-preserving payment system protocols remains largely unresolved. This delegation creates a critical cross-layer vulnerability, where even perfectly private on-chain transactions can be de-anonymized through timing analysis, transaction ordering, volume inference, or network topology correlation. Existing protocols often assume independence between transaction content and communication patterns, which adversaries can exploit. For instance, gossip-based mempool propagation leaks significant metadata about the origin and timing of transactions, enabling traffic correlation attacks. Emerging research directions include the design of privacy-preserving mempools that incorporate indistinguishable broadcasting, cover traffic, oblivious relaying, or dandelion-style (Fanti et al., 2018) diffusion to mask transaction origins. These approaches aim to reduce the information available to global or local adversaries observing network traffic (e.g., (Diaz et al., 2021; Bebel and Ojha, 2022)). However, these solutions often introduce latency, bandwidth overhead, or synchronization issues, leading to difficult trade-offs between efficiency, liveness, and anonymity guarantees. Addressing network anonymity in digital payment systems requires cross-layer designs that align messaging protocols with the cryptographic guarantees of the ledger. This remains a highly active research area in adversarial environments.
6. Conclusion
Designing digital payment systems that balance user privacy with institutional auditability remains one of the most intricate and impactful challenges in modern cryptographic and systems engineering. Over the past decade, crypto-asset systems and CBDC prototypes have pushed innovation across cryptographic protocols, PETs, and system architectures. Yet, despite substantial progress, the design space remains fragmented, and no universally accepted blueprint has emerged.
This survey provides a structured and technically grounded overview of the evolving landscape of privacy-preserving digital payment systems. We systematize the literature across both decentralized and centrally managed designs, including crypto-asset systems and CBDC proposals, introducing a comprehensive taxonomy that connects privacy and auditability goals with underlying cryptographic primitives, protocol designs, and system architectures. In doing so, we bridge conceptual objectives (such as anonymity, confidentiality, and auditability) with concrete technical mechanisms and design models.
Our analysis traces the evolution of privacy-preserving digital payment systems across three generations, reflecting an increasing interest in addressing privacy-accountability trade-offs. We underscore that privacy is not a monolithic property but a multifaceted design goal shaped by system requirements and regulatory constraints. Finally, we outline several open research challenges that call for interdisciplinary collaboration across cryptography, distributed systems, economics, and policy. Key directions include the development of scalable and composable privacy primitives, architectures that balance auditability with confidentiality, and formal models that capture realistic adversarial and compliance scenarios. As digital payment technologies progress from academic prototypes to deployed systems, balancing strong privacy guarantees with regulatory compliance will require ongoing technical advances and interdisciplinary efforts.
Acknowledgements.
We thank Marco Benedetti, Claudia Biancotti, Giuseppe Galano, Sara Giammusso, Antonio Muci, and Fabiana Rossi for their insightful discussions in the early stages of this work and for their valuable feedback throughout its development.
[…]
Supplementary Material to:
A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability
MATTEO NARDELLI, FRANCESCO DE SCLAVIS, MICHELA IEZZI, Bank of Italy, Italy
Appendix A Appendix
This appendix provides supplementary material for A Hitchhiker’s Guide to Privacy-Preserving Digital Payment Systems: A Survey on Anonymity, Confidentiality, and Auditability, which complements the main analysis but is excluded from the core sections to preserve narrative flow. Specifically, it includes: (i) the research methodology used for the literature review (Section A.1); (ii) a breakdown of publication venues for the works reviewed in this survey (Section A.2), and (iii) an analysis of mixing protocols (Section A.3) that complements the analysis in Section 4, with particular emphasis on unlinkability techniques discussed in Section 4.3.5. These materials are presented as tables and offer additional context and technical detail for readers interested in the broader research landscape and protocol-level design trade-offs. Furthermore, in Section A.4, we briefly present interesting yet underexplored cryptographic tools that can be used to design privacy-preserving digital payment systems.
A.1. Selection and Classification of the Reviewed Contributions
To identify the relevant contributions we analyzed in the paper, we adopted the following approach:
  1. (1)
    We identified a core collection of works relying on direct authors’ knowledge of the research field;
  2. (2)
    We enlarged this collection through various iterations of backward snowballing (Wohlin, 2014) (i.e., using the references), which allowed us to identify related works not included in the first step;
  3. (3)
    Analyzing the selected papers, we identified the most frequent publication venues (i.e., journals or conferences with 2+ selected papers). Checking the list of papers published in these venues, we added further papers to the collection;
  4. (4)
    As a final step, we performed keyword-based searches on publication databases (i.e., Google Scholar, DBLP, Elsevier Scopus) using combinations of terms such as privacy, auditability/accountability, cryptocurrency/blockchain, crypto-assets, payment systems/CBDC, and specific mechanism keywords (e.g., zero-knowledge proofs, anonymous/confidential transactions, conditional/revocable anonymity, tracing/AML compliance).
After searching and scraping, we selected and reviewed about $70$ publications for this survey. Section A.2 provides information about the most popular publication venues. Table LABEL:tab:sota-cryptocurrencies summarizes the analysis of work related to privacy preserving digital payment systems and crypto-asset systems, whereas Table LABEL:tab:sota-mixing reports the analyzed mixing protocols.
A.2. Most Popular Publication Venues
Table 3 summarizes the most common publication venues among the reviewed works. We report only the most frequent venues (with $2+$ works), including journals (e.g., IEEE Transactions on Dependable and Secure Computing, Future Generation Computer Systems, IEEE Transactions on Information Forensics and Security), international conferences (e.g., Financial Cryptography and Data Security (FC), ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P)), and other (non-peer reviewed) sources, such as pre-prints and whitepapers.
Table 3. Most popular publication venues among the reviewed works.
Venue Type Number of Publications Frequent Venues ($\geq$ 2 works)
Journal 16
IEEE Trans. Dependable Secur. Comput. (6)
Fut. Gen. Comp. Sys. (2)
IEEE Trans. Inf. Forensics Secur. (2)
Conference 33
FC (8)
ACM CCS (5)
ESORICS (3)
IEEE S&P (3)
AFT (2)
ASIACRYPT (2)
CRYPTO (2)
Other 16
IACR Cryptol. ePrint Arch. (8)
Whitepapers (7)
A.3. Analysis of Mixing Protocols
Similar to Table LABEL:tab:sota-cryptocurrencies, which surveys privacy-preserving digital payment and crypto-assets systems, Table LABEL:tab:sota-mixing focuses specifically on mixing protocols. The table serves as a reference to facilitate comparison across existing approaches in the literature. The corresponding symbols in Table LABEL:tab:sota_mixing_legend extend those introduced in Table 1.
Table 4. Additional labels used in Table LABEL:tab:sota-mixing beyond those defined in Table 1.
Group Legend
Protocol
No-Trust Assumption Accountable
Other
Cryptographic DC-net Dining cryptographers network HDC-net Harmonized DC-net
$\hookrightarrow$ Primitives MPC Multi-party computation TEE Trusted Execution Environment
VS Verifiable shuffling
Table 5. Classification of existing mixing protocols (see notation in Table LABEL:tab:sota_mixing_legend).
Protocol Privacy Other
Paper
Year
Decentralized
Blockchain Compatibility
No-Trust Assumptions
Implemented
Different Denominations
Sender Anonymity
Receiver Anonymity
Transaction Value Confidentiality
Unlinkability
Cryptographic Primitives
CoinJoins (Maxwell, 2013) 2013 $\times$ $\times$ $\times$ $\times$ $\times$ DS
CoinShuffle (Ruffing et al., 2014) 2014 ✓  (B) $\times$ $\times$ $\times$ $\times$ $\times$ DS, PE
Mixcoin (Bonneau et al., 2014) 2014 $\times$ ✓  (B) $\times$ $\times$ $\times$ $\times$ $\times$ DS
Xim (Bissias et al., 2014) 2014 ✓  (B) $\times$ $\times$ $\times$ $\times$ PE
Blindcoin (Valenta and Rowan, 2015) 2015 $\times$ ✓  (B) $\times$ $\times$ $\times$ $\times$ $\times$ BS
CoinShuffle++ (Ruffing et al., 2016) 2016 ✓  (B)  [^33] https://github.com/decred/cspp $\times$ $\times$ DC-net
TumbleBit (Heilman et al., 2016) 2016 $\times$ ✓  (B) $\times$ $\times$ MPC, ZK
SecureCoin (Ibrahim, 2017) 2017 ✓  (B) $\times$ $\times$ $\times$ $\times$ SA
ValueShuffle (Ruffing and Moreno-Sanchez, 2017) 2017 ✓  (B) PE, SA
CoinParty (Ziegeldorf et al., 2018) 2018 ✓  (B) $\times$ $\times$ $\times$ $\times$ $\times$ DSt, Mixnets, MPC
Liu et al. (Liu et al., 2018b) 2018 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ RS
Möbius (Meiklejohn and Mercer, 2018) 2018 ✓  (E) $\times$ $\times$ RS,SA
Obscuro (Tran et al., 2018) 2018 $\times$ ✓  (B)  [^34] https://github.com/BitObscuro/Obscuro $\times$ $\times$ $\times$ TEE
MixEth (Seres et al., 2019) 2019 ✓  (E)  [^35] https://github.com/seresistvanandras/MixEth $\times$ $\times$ $\times$ $\times$ VS
AMR (Le and Gervais, 2021) 2021 ✓  (E) $\times$ $\times$ $\times$ $\times$ $\times$ ZK
BlindHub (Qin et al., 2023) 2023 $\times$ ✓  (B)  [^36] https://github.com/blind-channel/blind-hub $\times$ $\times$ $\times$ BS*
UCoin (Nosouhi et al., 2023) 2023 ✓  (B) $\times$ $\times$ $\times$ $\times$ $\times$ HDC-net
Wang et al. (Wang et al., 2023) 2023 $\times$ $\times$ $\times$ $\times$ MT
A.3.1. Protocol
System Architecture
Both centralized (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015; Qin et al., 2023) and (Heilman et al., 2016; Liu et al., 2018b)) and decentralized solutions (e.g., (Ziegeldorf et al., 2018; Bissias et al., 2014; Ruffing and Moreno-Sanchez, 2017; Nosouhi et al., 2023; Wang et al., 2023) and (Ruffing et al., 2014)) have been proposed to implement mixing protocols. In centralized solutions, a third party (mixer or tumbler) provides the mixing service for a fee. While such designs are simple and support large anonymity sets, they introduce risks such as availability, theft, and potential privacy breaches. Therefore, different trust models have been proposed to mitigate these concerns (see Section 4.1.4). Decentralized mixers remove the need for a trusted party by using cryptographic protocols or peer-to-peer coordination, enhancing security and censorship resistance, but they often incur higher complexity, require user coordination, and typically support smaller anonymity sets (e.g., Xim (Bissias et al., 2014), CoinShuffle++ (Ruffing et al., 2016), and ValueShuffle (Ruffing and Moreno-Sanchez, 2017)). Both centralized and decentralized approaches face challenges such as Sybil attacks [^37] In a Sybil attack, an adversary injects fake participants into the mix to degrade anonymity. and regulatory scrutiny.
Trust Assumptions
As regards centralized mixing protocols, since the third party represents a single point of failure, different protocols focused on different threat models, where the mixer may be trusted and accountable (e.g., (Bonneau et al., 2014; Valenta and Rowan, 2015)) or untrusted (e.g., (Heilman et al., 2016)). For instance, Mixcoin offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Conversely, Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. The rise of off-chain solutions has let to the exploration of alternative mixer definitions. For instance, BlindHub (Qin et al., 2023) implements mixing within payment channel hubs, where the hub acts as the tumbler. It employs blind signature-based protocols to hide users’ addresses from the hub, ensuring privacy.
A.3.2. Privacy Features
Unlinkability
Different mixing protocols have been proposed (see Appendix A, Table LABEL:tab:sota-mixing). The naïve approach for mixing transactions consists in a centralized service that collects transactions, and publishes on the ledger a single large transaction with multiple inputs and outputs. This hides the information regarding which inputs flows towards which output, thus limiting (if not preventing) linkability between sender and receiver. The Mixcoin centralized protocol considers an accountable mixer, which offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Blindcoin (Valenta and Rowan, 2015) extends this protocol by considering blind signatures for the warrants; this hides the link between input and output addresses of a user even from the mixer, achieving full unlinkability. Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. In BlindHub (Qin et al., 2023), mixing is performed within payment channel hubs, where the hub acts as tumbler and uses blind signatures to hide the customers’ addresses from itself. Obscuro (Tran et al., 2018) is a mixing protocol that utilizes Trusted Execution Environments (TEEs) and enables participants to verify the mixer’s integrity via remote attestation. [^38] Obscuso is specifically designed to defend against two threats: the participation rejection attack, where malicious users try to manipulating mixing by withdrawing their deposits before the TEE reads them, and the blockchain forking attack, where adversaries present stale blocks to mislead the mixer. Numerous papers also investigate decentralized protocols. CoinJoin (Maxwell, 2013) is a technique that allows multiple users to combine their transactions into a single joint transaction; however, the author only describes the mixing idea. CoinShuffle (Ruffing et al., 2014) allow to combine multiple users’ inputs and outputs in a single transaction but may compromise anonymity if transaction values differ or the anonymity set is small. Later, Ruffing et al. (Ruffing et al., 2016) proposed CoinShuffle++, which uses a more efficient mechanism for anonymity, i.e., Dining Cryptographers Networks (DC-nets) instead of mix-nets: while a mix-net requires sequential processing, hence a number of communication rounds linear in the number of participants, DC-nets enable to process mixing in parallel, hence it requires only a constant number of communication rounds. Xim (Bissias et al., 2014) is a multi-round protocol for anonymously finding mix partners based on advertisements placed on the blockchain, with the property that no outside party can identify participants that pair up. SecureCoin (Ibrahim, 2017) creates an aggregated temporary deposit address and requires the collaboration of involved peers to generate novel destination addresses and collectively redistribute coins. Leveraging Confidential Transactions (which nonetheless is available only in a soft fork of Bitcoin), ValueShuffle (Ruffing and Moreno-Sanchez, 2017) provides full privacy besides parties anonymity. Confidentiality enables mixing funds of different values without compromising unlinkability. CoinParty (Ziegeldorf et al., 2018) is a decentralized mixing protocol that distributes the mixing process across a group of semi-trusted parties, called mixing peers. They use an MPC protocol to jointly perform sensitive operations (i.e., encryption, decryption, and address shuffling), ensuring that no single party can link input and output addresses. Each participant’s output address is encrypted and shuffled multiple times by different peers. Threshold signatures is used to enable claiming funds upon agreement of a majority of mixing nodes (thus improving protocol robustness). It also includes mechanisms to detect and exclude misbehaving participants, increasing the overall protocol reliability.
Decentralized tumblers have also been designed using smart contracts on Ethereum. Möbius (Meiklejohn and Mercer, 2018) uses ring signatures and stealth addresses to obfuscate the senders and recipients associations. First, senders furnish both their funds and stealth keys to the smart contract; then, each recipient creates a ring signature to withdraw their funds from the contract and transfer them to an ephemeral address. The size of the anonymity set is limited to the size of the ring, and the gas cost of the withdrawing transaction increases linearly with the size of the ring. MixEth (Seres et al., 2019) aims to minimize gas costs by utilizing Neff’s verifiable shuffle (Neff, 2001) and off-chain messages. Once the sender deposits coins, the MixEth contract enters rounds of shuffling and challenging, during which anyone can challenge the correctness of the preceding shuffle. When recipients believe that sufficient shuffling has been performed, they can withdraw their funds from the MixEth contract. Although the shuffling occurs off-chain, MixEth stores the newly shuffled public keys and all the necessary information within the contract to enable anyone to verify the shuffle’s correctness and continue public key shuffling after the corresponding challenge round.We also mention AMR (Le and Gervais, 2021) which disrupts the traceability connection between coins deposited and withdrawn by the same user. Participants deposit a predetermined quantity of coins into a smart contract, which establishes a Merkle tree structure over the deposits. To withdraw coins, a zk-SNARK prove coins inclusion in the Merkle tree. AMR leverages lending platforms to provide users with the opportunity to accrue interest on their deposited assets.
Mixing protocols are still vulnerable to transaction graph analysis to some extent. For example, if the mixing process does not involve adequate randomization or if there are patterns in the amounts or timing of transactions, likely connections can be identified, even after several rounds of mixing. To maximize privacy, participants need to combine mixing protocols with other privacy-preserving techniques (e.g., stealth addresses, payment channels) and follow best practices (e.g., distributing coins across multiple addresses) to reduce the risk of being traced through graph analysis. Further details on mixing solutions for Bitcoin and Ethereum, e.g., in (Arbabi et al., 2023), and on transaction graph analysis, e.g., in (Ron and Shamir, 2013; Chan and Olmsted, 2017) and in (Fleder et al., 2015).
A.4. Emerging Cryptographic Primitives
In addition to the cryptographic tools discussed in the main survey, several advanced primitives offer promising tools for future research in privacy-preserving digital payment systems. Although not yet widely adopted in practical systems, these techniques may enable more flexible or robust privacy features in regulated or decentralized contexts.
A.4.1. Anonymous Credentials
Systems like Idemix (Camenisch and Van Herreweghen, 2002) and Coconut (Sonnino et al., 2020) allow users to prove possession of attributes (e.g., age, residency, KYC compliance) without revealing their identity. These systems support selective disclosure, enabling users to present only the minimal information required for a given interaction; for example, proving age above $18$ without disclosing birthdate. This selective transparency could be very important for digital payment systems subject to regulatory constraints (e.g., AML, KYC), as users can demonstrate compliance without full deanonymization. Moreover, advanced anonymous credential schemes could support unlinkability, preventing different credential uses from being correlated. However, at the time of writing, efficiently revoking credentials (e.g., after misuse or expiration) while preserving privacy is nontrivial and typically requires accumulator-based techniques or dynamic revocation protocols. As digital payments increasingly intersect with real-world identity frameworks, anonymous credentials offer a promising path for balancing privacy and regulatory needs.
A.4.2. Cryptographic Accumulators
Accumulators provide compact representations of sets along with efficient membership proofs. They are useful for constructing privacy-preserving mechanisms such as spent coin tracking, revocation, and anonymous group membership verification. Although some systems already use Merkle trees to prove coin or user inclusion in a list of valid or authorized entities (e.g., (Sander and Ta-Shma, 1999; Gross et al., 2021; Xue et al., 2023; Wang et al., 2023; Elfadul et al., 2024)), we believe that accumulators remain underutilized—particularly in the design of revocation mechanisms, where they could offer efficient and privacy-preserving solutions.
A.4.3. Functional Encryption
Functional encryption (Boneh et al., 2010) allows specific functions to be computed on encrypted data without revealing the underlying inputs. In contrast to traditional public-key encryption, where decryption reveals the entire message, functional encryption supports function-restricted decryption keys. In the context of digital payment systems, this could enable selective disclosure, e.g., allowing regulators to verify policy compliance or transaction conditions without full access to transaction details. For instance, a regulator could decrypt only whether a transaction exceeds a policy-defined threshold without learning the actual value. Potential applications may include: (i) allowing a central bank to verify that a transaction is below a predefined limit, without revealing the amount; (ii) enabling eligibility verification for financial services based on encrypted KYC attributes; and (iii) constructing privacy-preserving tax collection mechanisms that expose only aggregate (or policy-relevant) outcomes. Despite its promise, functional encryption remains largely theoretical for high-complexity functions and faces efficiency and key management challenges in practical deployments.
A.4.4. Time-Lock Encryption
Time-lock encryption (Liu et al., 2018a) delays access to encrypted information until a predefined time has elapsed. Potential applications include time-delayed auditability, sealed transactions, or rate-limited disclosure of account (sensitive) data.
A.4.5. Witness Encryption
In its canonical form, witness encryption (Garg et al., 2013) allows one to encrypt a message with respect to an instance $x$ of a (nondeterministic polynomial time $NP$) language, $L\in NP$, such that decryption is possible if and only if the decryptor possesses a valid witness $w$ for the statement $x\in L$. This capability opens the door to a number of expressive access control mechanisms that could be beneficial in the context of digital payment systems. One promising application is signature-based witness encryption, where the $NP$ statement encodes a threshold condition over a set of public keys. For example, the instance $x$ could describe a set of $n$ public keys, and the corresponding witness would be any subset of $k$ valid signatures under distinct keys. This effectively enables a form of threshold encryption—where decryption requires authorization from at least $k$ participants—without any setup phase for correlated key generation or key sharing, unlike traditional threshold crypto-systems. While a downside is that ciphertext size can grow with the complexity of the underlying relation (e.g., linear in $n$), recent constructions have demonstrated ciphertexts of size $O(\log n)$, under specific assumptions (i.e., indistinguishability obfuscation for Turing machines), making this approach increasingly practical (Avitabile et al., 2024). Another line of research could explore witness encryption with hidden statements, where the encrypted message can only be decrypted by someone knowing a witness for a secret statement, while the statement itself remains concealed. This property could offer a powerful tool for regulatory compliance. For example, policy enforcement could be realized by requiring users to provide ZKPs of compliance (i.e., knowledge of a valid witness), without revealing the rule being enforced. Such a mechanism may be appealing to regulators who wish to enforce certain constraints (e.g., AML/KYC policies, tracing) without disclosing the exact criteria (e.g., (Mu and Vasudevan, 2024)). To this end, users must include a secret witness with each transaction, attesting to their compliance with the policy. To conclude, witness encryption could provides a rich foundation for designing advanced privacy and access control features, especially in contexts where minimal trust assumptions, fine-grained control, and privacy coexist.
These emerging primitives highlight exciting directions for foundational research in the design of privacy-preserving digital payment solutions. While most remain experimental, their potential to address regulatory, architectural, or usability gaps makes them worth further exploration.
[…]
Additional References
  • (1)
  • Avitabile et al. (2024) Gennaro Avitabile, Nico Döttling, Bernardo Magri, Christos Sakkas, and Stella Wohnig. 2024. Signature-based Witness Encryption with Compact Ciphertext. IACR ePrint 2024/1477.
  • Boneh et al. (2010) Dan Boneh, Amit Sahai, and Brent Waters. 2010. Functional Encryption: Definitions and Challenges. IACR ePrint 2010/543.
  • Camenisch and Van Herreweghen (2002) Jan Camenisch and Els Van Herreweghen. 2002. Design and implementation of the idemix anonymous credential system. In Proc. of ACM SIGSAC CCS’02. ACM, 21–30. doi:10.1145/586110.586114
  • Fleder et al. (2015) Michael Fleder, Michael S. Kester, and Sudeep Pillai. 2015. Bitcoin Transaction Graph Analysis. arXiv 1502.01657. doi:10.48550/arXiv.1502.01657
  • Garg et al. (2013) Sanjam Garg, Craig Gentry, Amit Sahai, and Brent Waters. 2013. Witness Encryption and its Applications. IACR ePrint 2013/258.
  • Heilman et al. (2016) Ethan Heilman et al. 2016. TumbleBit: An Untrusted Bitcoin-Compatible Anonymous Payment Hub. Crypt. IACR 2016/575.
  • Ibrahim (2017) Maged Hamada Ibrahim. 2017. Securecoin: a robust secure and efficient protocol for anonymous bitcoin ecosystem. Int. J. Netw. Secur. 19, 2 (2017), 295–312.
  • Le and Gervais (2021) Duc V. Le and Arthur Gervais. 2021. AMR: autonomous coin mixer with privacy preserving reward distribution. In Proc. of ACM AFT’21. ACM, 142–155. doi:10.1145/3479722.3480800
  • Liu et al. (2018a) Jia Liu, Tibor Jager, Saqib A. Kakvi, and Bogdan Warinschi. 2018a. How to build time-lock encryption. Designs, Codes and Cryptography 86, 11 (2018), 2549–2586. doi:10.1007/s10623-018-0461-x
  • Liu et al. (2018b) Yi Liu, Xingtong Liu, Chaojing Tang, Jian Wang, and Lei Zhang. 2018b. Unlinkable Coin Mixing Scheme for Transaction Privacy Enhancement of Bitcoin. IEEE Access 6 (2018), 23261–23270.
  • Meiklejohn and Mercer (2018) Sarah Meiklejohn and Rebekah Mercer. 2018. Möbius: Trustless Tumbling for Transaction Privacy. Proc. Priv. Enhancing Technol. 2018, 2 (2018), 105–121. doi:10.1515/POPETS-2018-0015
  • Mu and Vasudevan (2024) Changrui Mu and Prashant Nalini Vasudevan. 2024. Instance-Hiding Interactive Proofs. IACR ePrint 2024/776.
  • Neff (2001) C. Andrew Neff. 2001. A verifiable secret shuffle and its application to e-voting. In Proc. of the ACM CCS’01. ACM, 116–125. doi:10.1145/501983.502000
  • Ruffing et al. (2014) Tim Ruffing, Pedro Moreno-Sanchez, and Aniket Kate. 2014. CoinShuffle: Practical Decentralized Coin Mixing for Bitcoin. In Proc. of ESORICS’14. Springer, 345–364. doi:10.1007/978-3-319-11212-1_20
  • Tran et al. (2018) Muoi Tran, Loi Luu, Min Suk Kang, Iddo Bentov, and Prateek Saxena. 2018. Obscuro: A Bitcoin Mixer using Trusted Execution Environments. In Proc. of the ACSAC’18. ACM, 692–701. doi:10.1145/3274694.3274750
  • Wohlin (2014) Claes Wohlin. 2014. Guidelines for snowballing in systematic literature studies and a replication in software engineering. In Proc. of EASE’14. ACM, Article 38, 10 pages. doi:10.1145/2601248.2601268